A Complete Bibliography of Publications in Transactions of the American Fisheries

Society: 1980–1989

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
             WWW URL: http://www.math.utah.edu/~beebe/

30 November 2021 Version 1.00

Title word cross-reference

14 [GA83]. _{12} [LL82]. α [JH82]. F_{1} [BH86, FLK^{+}87, KWP88, MM88a, Mac89, MPP82, YMH83]. F_{2} [FLK^{+}87]. × [AS83, AV84, MPP82, STS87, YMH83].

-ATPase [BE82, HCFE81, Joh83b, SBJD83].

12th [Dab89]. 17 [SK86]. 17-MW [SK86].

80-Year [IM83].

Abalone [FB83]. Abdominal [MS88a]. Ability [JST^{+}81]. Abiotic [DSP84]. Abundance [BA83, Bar82, BM83a, CB88a, Cha86, FCBTSR83, Goo85a, GDH83, HBD88, KO83, MJM^{+}83, Mil81, NP83, PS81a, QB85, TCKW83, WS82, Wat83]. Abuse [Ryd82]. Access [Bed87]. Accommodations [NN82]. Accumulated [MHA81]. Accumulation [MS88b, RG86, WMM81]. Accurately [LJL86]. Acid [FD89, JH83, KP87, MTB81, RJ85, THV87]. Acid-Exposed [RJ85]. Acidic [CMNS85, Hai81]. Acidified [MJL89]. Acidities [PMR86]. Acidity [CLH^{+}86, GMB^{+}87, WFM^{+}89]. Acids [HRN83, SMLM83]. Acquisition [GGS82]. Action [Van87a]. Actions [DT84b]. Activities [BL86, Cip83, SBJD83]. Activity [BE82, CB88b, DGK^{+}88, HCFE81, HM87b, Joh83b, Kol85, SWM87, SN84, WC83]. Acton [PWM81a]. Acuity [MCR84]. Acute [BSS86, BS87, BHN84, BMHW81, HTMZ85, JH83, MSPS81, MBB^{+}84, SW88b, TR83, TRP83]. Adaptation [Vet82]. Adenine [VHH86]. Adenosine [KMTU84]. Adhesive [NM88]. Adjacent [WO86]. Adjustments [Eri84]. Administered [MS88b]. Administration [GRSD83]. Admixture [CJ85]. Adult [BWQ84, BQW^{+}86, BP83, CFCSI84, FG89, HM87b, Mar87a, PEMM87, SH84, TS84, WQB82]. Adulthood [WM86]. Adversely [DKT88]. aeglefinus [KBDP85]. Aeration [CW82]. Aerator [AB82]. Aerators [SB83]. Aerial [NG81]. Aeromonas [CRF83, Cip83, CH86, DS85, HER81, TGRA82b, TGRA82a]. Affect [DKT88, RW86]. Affected [DHVG87, HMMJ81a, PMR82]. Affecting [Min89, Tan85]. Affects [NS84a]. African [Bay81]. after [BNW88, BQW^{+}86, CEG88, CW88, Ell86, FS85, Joh83b]. against [CH86, Dee81, WM86]. Age [BM87, Bar85, Bea84, BMW89, Bea81, BM83b, BM88, Bow89, BC89, Cam84, CLSRS89, FC87, FK88, HE84, HJ83, HG85, HBB84, HH87, HSM85, Joh83a, KT87, LN87, MHLL87, MMNW81, Mil82, Mil83, MPSC89, NG86, PLC86, Rim85, RH82, Ros88, SWM87, SK81, Ser82a, TGRA82b, TGRA82a, War84]. Age-0 [FK88, MMNW81, Mil82, Mil83, MPSC89, Rim85, SWM87, Ser82a, TGRA82b, TGRA82a]. Age-1 [KT87]. Age-Determination [BC89]. Agglutinating [HER81]. Agglutinins [Cip83]. Aggregate [Bow81]. Aggregation [PM81]. Aggregations [Arg82]. Aging [PD88]. Agonistic [RM87a]. Air [SH86]. Alabama [Boy82, HSR^{+}86, Mar83, Mos85, PM83, TSD81]. Alabama-Georgia [PM83, TSD81]. Alaska [Dol86, Dol87, HMSK86, HG89, KG89, LF89, MM88b, MRG87]. Alaskan [BM87, BSS^{+}88, Cra84, GSMH88, Har86, WB85]. Albacore [Bea81]. Alberta [OIWD81]. Alevins [AB83, NB85, PMR86]. Alewife [FD85, HHB81a, KN81, MB82]. Alewives [Arg82, BO89, BMM^{+}87, HS89, NN82, OS86, SB86c]. Alga [Ger84]. Algorithm [HH87]. Alkalinity [HB81]. All-Female [HDGE82]. Allele [HDS86]. Alleles [CB89]. Alleviation [CTSD84a]. Allocation [Hig87, Hil89]. Allozyme [KM87a, KM87b, MBM89]. along [BA85, Res89, RPI89, SHCMS83]. Alosa [LKF82, Mil81]. Alteration [YHW89]. Alterations [JH83, TAS86]. Alternative [CB89]. Aluminum [CLH^{+}86, KP87, WFM^{+}89]. Amazon [Bay81]. Ambient [PMR86, RS83, SS84]. America [FCL^{+}83b, QB85, Sco88]. American [AR82, BHR85, CM83, CWCR89, CB83, DGK^{+}88, Edi88, Enn84a, Enn84b, Enn86, FL81, FI88, HE84, HSB^{+}83, HBB84, Jes87, LF83, MBM89, Noa88, OK86, RC87, SB86b, TK85]. americanus [Enn84a, Enn84b, Enn86]. Amino [HRN83]. Ammonia [ACG^{+}82, HA83a, HA83b, SL86, SFS83, TR83, TRP83, TRL^{+}84, TRM^{+}86]. Ammonium [HB81]. among [FW86, HDS86, HHEM86, HKA83, KM87a, KMKM89, LM82, MM88a, Mac89, SÍW89, WJ87, WB85]. Anadromous [BA85, LKF82, Noa88, RPI89]. Analyses [MPP82]. Analysis [AV84, BB83, BCNW87, BOWY83, Bro85, CM83, EMC^{+}85, Fle85, FB83, GV88, HWGF82, HO83, Hus87, Ker82, LB82, LM81, MHLL87, Mar87b, PH89, RF88, RCB85, RCH87, RNT85, SHI84, Ser86, TC89, TMF83, Ush87b, ZS88, ZSNC87]. Analytical [MYA81]. Analyzing [EH86]. Anatomy [KBDP85]. Anchor [VO87]. Anchoveta [Lon88]. Anchovy [Bro85, VAKT87]. Anesthetic [AGE82, LGP83]. Aneuploidy [BSDD86]. Angler [SK81]. Anglers [GMW83, Mat81, SJR87]. Angling [DP87, GC81, Hud84]. Anguilla [VJ86]. anguillarum [Dee81]. Annual [Dab89, DM81, Eri84, Eva84, Mar85, ZB84]. Annuli [Sch89, TT82]. Annulus [Pow82]. Anthracene [LB84]. Antibody [CRF83, HER81]. Antivirus [HM87b]. Apache [Mes88]. Appalachian [CLS87, LM85a, NP83, WS83]. Apparent [KBDP85]. Appearance [FHT84]. Application [FKY84, FI88, LO86, PT88, PH89, SWRM85, TMF83]. Applications [LNK87]. Appraisal [Mar87b]. Approach [Par84, SE87, THV87]. Aquaculture [Par84]. Aquatic [Hai81, HH82c, LM81, MBB^{+}84, PAH89]. Aquilla [MMI88]. Arctic [BM87, FP89, HC84, HG89, Kay89]. Area [EH86, HPG83, RW86, Her85]. Areas [HSR^{+}86, HHB81b, RVV85, WCM88]. argus [GLC82]. Arizona [Mar87a]. Arkansas [AB83, Bro86]. Arrow [MMR89]. Art [FCL^{+}83b]. Artificial [GC87, HCFE81, LJL86, Mes88]. Ash [WQB82]. Aspects [DT84b, Esh87, PWM81a]. Aspidoparia [MM85]. Assemblage [MB85, OT86]. Assemblages [HG87]. Assess [JST82, JST^{+}81, SSU86]. Assessing [HM86, MJE89, WRH^{+}85]. Assessment [BJCM87, Gal87, Hus87, MSC86, PD88, PG87, SE87, TDG^{+}87, TL87, Van87b, Wal81]. Assimilation [Bow81, Ger84, HN84]. Associated [CTSD84a, Gar83, HH82b, WT88]. Assumptions [Mab89]. Atlantic [MH84, Bar85, BA85, Cun89, DB86, EG86, GT85, Joh83b, KT85, Kol85, MDS84, NM82, ØKF85, PS81b, PMR82, PMR86, RPI89, Ros88, SÍW89, SW85, Sos82, SK86, Tis89]. ATPase [BE82, HCFE81, Joh83b, SBJD83]. atratuli [Kir81]. Auburn [Boy82]. Australian [BSM82]. Autumn [BO89]. Availability [AMP82, JSC88]. Available [HM87a, Res89]. Avoidance [BP83, DGS^{+}85, EB86, GCCL81, GGS82, RC87, SW88b]. Avoided [PB83]. Avoiding [GC87]. aztecus [HH82b, MZM89].

B [ZR81, LL82]. Back [Fed87]. Backbone [BHL^{+}85]. Bacterial [ACG^{+}82]. Bacteriophage [FJT89]. Bacterium [HB86]. Bahamian [CA83]. Balance [SL86]. Balances [WW86b]. Baltic [Eri84]. Band [Smi84]. Banded [Ros89]. Banding [PZI89]. Bank [Her85, KBDP85, OT86]. Barge [Hol86]. Barnacle [HMSK86]. Based [Bar85, Bor83, Car87, FKY84, FI88, GGS82, HD88, KM87a, KM87b, RCH87, ZS88]. Basic [Lof87]. Basin [QB85]. Basins [Bro86]. Basis [KMTU84, Ran87, SJR87]. Bass [AMP82, AV84, AW85, BL86, BM83a, BCNW87, Bod85, Bor83, BA85, BZP81, Bre88, BJCM87, BH86, BHL^{+}85, CMNS85, Car87, CTSD84a, CTSD84b, CWSM86, CVC85, CL82, Col84, CM81, CP81, CV89, CD83, CZCP84, Cou85a, Cou85b, Cou87, CC81, Dey81, EWE^{+}81, Fab87a, Fab87b, FG89, FLK^{+}87, FB88, FS83, FH89, Gof86, GCC85, Goo85a, Goo85b, GO86, GA85, HPHF85, HZB^{+}89, HSR^{+}86, HER81, HB86, HCG89, HO83, HSM85, INKP87, JBL88, KW89, KE85, KLJ84, Ker83, KBH85, KHS81, Koh81, KWP88, MMI88, MM88a, Mac89, MBM86, MWMSH85, MHS85, Mat85, MHEG89, MW81, MJL89, MO88, MCHL81, MK89, MHH^{+}82, MW86, MS82, MS84a, Mil82, Mil83, MRC81, MRN83, MWSR83, Mos85, PN87, PHD85, PCW83, PFT^{+}85, PRCM84, Ran86, RW81a, RNT85, RM87b, SRM83]. Bass [SS82a, SJ82, SFW83, SSGT84, SZ85, SCSR86, Ser82b, Ser84, SHBM^{+}81, SWRM85, SA82, SHM83, SCH81, SKMK85, Sto86, SHCMS83, TT82, TSD81, TR89, Uph89, WOR85, WJ87, WC86, WCSM86, WF81, WDS85, Wre84]. Bay [FS85, Gar83, Gof86, Goo85a, MHLL87, Mar83, MWSR83, Rec86, RM87b, VAKT87, Fab87a, HPHF85, HZB^{+}89, Koh81, LB82, PFT^{+}85, WMRH86]. Beaufort [Cra84, HC84]. Behavior [Bac84, BJCM87, Enn84b, Fos85, Gal86, GCCL81, HO83, JV87, Kav84, KW89, LO83, MK89, OK86, RM87a, WS88, WQB82]. Behavioral [HA86, MS89]. Behaviour [EB86, Eri84, Gib84]. Below [Gri81, GCC85, RHU82, Rug86]. Benefit [PM83]. Benefits [Gre87]. Benthic [Fis87, IZ83]. Between [HFKR81, MPSC89, PS81a, AK84, BB83, BE82, CC85, Cip83, CH86, CL82, DCR89, GV83, Goo85a, GC87, KG83, LM85b, LD86, MHH^{+}82, PM83, Pin82, Rig82, RM87a, SS82a, ST83, Sos82, WD88, WO86]. Bias [GC84]. Bias-Elimination [GC84]. Big [IM83]. Bighead [AS83, BB83, MPP82, YMH83]. Bight [EG86, SW85, MH84, SP88]. Bimodal [DC82]. Binary [TS84]. Binary-Coded [TS84]. Bioassay [MJE89]. Biochemical [AM89, BWMB88, BMW89, BHL^{+}85, HSR^{+}86, MPP82, PCW83, SLJ^{+}84, WCSM86]. Biodiversity [Fle89]. Bioeconomic [Ret87]. Bioenergetic [Dab89]. Bioenergetics [BST89, Car87, JSC88]. Biological [AR82, HMJ87b, ZSNC87]. Biology [BM83b, CM81, FL81, MMR89, RHU82, SW85]. Biomass [BR86, CFCSI84, Cou81, GMB^{+}87, HHB81a, ZS88]. Biomass-Based [ZS88]. Biotic [DSP84, FKY84, JST^{+}81, KYFS87, LO86]. Black [HH82a, MJL89]. Black-Spot [HH82a]. Blackfish [JV87]. Bladder [MMB81]. Bloater [RCB85, RCH87]. Bloaters [CC84]. Block [SCD81]. Blood [EMC^{+}85, MJL89]. Blue [BOWY83, BOY84, SZ85, ZG89]. Blueback [CB83, KP87, OK86, TK85]. Bluefish [FGB^{+}88]. Bluegill [TSC84]. Bluegills [BH83, Bar82, FK88, FS83, Gar81, GSC81, GB86, GC87, HA86, JBL88, SS82a, SJ82, SZ85, SJ89a, Sch89, WH82b]. Bluehead [DKT86]. Blueline [RH82]. Bluntnose [Gal83]. Boat [HWD88, Mat81]. Body [BWMB88, BSN88, GMB^{+}87, GA85, MHH^{+}82, Mil82, MPSC89, WOR85]. Bolivia [WT88]. Bone [HMMJ81b, HMMJ81a, HR88, MMB81, MHH^{+}82]. Bonefish [CA83]. Bones [HDLG88]. Bonytail [MBM89]. Book [Ano81a, Ano81b, Ano81c, Ano81d, Ano81e, Ano82a, Ano82b, Ano82c, Ano82d, Ano82e, Ano82f, Ano83a, Ano83b, Ano83c, Ano83d, Ano83e, Ano84a, Ano84b, Ano84c, Ano84d, Ano84e, Ano84f, Ano85a, Ano85b, Ano85c, Ano85d, Ano85e, Ano85f, Ano86a, Ano86b, Ano86c, Ano86d, Ano86e, Ano87a, Ano87b, Ano87c, Ano87d, Kre86a, Ush87a]. borealis [PLC86]. Bottom [Arg82, DRFL89]. Bottom-Sampling [DRFL89]. Brackish [Eri84, HE84]. Brackish-Water [HE84]. Braided [GD85b]. Brain [ZSNC87]. Branched [HRN83]. Branched-Chain [HRN83]. Brazil [Bay81]. Breaks [SH86]. Bream [CFCSI84, Tan85]. Breeding [BSYW84]. Brevoortia [SWT^{+}85]. Briarosaccus [HMSK86]. Brief [BQW^{+}86]. Bright [EMC^{+}85]. Bright-Field [EMC^{+}85]. British [Bea84, BH88, FC87, LF89, MMR89, MR89, SJ89b, BWMB88, MM88b]. Bromide [AQM^{+}82]. Brood [DSHB83, MBM89]. Broods [MHS88]. Brook [CHW87, CLH^{+}86, CG84, CCP87, FW86, GS82b, HMJ87a, HMJ87b, JB87, KR85, LM85a, Mar84, MO86, STS87, SB86b, Wat83, WS83, WM83, YHW89]. Brown [Bac84, BSP87, CLS87, FW86, HJ83, Heg88, HH82b, Jon85, KM87b, Mar84, Mes88, MZM89, STS87, SD83, Sos82, STP82, Wat83, WCM88, WM83]. Browns [KBDP85]. Bubble [Hau86]. Budget [CCP87]. Budgets [Boy85a, PMR86]. Bullheads [BSP87]. Buoyancy [Sos82]. Buoyant [LJL86]. Burbot [BSS^{+}88, MJM^{+}83]. Burdens [WOR85, YMH86]. Burst [Fui86, TM85a]. Burst-Swimming [Fui86]. By-Products [Fos85].

C [GA83, HMMJ81a]. C-Glycine [GA83]. Cadmium [FV82, HMJ87a, HMJ87b]. Cadmium-Saturation [HMJ87a]. Caged [Gal83]. Calcein [WBD87]. Calculating [Car82, SB83]. Calculations [AB82]. California [BVBM87, BC89, Erm86, MB85, SZS82, SKMK85]. callosus [HMSK86]. Caloric [MR83]. Canada [FP89, Joh89, RM87b]. Canadian [BM87, BP87, CM83]. Canal [KHS81]. Cannibalism [Bro85]. Canopy [MHA81]. Canyon [KZ83]. Capabilities [SJ81]. Cape [Bro85, HHB81b]. Capelin [Nak87]. capensis [Bro85]. Captive [GD85a, Gal86, Hic82]. Capture [BOWY83, BOY84, McG85, MCR84]. Carbon [ACG^{+}82]. Care [BH83]. Carnation [BH88]. Carolina [FB88, HE84, BM83a, BOWY83, HHB81b, Lem85, Pow82, RH82, Ros88, Ros89, WCM88]. Carp [AS83, ATH86, BB83, BSYW84, JSYW83, LVNM82, LNV83, MPP82, MTBB81, PAH89, TC89, WW86b, YMH83]. Case [BP87, HWD88, SHM83]. Catadromous [Noa88]. Catch [BA83, CC85, Jen85, Mac82, MI89, SJR87]. Catch-Curve [Jen85]. Catchability [EH86, LKF82, Nie83, PS81a, SHP^{+}85]. Catches [PCC85]. Catching [DP87]. Catfish [AB83, Boy85a, BB84, BS81, BB82, DCR89, DSP84, DRRP82, DSHB83, DEAT87, EMC^{+}85, FCL^{+}83a, GS82a, GRSD83, HDS86, HMMJ81a, HM87b, HBD88, LM85b, LLBA86, MS86, MS88a, MB83, ST83, SL86, SGK^{+}89, SMLM83, SM84, TGRA82b, TGRA82a, WBH85, WLC81, WLC82]. Cation [PMR86]. Catostomid [Sny83]. Catostomidae [Sta88]. Catostomids [Sco88]. Caused [HB86, New81]. Causes [JM82, LM82]. Causing [Dia87]. Central [Car83, MW86, NM82, Rat81, CLSRS89, OWNH88]. Centrarchid [Car82]. Century [Cha86]. Ceratomyxa [BSZF83, HHEM86, Rat81]. Chain [Dup82, HRN83, WRH^{+}85]. Challenge [MPAR86, MO86]. Champlain [FL81, LF83]. Change [Lon88]. Changes [BW89, Bow89, BM83c, Col84, Edi88, HDS86, HA86, HG87, IM83, Joh83b, KW89, Lue86, MS89, NB85, New81, PTMG89, SW88b, WCM88, ZB84]. Changing [Hig87]. Channel [AB83, BSN88, Boy85a, BB84, BS81, BB82, DCR89, DSP84, DRRP82, DSHB83, DEAT87, EMC^{+}85, FCL^{+}83a, GS82a, GRSD83, HDS86, HMMJ81a, HM87b, HBD88, LM85b, LLBA86, MS86, MS88a, MB83, ST83, SL86, SGK^{+}89, SMLM83, SM84, Tay88, TGRA82b, TGRA82a, WBH85, WLC81, WLC82]. Channels [Mes88, Rim85]. Char [BM87, FP89]. Character [Mab89]. Characteristics [BW89, Dol87, Kus81, LF89, LA84, Ran86, RCH87, SCSR86, VJ86]. Characterization [CTSD84a]. Characterizing [TB87]. Characters [BS81]. Chemical [Boy85a]. Chemosensory [Fos85]. Chesapeake [HZB^{+}89, Fab87a, Goo85a, HPHF85, KHS81, Koh81, PFT^{+}85, WMRH86]. Chicken [TC89]. Chinese [LMW81]. Chinook [BSS86, BMW89, BNF82, BQW^{+}86, BE82, CWW^{+}86, CW88, EH83, FV82, HA83a, Har86, HCFE81, HGP87, KMTU84, MPAR86, MSBB88, MS88b, MR89, NG86, NB85, Ors87, Rat81, SW88b, Tay88, WQB82]. Chloride [ST83]. Chlorinated [GCCL81, Mer86, OIWD81]. Chlorine [BB84]. Chlorophyll [JH82]. Chlorophyll- [JH82]. Choice [BVBM87, MCR84]. Cholinesterase [ZSNC87]. Choptank [Uph89]. Chosen [SD83]. Chromosome [PZI89]. Chromosome-Banding [PZI89]. Chronic [BMHW81, DBBF83, FLK^{+}87, SBE84, TRL^{+}84, TRM^{+}86]. Chronically [CGT82, HR88]. Chronobiologic [Par84]. Chronology [FHT84]. Chub [GV83, HR88, KZ83, PH89]. Chubs [GS82b, RB86]. Chum [Bea84, SSU86, WM86]. Cichlid [DTLK84]. Ciguatera [DKT86, DKT88]. cimbrius [Her85]. Circadian [Mei84, NS84a, WM84]. Cisco [NJS86]. Class [Car86, CP81, KBDP85, Ser86]. Classification [Fab87b]. Claw [SH86]. Cleaning [Dol86]. Climatic [SR82]. Closed [ACG^{+}82, NM88]. Clue [BO89]. Clupea [DRV85, GT85]. Clupeid [MD86, SP87]. Coal [DBBF83, DGS^{+}85, MBB^{+}84]. Coal-Derived [MBB^{+}84]. Coal-Liquid [DBBF83]. Coast [Res89, BA85, MHH^{+}82, SRM83, War84]. Coastal [Cra84, Fab87a, GT85, HC84, Koh81, MYA81, RVV85, Tro89]. Cod [Bea81, DRV85, KT85, ØKF85]. Coded [TS84]. Coefficients [PS81a, SHP^{+}85]. Coexisting [CB83]. Coho [BSN88, BNW88, BWQ84, BH88, CCMD81, Dee81, Dol86, Dol87, FMS81, FW86, HHEM86, HDGE82, MHS88, MBRK81, Mol83, Ors87, PEMM87, PSBZ86, RS83, RSE87, RJS85, RM87a, SHI84, SBE84, SS82b, SS84, TM85a, TM85b, Tay88, TGC^{+}86, YMH86]. Cohort [KG89, Ser86, ZS88]. Cohorts [DC82]. Cold [ZG89]. Collected [Joh89]. Collecting [MSBB88]. Collection [FD89]. Colorado [AM89, BP85, CB88b, KZ83, PB83, SB86a, WRH^{+}85]. Columbia [LF89, Bea84, BWMB88, BH88, Cha86, FC87, MSBB88, MM88b, MMR89, MR89, Rug86, SJ89b]. columnaris [JP81, MBM86]. Comanche [EE89]. Combined [BB84]. Coming [Fed87]. Comments [Ros88]. Commercial [Bow89, BB82, ED88a, Goo85a, GAT82]. Committee [TDG^{+}87]. Common [MTBB81, Noa88]. Communication [MCH87]. Communities [FKY84, MHA81]. Community [Cou81, DTLK84, GC81, TMF83]. Comparative [BC89, BH86, BHL^{+}85, CD83, Gal87, Hic82, ML82, MTB81, MBB^{+}84]. Comparison [Bay81, Boy81a, Cam84, DRFL89, FF86, GC83, GCCL81, Jen85, KM87a, PEMM87, PT88, Smi85, VHH86]. Comparisons [EE89, Sos82]. Compartment [CH85a]. Compendium [Tro89]. Compensation [Vet82]. Competition [CL82, FW86, OT86]. Competitors [GMB^{+}87]. Complete [RW83]. Components [MBB^{+}84, MBRK81]. Composition [BHL^{+}85, HG85, HH87, Mac82, MMB81, Mil82, Rad84, TAS86]. Computer [GGS82]. Computer-Based [GGS82]. Concentration [JH82, Mil82, RS83, SS84]. Concentrations [BSP87, CGT82, CMNS85, CW88, DHSA83, GB86, Kol85, PB83, SBJD83]. Concept [DRH88]. Concepts [JO82, Ryd82]. Condition [Con89, Daw88, FGB^{+}88, KM84, McG85, PC85, WJ87, WH82b]. Conditions [AMP82, CLL86, CHW87, Mei84, PSBZ86, RTV84]. Conference [Dab89, EB86, Sny84, Mar85]. Confinement [CTSD84b]. Confirmation [BSDD86]. Connecticut [BK85, CB83, OK86]. Consequences [Hai81, KN81]. Consideration [YH82]. Considerations [ZSNC87]. Considered [RW86]. Constant [MI89, STP82]. Constant-Catch [MI89]. Consumers [Ger84]. Consumption [Bre88, Car87, HCG89, LO83, MF81, Rug86, SA82, SB86c]. Containing [SMLM83]. Contaminant [BSP87, HZB^{+}89]. Contaminants [HPHF85, HM86, MHH^{+}82, PHD85]. Contaminated [CV89, RB86]. Contamination [Mer86]. Content [MR83, RB86, SP87, TC89, WJ87]. Contents [Mar87a]. Continental [SWT^{+}85]. Contingent [Hoe87, RTP85]. Continued [CEG88]. Continuous [Bou82]. Contrast [Hin85]. Contributing [SP88]. Contribution [BH88, Fab87a, SW88a]. Control [ACG^{+}82, LNV83, TAS86]. Controls [JM82]. Conventional [JST82]. Conversion [DP85, SB86c]. Convert [HH87]. Cooling [SJ81].

Cooling-Water [SJ81]. Coolwater [LO86]. Cooperation [Tis89]. Coos [MWSR83]. Copper [Bod85, FV82, GGS82, HA86, Kol85]. Coregonine [Dab89]. Coregonus [RCB85]. Corneal [Bod85]. Correct [HH87]. Correction [SB83]. Correlates [SB86b]. Correspondence [GV88]. Corroboration [BST89]. Corticosteroid [DSP84, DP86]. Cortisol [CW88, RS83, SN84]. Cost [BS87, Rei83, SD87, THV87]. Cottonseed [DRRP82]. Counts [NG81]. Cover [Mes88, SZ85]. Crab [BOWY83]. Crabs [BOY84, HMSK86, Res89, RC87, SH86, SJ89b]. Crappie [MJL89]. Creek [BHR85, BH88, Erm86, GS82b, HR88, HSB^{+}83, Kwa81, RB86, SSS86, DCR89]. Creeks [WO86]. Crevice [Gal86]. Criteria [AAC88, MB85]. Critical [Cha88, FLK^{+}87]. Croaker [Ros88]. Croakers [Bar85]. Crossbred [DSHB83]. Crosses [HHEM86]. Crossflow [GW83]. Crude [BQW^{+}86, MR83, WMM81]. Cryogenic [Ker83]. Cryopreserved [KBH85]. Ctenophores [BL83b]. Cucumber [BSM82]. Cues [Fos85]. Cui [SWCB83, Sco88]. Cui-ui [SWCB83, Sco88]. Cultural [SJR87]. Culture [AMCD85, FCL^{+}83a, MMY^{+}82, SHI84, SFS83, WF81]. Cultured [LM82]. Culturing [DS85]. Cumberland [Wal86b]. Cumulative [BSS86]. Cunners [Hic82]. Current [LVNM82]. Curve [CH85a, Jen85]. Cutthroat [BSN88, LAPK87, Lue86, Mar84, MS89, MG88, MYA81, SB86a, Tro89, WMM81, WFM^{+}89]. Cyanoacrylate [NM88]. Cycle [DM81, DGK^{+}88, Eva84]. Cycles [BL83a, Eri84, Hon84, ZB84]. Cyclical [Rec86]. Cyclophosphamide [CRF83]. Cyprinid [FCL^{+}83b, MBM89]. Cyprinidae [MM85]. Cyprinids [Sco88]. Cytological [ATH86]. Cytometric [BSDD86]. Cytometry [AS83, EMC^{+}85, TC89].

Dace [MBC85]. Daily [BNF82, DSM85, DGK^{+}88, EC86, Hon84, MS82, MS84a, TY89, ZB84]. Dam [RHU82, Rug86, SK86]. Damage [Bod85, THV87]. Dams [MSBB88]. Danios [Fui86]. Daphnia [HW81, MF83]. Darters [GD85a]. Data [GGS82, GV88, KM87a, KM87b, Pra88, Res89]. Day [SA82]. Day-to-Day [SA82]. Deaminase [KMTU84]. Debris [AK84, BW89, Ell86]. Decades [Lon88]. Dechorionation [HSG^{+}88]. Decisive [Bal84]. Decline [SKMK85]. Declining [GCC85]. Deep [Fab88]. Deepwater [MJM^{+}83]. Deficit [CCP87]. Define [Cha88]. Definition [CM83]. Deformation [RJ85]. Delaware [KHS81]. Delayed [Hay86, RW81a]. Delivery [PTMG89]. Demersal [OT86, Tan85]. Demographic [Bro86]. Demography [EG86]. Densities [GC87]. Density [DA82, FMS81, MCK86, RA86, Ser82a, Ser82b, SHP^{+}85, SBE84, TSC84, WH82b]. Density-Dependent [RA86, SHP^{+}85]. Dentition [Wal86a]. Dependent [CD83, RA86, SHP^{+}85]. Depensation [KE85]. Depletion [MS89, MS88b, WH84, Wet88]. Deposition [Smi84, STP82]. Depth [MHS85]. Depuration [AQM^{+}82, LB84]. Derived [MBB^{+}84]. Design [Hoe87]. Detect [DS85]. Detected [FJT89]. Detection [KC89, PEMM87]. Determinants [Tay88]. Determination [BM88, BC89, Cam84, Hil88, KLM86]. Determinations [FD89, MB85].

Determine [HSM85, SK86]. Determined [AS83, BB83, DT87, MS84a]. Determining [Mab89]. Detrended [GV88]. Detrital [Bow81]. Detritus [MW88]. Developing [MB85]. Development [BL86, Bor83, BM83c, CEG88, ED86, EB86, HR88, HZ89, KT85, MMB81, MHH^{+}82, MJ83, MRC81, MR82, RW83, RW81a, SSS86, WLC82]. Developmental [BNF82, LAK85, MMR89, PSBZ86, RJS85, STS87]. Device [Bou82]. Dewatered [NB85]. Dewatering [BNF82, RW83]. Dewaterings [BNF82]. Diagnostic [HDLG88, Wal86a]. Dicentrarchus [SSGT84]. Die-off [KN81]. Diel [AB83, Bra86, Fui86, SN84]. Diel-Feeding [Bra86]. Diet [Bra86, Daw88, FGB^{+}88, HN84, IZ83, JTDM87, KE85, KT87, LD86, Rei83, Ros89, Sto86, Wal81]. Diet-Overlap [Wal81]. Dietary [CV89, DP85, DRRP82, GS82a, GV88, HRN83, LL82, Mil82, Mil83, WOR85]. Diets [BL83b, MW88, SMLM83]. Differences [Car83, CH86, Mac89, PZI89, RM87a, WB85]. Different [Car83, HKA83, MW81]. Differential [CLL86, WC86]. Differentiation [BCNW87, BMU87, HR87, KM87a, KM87b, KMKM89]. Digestion [Bow81]. Digestive [BL86, Mar87a, MW88, WGR87]. Dilations [HM85]. Dimensions [Bro87]. Dimorphism [Sta88]. Dioxide [ACG^{+}82]. Dipeptidase [CB89]. Diploid [WW86b]. Direct [GAT82]. Direction [CWCR89]. Directly [GMW83]. Discharge [Gri81, New81, PM81, Rim85]. Discrete [Car86]. Discriminate [Han82]. Discrimination [Fab87b]. Discussion [Bed87]. Disease [HH82a, Hau86, HM87b]. Dispersal [Mar86]. Dispersed [BQW^{+}86]. Dispersion [SHCMS83]. Dispersions [DBBF83]. Displacement [Har87, Heg88]. Dissolved [Bou82, Cob82, Col84, Cou85b, PB83, ST83, WH84]. Dissolved-Gas [Col84]. Dissolved-Oxygen [WH84]. Distribution [BO89, Bow89, CWW^{+}86, CVC85, FG89, FCBTSR83, GDH83, Her85, HS83, Hol86, HKA83, KO83, MJM^{+}83, Mar83, Mat84, MHS85, MS84a, Mil81, RG86, Rim85, Sco85, SHBM^{+}81, TJ83, Tyu87, WS82]. Distributions [BA83, DC82, WCM88, WDS85]. Disturbances [BSS86]. Diurnal [TCKW83]. Divergence [LAPK87]. Diversion [Fle85]. Diversity [FCBTSR83]. Diving [RTP85]. DNA [FJT89, TC89]. Do [Ret87, RCB85]. Does [RW86]. Dogfish [NMC85]. Dolly [Dol86, Ell86]. Dominance [CP81, CG84]. Dosages [AGE82]. Dover [KTKD82, PD88]. Downstream [Har87, MS84b]. Drawing [FG83]. Drift [AB83, CP86, HDS86, NJS86]. Drum [CLSRS89, HA83b, JM82, LHA84, Mat84, Ros89]. Drums [HKA83]. Dry [McG85]. Dual [Ken81]. Due [Hau86, MSC86]. Dummy [CH85b, MS86, SM84]. Dungeness [SJ89b]. Duration [BB84, CC85, GO86]. during [BM83c, ED86, FS85, FGB^{+}88, JM82, KM84, KG89, KR85, MHS85, RS83, RJS85]. Dwelling [PRCM84, TR89]. Dye [PT88]. Dynamics [AQM^{+}82, CC81, FD85, FS85, GLC82, KG89, Kwa81, LM87, OS86, Ros88, SSS86, SSU86, SB86c, Tyl88, VAKT87].

Early [Ald85, Bal84, BJCM87, BM83c, CFCSI84, CLH^{+}86, CP83, CCP87, Dab89, KHS81, KBDP85, KR85, Mar85, MO88, NN82, Smi89, TCS88, TSD81, WW86a]. Early-Life [CP83, NN82]. East [MHH^{+}82]. East-Coast [MHH^{+}82]. Eastern [Kwa81, BHL^{+}85, FCL^{+}83b, TJ83]. Ecological [CP83, CC84, LM81, MENG86, SZS82, SWRM85, WOL83]. Ecology [CA83, CB83, HC84, KZ83, Lac89, LM88, MRTR82, PWM81a]. Economic [AR82, BBW87, HWD88, Mar87b, PM83, PG87, Ran87, TC87, Ush87b, Van87a]. Economics [LNK87]. Ecosystem [Hig87, Lon88]. Ecosystems [CL82, Hai81, TC87]. Ectoparasite [Kir81]. Editorial [Ken89]. Edwardsiella [HB86]. Eel [HSB^{+}83, SB86b, VJ86]. Eels [BHR85, DGK^{+}88, FL81, HE84, HBB84, Jes87, LF83]. Effect [CRF83, Daw88, HCG89, JB87, LD86, Mol83, MD86, SWM87, Wet88, WLC82, ZG89]. Effective [AGE82, FCL^{+}83a, Lof87]. Effects [AR82, BNF82, Bre88, BB84, BMHW81, CWW^{+}86, Cha88, Col84, CW88, CPML85, CCMD81, DBBF83, DKT86, DP85, Dol86, DRRP82, EWE^{+}81, FMS81, FD89, Fui86, Gar81, GB86, GD85b, GC81, Goo85b, GA83, GO86, Gri81, HZB^{+}89, HH82a, Hay86, HM87a, Heg88, HFDL83, Hin85, Hol86, HA83b, JP81, KMC87, KMCB84, Kol85, LNV83, LKF82, LM87, MBM86, MMNW81, MSBB88, McG85, MMB81, Mes88, MJ83, Mil82, Mil83, MR83, MP84, MRC81, MRN83, MHA81, NG86, OIWD81, PSBZ86, PAH89, PH89, RSE87, RW83, RF81, Rim85, RW81a, SFW83, SZ85, SSS86, SBE84, SH86, SFS83, SN84, TAS86, TGC^{+}86, TGRA82b, TGRA82a, TSC84, Uph89, WSL86, WH84, WOR85, WH82b, WMM81, WC83]. Efficacy [BOY84, LGP83]. Efficiencies [MW88]. Efficiency [JV87, Mil82, MD86, PV88, SCD81]. Effluent [SWRM85]. Effluents [OIWD81]. Effort [BA83, PM83]. Egg [BWMB88, Bor83, FI88, KG89, KT85, RW83]. Egg-Hatching [RW83]. Eggs [DRV85, HSG^{+}88, Hay86, HA83b, HPG83, KP87, KMCB84, LVNM82, Mar83, Mer86, MRC81, MR82, MRN83, MD81, NB85, NS84b, RJ85, Ser84, SHBM^{+}81, Uph89]. Eighth [Mar85]. Electric [BLD89]. Electrofishing [MS89, PCC85]. Electrolyte [CMNS85]. Electrophoresis [CWSM86]. Electrophoretic [CJ85, Fab87b, HWGF82, KMTU84, RNT85]. Elemental [BCNW87]. Elevated [MW81, WFM^{+}89, Wre84]. Elevation [CHW87]. Eleven [CB88b]. Elimination [GC84]. Elvers [SB86b]. Embryo [SP85]. Embryonic [KC89]. Embryos [CEG88, HFDL83, JB87, MTBB81, NB85]. Emerald [BB84]. Emerged [TM85a]. Emergence [CB88a]. Emergents [TS84]. Emigrating [MSBB88]. Emphasis [AW85]. Empirical [BGC81, SE87, ST83, WDS85]. Enchelyopus [Her85]. Enclosure [Kus81]. Encroachment [LM85a]. Endangered [AM89, EE89, Lac89, MBM89, PB83, SWCB83]. Endemic [WFM^{+}89]. Endogenous [Qua85]. Endotoxin [TGRA82a]. Energetics [MF81, SB86c]. Energetics-Modeling [SB86c]. Energy [AMP82, CCP87, FD85, Qua85, TGC^{+}86, WW86b]. England [KS81]. English [KTKD82, KT89, MBC85, Sco85]. Engraulis [Bro85]. Enrichment [PFT^{+}85]. Entrainable [KLJ84]. Entrainment [BGC81, JST82]. Entry [BE82]. Enumeration [NG81]. Environment [CG84, DA82, ZR81]. Environmental [Cou85a, Cou85b, Goo85b, PH89, SB86b, SP88, Uph89, Van87a]. Enzymatic [HSG^{+}88]. Enzyme [BL86]. Epiphytic [SJ89a]. Epithelial [JH83]. Epizootiology [Rat81]. Epoch [PH89]. Equations [KLJ84]. Erie [BP87, Gof86, HM87a, HWD88, KMC84, RC82]. Errata [Ano81f, Ano84g, Ano85g, Ano86f, Ano86g, Ano87e, Ano87f]. Erratum [Ano84h]. Eruption [BNW88]. Erythrocytes [TC89]. Erythromycin [MS88b]. Escanaba [SK81, Ser82a, Ser86]. Escape [ED86]. Escapement [Han82]. Esocids [WS88]. Essential [HRN83]. Establishing [Mes88]. Establishment [RF81]. Estimate [CGT82, Car87, HDLG88, HG84, New81, Pau82, WOL83]. Estimated [Fab87a]. Estimates [Arg82, Fra81, HCG89, HH87, Res89, TCKW83]. Estimating [BGC81, EC86, GMW83, Han82, THV87]. Estimation [CH85a, FB83, HHB81a, Jen85, PMB82]. Estimator [RMC85]. Estuaries [BGC81, RTV84]. Estuarine [Dee81, FGB^{+}88, MWG82, MZM89, NS84b]. Estuary [BZP81, HHB81b, KO83, SHBM^{+}81, SKMK85, Vet82, WO86, Dey81, MWMSH85]. Etomidate [AGE82, LGP83]. Eualus [MP84]. European [VJ86]. Eustrongylides [WMRH86]. Eutrophication [HM87a]. Evacuation [SWM87]. Evaluate [HMMJ81b, JSC88, MNT83]. Evaluating [RCB85]. Evaluation [ATH86, BM87, BS81, EH83, HMJ87b, HJ85, MB83, OM82, PN87, PCW83, RC82]. Evaporation [Boy85b]. Even [GSMH88]. Even-Year [GSMH88]. Events [Bal84, DT87]. Evidence [CLS87, CJ85, HDGE82]. Evoke [BSS86, MS89]. Evolution [Ney81]. Examination [Hil88]. Exchange [MB83, WO86]. Exclusion [SJ81]. Exotic [LM85a]. Expansion [And87]. Experience [DF87, Hud84, TSC84]. Experimental [AMCD85, BOY84, Boy82, CW88, GSC81, HO83, HDGE82, MSC86, YMH86]. Experiments [Fle85, Jen81]. Explicit [Jen84]. Exploitation [Bow89, FS85, Kwa81, RHU82, SK81]. Exploited [HG85]. Exploration [RCH87]. Exploratory [KT89, OT86]. Exposed [Bod85, FF86, HA86, JH83, MJL89, MBRK81, RJ85, SW88b, WBH85]. Exposure [BQW^{+}86, BB84, BJCM87, RSE87, SH86]. Exposures [SSGT84]. Expulsion [CH85b, MS86, SM84]. Extended [Pec82]. Extension [NN82]. External [Ker82]. Extracellular [TGRA82b, TGRA82a]. Eye [Fab87a]. Eyes [KC89].

Factor [HN84, McG85, WJ87]. Factors [EWE^{+}81, KO83, MMI88, Min89, RMC85, Ser82a, SB83, SP88, Tan85]. Failure [KBDP85]. Fall [DPR89, GS82a]. Fall-Run [DPR89]. Fall-Winter [GS82a]. Families [BS82, YMH86]. Farming [Sti89]. Fat [MR83, STP82, WM84, WJ87]. Fathead [AQM^{+}82, BMHW81, DBBF83, DGS^{+}85, GB82, MMB81, TRP83, TRM^{+}86, WRH^{+}85]. Fatty [SMLM83]. Fear [HHB81b]. Fecundity [DPR89, GB82, Gal83, GD85a, Gal86, Nak87, RB86]. Fed [IN82]. Feed [Mil82]. Feeding [AB83, BL86, Bra86, BJCM87, CA83, CB83, DMO82, EB86, EWE^{+}81, Gar81, Ger84, HO83, IZ83, IN82, JV87, KT85, LO83, Lue86, MM85, ML82, MRTR82, MBC85, MD86, MW88, NS84a, PWM81a, Rei83, RW81a, SWM87, Sco85, SD83, SN84, STP82]. Female [DT84a, HDGE82, SJ89b, MPP82, STS87, YMH83]. Feminization [GRSD83]. Fertilization [Boy81a, CPML85]. Fertilizers [Boy81b, HB81]. Field [CZCP84, EMC^{+}85, FF86, GCCL81, Kru88, MJE89, ML82]. Filter [DMO82, JV87, MD86]. Filter-Feeding [DMO82, JV87, MD86]. Filtering [MD86]. Fin [Bea81, Sta88]. Fin-Ray [Bea81]. Final [SS82b]. Fine [CCMD81, PTMG89]. Fines [Cha88]. Fingerprints [FJT89]. Finned [Daw88]. First [BL86, Fra81, INKP87, Lem85, MM85, NG86, Ser82b, Ser84, TS84, WJ87]. First-Order [Lem85]. First-Summer [Ser84]. First-Year [INKP87, NG86, Ser82b, WJ87]. Fish [AGE82, ACG^{+}82, AK84, BA83, Bay81, BLD89, BLR81, BC84, Boy81a, Boy81b, Boy82, CGT82, CE83, CEG88, Cob82, Cou81, Cra84, Dab89, DS85, DP86, DP87, EB86, Edi89, Erm86, EC86, FKY84, FF86, Fis87, Fle85, FCL^{+}83b, GC83, Ger84, GCCL81, GD85b, GC81, GC84, GA83, Gri81, HMMJ81b, HM86, Han82, HFKR81, Het84, HG84, Hil88, HJ85, Hon84, HG87, HB81, IM83, JO82, JST82, Joh89, Kav84, KG83, KS81, KN81, Kus81, LJL86, Lem85, LM86, LMW81, LGP83, MJE89, Mar83, Mar85, Mar86, MMNW81, Mat81, MHS85, MO88, MBM89, Min89, MD86, MS84b, New81, Ney81, NS84a, NS84b, OIWD81, OT86, PN87, PMB82, QB85, RW86, Rah81, RTV84, RW81b, SR82, SJ81, SSS86, Sny84, SHM83, TAS86, Tay84, TCKW83, Tyl88, Wet88]. Fish [WBD87, ZR81]. Fish-Nursery [RTV84]. Fish-Sampling [BLD89]. Fish-Stock [BA83]. Fished [GC81]. Fisheries [And87, BM83b, BP87, Bro87, EH86, Fed87, Gal87, Hoe87, Hus87, Ker82, Lof87, MCH87, MRG87, Noa88, PG87, Rig82, TDG^{+}87, Van87b]. Fishery [BBW87, CM83, Con89, DF87, ED88a, ED88b, GM87, Gre87, JST82, Koh81, Mar87b, Ret87, TMF83, Ush87b, WH82a, WH84, Yar87]. Fishery-Landing [CM83]. Fishes [Ald85, Bal84, BVBM87, BP85, Car82, DKT88, EE89, FG83, FD89, FHT84, Gib84, HDLG88, Har87, HS83, Hol86, KWDS81, Kru88, LM87, Man84, Mar85, MSPS81, MWG82, Mer86, MB85, NP83, Noa88, Nob81, OM82, PB83, RG86, SK85, SZS82, Smi89, SK84, SKC81, SP87, Tho86, WW86a, WMRH86, WT88]. Fishing [Bed87, Edw89, Goo85b, HWD88, PM83, PS81a]. Fissures [KC89]. Fitting [CH85a]. Five [Boy81a, MBM86, PEMM87]. Fixed [KWDS81]. Flannelmouth [BMU87]. Flathead [DCR89]. flavescens [Hin85]. Flavor [BB82]. Flavors [LLBA86]. Flexibacter [JP81, MBM86]. Flooded [BH88]. Flooding [Har87]. Floodplain [FK88]. Florida [ML84, CMNS85, ED88a, ED88b, FLK^{+}87, GLC82, HSM85, INKP87, Joh83a, KWP88, LNV83, MMI88, MM88a, MW86, PCW83, Res89, SJ89a, WJ87, WC86]. Flounder [Bow89, Cam84, KMCB84, Pow82, Pre89]. Flow [AS83, BSDD86, CWW^{+}86, EMC^{+}85, Heg88, MB85, TC89]. Flow-Cytometric [BSDD86]. Flowing [PAH89]. Flows [GD85b, OM82]. Fluctuating [AMP82, STP82, TGC^{+}86, ZR81]. Fluctuations [Fui86, Heg88]. Flumes [CW88]. Fluorescent [PT88, WBD87]. Focusing [Fab87a]. following [OS86]. fontinalis [KR85]. Food [BZP81, CLS87, Car87, DP85, EWE^{+}81, GV83, HM87a, HCG89, HHB81b, MWMSH85, Mes88, MF81, PS81b, RMC85, RM87b, SWM87, SJ89a, SA82, SB86c, Str82, WGR87, WRH^{+}85, YMH83]. Food-Chain [WRH^{+}85]. Foods [LD86]. Forage [KS81, Ney81, Nob81, SFW83, SKC81, SP87, WB81]. Forage-Fish [KS81, Ney81]. Foraging [Bac84, Rug86]. Forecasts [Edw89]. Forestry [Gal87]. Forgotten [BM83b]. Fork [PTMG89]. Form [BSN88]. Formation [Pow82, Rad84, Sch89, TT82]. Formats [Hoe87]. Formulation [Dab89]. Four [BNF82, BS81, BSZF83, DP85, EH83, KG83, LNV83, MHS88, OGS^{+}83, RW81b, Sta88]. Fourbeard [Her85]. Fourteen [DP86]. Fraction [DGS^{+}85]. Fractional [GB82, Gal83, GD85a, Gal86]. Framework [TL87]. Fraser [MR89]. Free [Bac84, BB84]. Free-Ranging [Bac84]. Frenchman [Gar83]. Frequencies [BSP87]. Frequency [BA83, CGT82, GB82, Gal83, GD85a, HDS86, HCG89, WCM88]. Fresh [Cou85a, Eri84, Joh83b, MBRK81, PHD85]. Freshwater [BCNW87, BL83a, HSR^{+}86, JM82, Jon85, MSPS81, Mat84, RTV84, SB86b]. Fry [BS82, CD83, HW81, MBRK81, MD81, WF81]. Fuels [MBB^{+}84]. Function [And87, BW89]. Functional [HZ89, HFDL83, WW86a]. Functions [CC81]. Fundamental [Ryd82]. Fundy [RM87b]. Furunculosis [CH86]. Fusion [DEAT87]. Future [Cun89, Tis89].

Gadus [DRV85, KT85, ØKF85]. gairdneri [RJ85]. galilaea [DHVG87]. Galveston [MHLL87]. Gambusia [EE89]. Gammarus [MMY^{+}82]. Gas [CW82, Col84, Hau86]. Gases [Bou82]. Gasometer [Bou82]. Gavins [RHU82]. Gear [BLD89, DP87, TCKW83]. Gears [GC83]. Gene [DEAT87]. Genesis [DC82]. Genetic [AV84, BMW89, Bro86, CJ85, CWSM86, EE89, Edi89, GSMH88, HR87, HG89, KMTU84, KM87a, KMKM89, LB82, LAK85, LAPK87, MPP82, MNT83, PCW83, RNT85, SHI84, SSU86, WCSM86, WB85]. Genetics [AM89, BWMB88, FBCV82, HSR^{+}86, SLJ^{+}84]. Genotype [WD88]. Geographic [BC84]. Geomorphology [LHW87]. Georges [Her85, OT86]. Georgia [BHR85, HG85, HBB84, PM83, RTV84, TSD81]. Geosmin [LLBA86]. Geothermally [FBCV82]. Gill [BE82, DHVG87, ED88a, ED88b, HCFE81, HM85, JH83, Joh83b, MJL89, SBJD83, WOL83]. Gill-Net [ED88b]. Gizzard [DSM85, DMO82, HWGF82, MW88, PV88, PWM81a, PWM81b, SWM87, Sto86, VP88]. Glacial [BSS^{+}88, LF89]. Gland [Kav84]. Glochidia [Mol83]. Glycine [GA83]. Glycogen [VHH86]. Goldfish [HSG^{+}88, SN84]. Gonad [WLC82, ZB84]. Gonadal [DM81]. Gonadotropin [Hon84, ZB84]. Gonads [WM84]. Gossypol [DRRP82]. Grade [Tho86]. Grand [KZ83]. Granular [Boy81b]. Grass [AS83, ATH86, BB83, BSYW84, HKA83, JSYW83, LVNM82, LNV83, MPP82, PAH89, TC89, WW86b, YMH83]. Grayling [BSM82, HG89, Kay89, Sco85]. Grazing [DTLK84]. Great [BBW87, Bro87, DF87, Esh87, FW86, Fed87, HWD88, Mer86, MRG87, WS82]. Green [Ger84, Lem85, MJ83, SJ82, LB82, Tyu87]. Greenback [SB86a]. Ground [Tan85]. Grounds [MJE89]. Groundwater [SP85]. Group [Pra88]. Grouper [MH84, ML84]. Groups [HA86]. Growing [BH86]. Growth [AAC88, ACG^{+}82, Bar85, BM83a, Bow89, BH86, CB88a, CB89, CLSRS89, CD83, CC81, CCMD81, DSM85, Daw88, DP85, DT87, Dey81, Dia83, DA82, DRRP82, DEAT87, Fab87b, FB83, GS82a, GA83, GT85, GMB^{+}87, HDS86, HE84, HG85, HR87, Hin85, HA83b, HSM85, INKP87, Jen84, Joh83a, KE85, KG83, KBH85, KT85, KTKD82, KR85, LHA84, LD86, LO83, MHLL87, MH84, MRTR82, MS82, MS84a, Mil82, Mil83, MF81, MPSC89, MBRK81, MR83, Mol83, NG86, NM82, NS84a, PLC86, PMR86, Pow82, Rec86, Res89, RH82, RA86, Ros88, Sco88, Sco85, Ser82a, Ser82b, SBE84, SGK^{+}89, SJ89b, SFS83, STP82, TCS88, TSD81, TJ83, TY89, WD88, War84, WM86, WS83, WH82b, WLC82, WC83, YMH83, ZS88]. Growth-Invariant [Fab87b]. Gulf [Bar85, CFCSI84, CLSRS89, FCBTSR83, KG89, NM82, BP84, DT87, Mei84, RPI89, RTP85, SWT^{+}85, WM84]. Gulls [Rug86]. Gynogenetic [JSYW83]. Gyrodactylus [Kir81].

Habitat [AK84, BOY84, BSN88, BR86, Bra86, BSS^{+}88, CHW87, CB88b, Cou87, DB86, Dol87, DMH84, FHT84, HJ83, Heg88, HGP87, JBL88, LHW87, LJL86, LM85b, LF89, MJE89, MHEG89, MG88, PN87, PTMG89, PFT^{+}85, Ran86, SB86a, SCSR86, SH84, TR89, Tyu87, VP88]. Habitats [Fis87, GD85b, HHN87, SCD81]. Habits [BZP81, HHB81b, Lue86, ML82]. Haddock [KBDP85]. Hake [Gar83, LO83]. Hakes [Her85, RA86]. Handbook [Sti89]. Handling [Pra88, SW88b, WM84, WC86]. harengus [DRV85, GT85]. Harvest [BA85, JH82, KBH85, MMY^{+}82]. Hatcheries [BHL^{+}85]. Hatchery [Bac84, BMU87, CLL86, EE89, FMS81, Han82, LAK85, MRTR82, MBM89, SW88a, Sos82, WS87]. Hatchery-Reared [MBM89, WS87]. Hatching [RW83]. Hauling [CTSD84a]. Head [SK86, TK85]. Headwater [SB86a]. Healing [NM88]. Health [Gri81]. Heat [TJS81]. Heated [FBCV82, GCCL81, PM81]. Heavy [HR88]. Hecate [FC87]. Held [CW88]. Helens [BNW88]. Helicopter [Hau86]. Help [Ret87]. Hematological [SSGT84]. Hematology [Hic82]. Hemiuroid [MM88b]. Herbicides [FF86]. Herbivorous [Ger84]. Hermaphroditism [EG86, MWSR83]. Herring [CB83, DRV85, GT85, Hay86, KP87, McG85, OK86, TK85]. Heterogamety [HDGE82]. Heterogeneity [Kre86b, Str82]. Heterotrophic [CPML85]. Heterozygosity [ZR81]. High [CHW87, GB86, LM82, MYA81]. High-Elevation [CHW87]. Highly [HGP87]. Hind [MH84]. Histological [MS88a]. Histopathology [BHL^{+}85, VO87]. Historical [HS89]. History [Ald85, BP87, CFCSI84, Dab89, DT87, HC84, Jon85, KZ83, KHS81, KBDP85, Kru88, KR85, Mar85, MO88, NMC85, Smi89, Tro89, VJ86, VAKT87, WW86a]. Homarus [Enn84a, Enn84b, Enn86]. Home [BHR85, BWQ84, FS83, MW86, SFW83]. Home-Range [SFW83]. Homing [BQW^{+}86, MW86, WQB82]. Homozygous [CB89]. Hormone [DSP84, DEAT87, Hon84, SBJD83]. Hormones [GRSD83]. hoyi [RCB85]. Hudson [Dey81, Fab87a, MCHL81]. Human [Bro87, DEAT87]. Humidity [NB85]. Humpback [KZ83]. Huron [Arg82, FS85, KT87, Rec86]. Hybrid [AS83, AV84, BH86, KWP88, MM88a, Mac89, MPP82, WW86b, WC86, YMH83]. Hybridization [BMU87, FJT89, FH89, MMI88, Pin82]. Hybrids [BB83, FLK^{+}87, HD87, HD88]. Hydraulics [BSN88]. Hydrazine [HFDL83]. Hydrocarbons [Mer86]. Hydroelectric [GW83, SK86, TK85]. Hydrology [Boy82]. hydrophila [HER81, TGRA82b, TGRA82a]. Hydrostatic [Gib84]. Hyperglycemia [MSC86]. Hypochlorite [MTB81]. Hypochlorous [MTB81]. Hypolimnetic [Gri81]. Hypothesis [Cou85b, PFT^{+}85, SWT^{+}85]. Hypoxia [SP88].

Icelandic [SÍW89]. Ichthyophthirius [WT88]. Ichthyoplankton [FCBTSR83, GC83, Hol86]. Idaho [HGP87, PTMG89]. Identification [BSM82, CWSM86, FCL^{+}83b, HD88, KM87b, Kru88, LAPK87, Mar84, Pra88, RC82, SW88a, Sny83]. Identify [HDLG88]. Illinois [AB83, FS83, Sto86]. Illumination [Hin85]. Illustrate [FG83]. Immersion [Het84]. Impact [DTLK84, Mar87b, MF83, PFT^{+}85, PG87, SWRM85, Van87b]. Impacts [HS83]. Impairment [BJCM87]. Impingement [JST82]. Implantation [JSYW83]. Implanted [MS86, SM84]. Implants [MS88a]. Implications [CCP87, FW86, MYA81, MRG87]. Importance [Bea81, BZP81, DS85, DA82, GA85, HM86, MWMSH85, MENG86, Sto86]. Impoundment [CVC85]. Impoundments [Nob81]. Imsa [VJ86]. In-state [Edw89]. Inadequate [Lof87]. Incidence [MYA81]. Incisions [NM88]. Include [And87]. Incorporates [ZS88]. Increase [MHS88]. Increased [Dee81, PTMG89, ZR81]. Increasing [AR82, BW89]. Incremental [OM82]. Increments [DSM85, EC86, TY89]. Incubated [Car83]. Incubation [CEG88]. Independence [Eva84]. Independent [DT84b]. Indeterminate [Gal86]. Index [FKY84, GA83, JO82, JST^{+}81, KYFS87, LO86, RMC85, Ryd82, Str82, YH82]. Indexes [Wal81]. Indicated [WCM88]. Indication [Ser86]. Indicator [HMJ87b, LAK85]. Indicators [VHH86]. Indices [Con89, EH86, Kre86b, PC85, SR82]. Individual [RCH87]. Induced [CTSD84b, FK88, TJS81]. Induction [SS82b, WLC81]. Industrialized [BSP87]. Inexpensive [Bou82]. Infection [BSZF83, HHEM86, WMRH86]. Infectious [MO86]. Inferred [NG86]. Inflammatory [MS88a]. Inflows [Mat84]. Influence [BVBM87, Bar82, DGK^{+}88, Fos85, KE85, Mat84, MHEG89, MB83, MCR84, MCK86, MR82, PCC85, RS83, Rei83, RTV84, SJ82, Ser82a, SL86, SGK^{+}89, SS84, Str82, WD88, WQB82]. Influenced [SS82a]. Influences [CVC85, DSP84, FP89, SÍW89, TCKW83]. Influencing [RMC85]. Information [Bro87, Ken81]. Ingested [HDLG88, TB87]. Ingestion [DHVG87, MD86, SJ82]. Inhabiting [ZR81]. Inherited [RM87a]. Inhibitor [GA83]. Initial [RW81a]. Injected [Smi84]. Inland [EH86]. Inlet [Kay89]. Inorganic [Boy81b, PHD85, PMR86]. Input [Hus87]. Input-Output [Hus87]. Inshore [LF83, SZS82, Tan85]. Instability [LAK85]. Institutional [And87]. Instream [MB85, OM82]. Intake [Fle85, YMH83]. Intakes [SJ81]. Integrity [FKY84, KYFS87, LO86]. Intensity [CW88, LKF82]. Intensive [SFS83, WF81]. Interaction [DT84a, EH86, SS82a]. Interactions [CB88a]. Interactive [CLH^{+}86, Mil83]. Interannual [FGB^{+}88]. Intercepts [Car82]. Interdependence [BHN84]. Interdigitation [HD87, OGS^{+}83]. Intergrade [INKP87]. Internal [TC89, VO87]. International [Cun89, Tis89, Mar85]. Interpretation [HH82c]. Interrelationships [MPSC89]. Intersample [HCG89]. Interstice [JBL88]. Intertidal [Gib84, Mar86]. Intestinal [LL82]. Intraspecific [MM88a]. Introduction [CM81, TDG^{+}87]. Introductions [LM81]. Introgressive [BMU87]. Inundated [VP88]. Invariant [Fab87b]. Invertebrates [MO88]. Ions [MTB81]. Island [Fab87a, SB86b]. Isoelectric [Fab87a]. Isozyme [HDS86]. Issue [Lof87]. Issues [Bed87].

Jack [CP83]. Japan [SK85]. Jersey [Mil81, TAS86, VAKT87]. Joaquin [KO83, SKMK85]. Jolly [HG84]. Jolly-Seber [HG84]. Juvenile [BSS86, BS87, BSN88, BZP81, BHL^{+}85, BE82, CB88a, CZCP84, CC81, CCMD81, Dee81, DB86, Dol86, Dol87, EH83, Eri84, FV82, FGB^{+}88, Gar83, GW83, Har86, HD88, HGP87, HH82b, IN82, KLM86, LKF82, LO83, Lue86, MSBB88, MJL89, Mil81, MM88b, MR89, OK86, Ors87, PT88, PS81b, Rug86, SZ85, SK85, SW88b, Sos82, SS84, Tan85, TK85, Tay88, TGC^{+}86, WO86, WBD87, WC83, ZG89]. Juveniles [FW86, HM87b, LD86, ML82, MR83, ØKF85, RW83, Ser84, SHBM^{+}81, Uph89, WM86].

Kansas [Bro86]. Kaplan [SK86, TK85]. Keep [RC87]. Kelp [MP84]. Kelsey [SSS86]. Kenalog [CRF83]. Kepone [BMHW81, MMB81]. Keys [ML84]. Killifish [Bro86, Mei84, WM84]. King [HMSK86]. Kingfish [SW85]. kisutch [TM85a, TM85b]. Kokanee [MMR89]. Koyukuk [HG89].

Labeling [TY89]. Laboratory [Bro85, CZCP84, CCMD81, EWE^{+}81, FW86, FF86, GCCL81, Kru88, LHA84, ML82, MW88, Rah81, RW81a, Tay88]. Laboratory-Reared [EWE^{+}81]. Labrador [PLC86]. labrax [SSGT84]. Lack [Mar86, TCS88]. LaCloche [HPG83]. lacustris [MMY^{+}82]. Lag [Car86]. Lagoon [CFCSI84, FCBTSR83]. Lake [BSP87, BO89, BP87, CP83, DPR89, Esh87, FS83, FD85, Fos85, FP89, FS85, Gof86, GMB^{+}87, HW81, HJ83, HRN83, IM83, IN82, Kay89, KMC84, KT87, KS86, KMKM89, Kwa81, LA84, LM87, MMI88, MF81, MF83, MMR89, OWNH88, Pec82, PZI89, PWM81a, RW86, Rec86, RF81, RC82, SLJ^{+}84, SK81, Ser82a, Ser82b, Ser84, Ser86, SJR87, Sto86, TB81, TJ83, Ush87b, Arg82, Bra86, BMM^{+}87, Cou81, CC84, FL81, FG89, HHB81a, HM87a, HFKR81, HS89, HWD88, JST82, Joh89, JTDM87, KM87a, KM87b, KMKM89, LF83, LB82, MJM^{+}83, Man84, Mar87a, Mar87b, MHS85, OS86, RCH87, SW88a, Sny83, SKC81, SB86c, SHCMS83, WH82a, WH84, WH82b, WT88]. Lake-Fed [IN82]. Lake-Resident [FP89]. Lakes [AW85, BBW87, Bro87, CMNS85, Dia83, DF87, DA82, Esh87, FW86, Fed87, GC81, Har86, HWD88, JH82, LNV83, LMW81, Mer86, MW86, Min89, MRG87, Ney81, SR82, SJ89a, TMF83, WS82, WH82b, WB81]. Lamellar [HM85]. Lampreys [BM88]. Landing [CM83]. Landings [Goo85a]. Landlocked [NN82, RNT85]. Lands [BH88]. Large [Cha88, GC83, GD85b, IN82, KWDS81, MHEG89, Ors87, RF81, SCD81]. Largemouth [AMP82, BJCM87, BH86, CMNS85, Car87, CTSD84a, CTSD84b, CWSM86, CL82, CD83, FLK^{+}87, FS83, GA85, HSR^{+}86, HER81, HCG89, HO83, HSM85, INKP87, JBL88, KW89, KE85, KWP88, MMI88, MM88a, Mac89, MW81, MJL89, MK89, MW86, MS82, MS84a, PCW83, SS82a, SJ82, SFW83, SZ85, SA82, SCH81, Sto86, TT82, TSD81, WJ87, WC86, WCSM86, WF81]. Larva [BL86, Bor83, CP86, SJ81]. Larvae [Bod85, BC84, BMM^{+}87, Bre88, Bro85, CLSRS89, CB83, EB86, EWE^{+}81, FCL^{+}83b, GC83, GDH83, HPHF85, Hay86, HFDL83, Her85, Het84, HA83b, HPG83, JB87, KMC87, KP87, KMCB84, KT85, LHA84, MM85, Mar86, Mar84, ML82, McG85, MRC81, MRN83, MS84b, NJS86, NS84b, OGS^{+}83, PC85, RCH87, RVV85, RW81a, SHBM^{+}81, Sny83, TY89, Uph89, WOR85]. Larval [BH83, CE83, Dab89, EB86, EC86, FG83, FB88, FHT84, Fui86, GT85, HD87, Hin85, HS83, KM84, KG89, LM82, MJM^{+}83, Mar85, MWMSH85, Mat84, MBC85, PV88, Qua85, Rad84, RCH87, SK85, SWT^{+}85, Sny84, SHM83, Tan85, TB81, TCKW83, TJ83, VP88, WBD87]. Larval-Fish [TCKW83]. Latitudes [MW81]. Lawrence [RPI89]. Leaching [WRH^{+}85]. Legal [AR82]. Lemon [WGR87]. Length [Daw88, HH87, MM88a, OK86, WCM88]. Length-Frequency [WCM88]. Lengths [Car82, HDLG88]. Lens [Fab87a]. Leon [EE89]. Leopard [Smi84]. Lesions [SRM83]. Lessons [DF87]. Leuciscus [MBC85]. Leucocytes [EMC^{+}85]. Level [DP85]. Levels [MMNW81, MPAR86]. Life [Ald85, Bal84, BL83a, CFCSI84, CLH^{+}86, CP83, Dab89, DT87, Fab88, HC84, Jon85, KZ83, KHS81, KBDP85, KR85, Lac89, Mar85, MO88, NMC85, NN82, Smi89, Tro89, VJ86, VAKT87, WW86a]. Life-History [VJ86]. Life-Support [Lac89]. Light [CW88, KC89, LKF82, MCK86]. Lightly [HG85]. Likelihood [ATH86]. Limit [Wre84]. Limitation [CLS87]. Limitations [GM87]. Limits [BLR81, SJR87]. Limnology [Rig82]. Linear [HH87, RMC85, Str82]. Lipid [GS82a, Mil83]. Lipids [DM81]. Liquid [DBBF83, DGS^{+}85]. Little [KZ83, War84]. Littoral [LM87]. Littoral-Zone [LM87]. Load [JP81]. Lobster [AR82, CM83, CWCR89, Enn84a, Enn84b, Enn86, FI88, GLC82, Mac82]. Lobsters [RC87]. Local [LF83]. Localized [Wet88]. Locomotor [SN84, WW86a]. Locus [CB89, ZR81]. Log [HH87]. Log-Linear [HH87]. Logging [Ell86]. Logistic [Jen84]. Logs [MENG86]. Long [Erm86, Rec86, Gof86]. Long-Term [Erm86, Rec86]. Longevity [Sco88]. Longitudinal [HG87]. Longlines [GAT82]. Loss [Enn86]. Losses [BGC81, HB81]. Louisiana [SWT^{+}85]. Low [CWSM86, JB87, PMR82, Pre89, SK86, TK85, WSL86, HPG83]. Low-Head [SK86, TK85]. Low-pH [HPG83]. Low-Risk [CWSM86]. Low-Temperature [Pre89]. Lower [BC84, FK88, HWGF82, Mar83, ML84, MR89, SWCB83]. Luminescence [Fab88]. Lunar [Tay84]. Lures [Ors87]. Lymphocystis [MSB88]. Lymphocytes [FCL^{+}83a].

M13 [FJT89]. Mackerel [ED88a, ED88b, PH89]. Macroalgal [HN84]. Macrobenthos [Ell86]. Macrobrachium [CPML85]. Macroinvertebrates [SJ89a]. Macrozooplankton [RF81]. Maine [AR82, Gar83, GT85, MENG86]. Maintenance [Ger84]. Major [Bed87, JTDM87, PMR86, Tan85]. Malachite [MJ83]. Male [BH83, Gof86, HDGE82, Res89, SJ89b, MPP82, STS87, YMH83]. Management [And87, AW85, Bay81, BP87, Bro87, CM81, CP81, DF87, Hil89, Hoe87, KS81, Lof87, MRG87, Ney81, Nob81, PG87, Rig82, SSU86, Tis89, TMF83, Yar87]. Managers [Ret87]. Manipulation [MG88]. Manual [Kru88]. Manure [CPML85]. maraena [BSM82]. Marine [Edw89, Gib84, Het84, MM88b, Tyl88]. Maritimes [CM83]. Mark [Jen81, MS89]. Marked [WO86]. Marker [CWSM86, WBD87]. Markers [WCSM86]. Market [Fed87, Hig87]. Marking [EH83, Han82, Het84, MD81, PT88]. Marks [SSU86]. Marsh [RTV84, TAS86]. Marshes [TAS86]. Maryland [GCC85, Goo85a, Uph89]. Mass [OS86, ØKF85, TCS88, WT88]. Materials [Goo85b]. Matter [QB85]. Maturation [Dia83, SS82b]. Maturity [Bow89, NG86]. Maxima [FLK^{+}87]. maximus [Qua85]. Mdh [ZR81]. Mdh-B [ZR81]. Meadows [HKA83]. Meal [DRRP82]. Means [PD88]. Measure [BA83, MPAR86]. Measurement [HMJ87b]. Measurements [Car82]. Measures [Gre87, Smi85]. Measuring [BLR81, DHSA83, HMJ87a, Tis89]. Mechanical [BHL^{+}85, HMMJ81b, HMMJ81a]. Mechanisms [Dia87, KT89, MS86, RCH87]. Median [CH85a]. MEI [Ogl82]. Melanogrammus [KBDP85]. Menhaden [DT87, PS81b, SWT^{+}85]. Menippe [SH86]. mercedis [KO83]. mercenaria [SH86]. Meristic [RPI89]. Meristics [HD88]. Mesh [SJ81]. Mesosaline [HSR^{+}86]. Metabolic [BS87, Vet82, VHH86]. Metabolism [Eva84, PWM81b]. Metabolites [LB84]. Metal [HM86, RB86]. Metal-Contaminated [RB86]. Metallothionein [DEAT87, HM86, HMJ87a, HMJ87b]. Metallothionein-Human [DEAT87]. Metals [HR88, Joh89, MBM86]. Methemoglobin [ST83]. Method [Bea81, Car87, DHSA83, GMW83, HMMJ81b, HJ85, Mar87b, Pau82, Pra88, SD87]. Methodologies [MS89]. Methodology [BGC81, OM82]. Methods [ACG^{+}82, Cam84, EC86, Fab87b, FD89, Jen85, PT88, Res89]. Methotrexate [CRF83]. Methyltestosterone [JSYW83]. Mexico [Bar85, CFCSI84, CLSRS89, FCBTSR83, NM82, RTP85]. Michigan [CC84, Dia83, FD85, HHB81a, HS89, JST82, JTDM87, LB82, MJM^{+}83, Man84, RCH87, SW88a, SKC81, SB86c, TJ83, WS82]. Microbranchiospines [DHVG87]. Microdistribution [GC83]. Microhabitat [BVBM87, MB85]. Microhabitats [SD83, Tay88]. Microscopy [EMC^{+}85]. Midwater [Arg82]. Midwestern [JH82, KYFS87]. Migrant [Jon85]. Migrates [Koh81]. Migrating [CW88, Jes87, OK86, Rug86]. Migration [FG89, HCFE81, SB86b, Tan85, WCM88]. Migrations [Lei85, MW86]. Mile [CB88b]. Milkfish [TY89]. Minimum [AR82]. Minnesota [RW81b, Wat83]. Minnow [BHN84, BMHW81, GB82]. Minnows [AQM^{+}82, DBBF83, DGS^{+}85, Gal83, GSC81, MMB81, TRP83, TRM^{+}86, WRH^{+}85]. Minor [BH88]. Mirex [BMHW81, MMB81]. Mississippi [BC84, FK88, HWGF82, HBD88, HS83, Hol86, HHN87, PCC85, SH84]. Missouri [DCR89, JST^{+}81, MMNW81, RHU82, WH82a, WH84]. Mitogenic [FCL^{+}83a]. Mixtures [FV82]. Mobile [Mar83]. Model [BST89, CH85a, Dab89, FI88, HG84, HH87, Jen84, OT86, PN87, WDS85]. Modeling [Car87, DRH88, JSC88, SB86c]. Modelling [SWRM85]. Models [GC84, JST82, MH84, Ret87]. Modification [MHA81]. Mohave [Mar87a]. Molinate [FF86]. Monitored [GGS82]. Monitoring [Bou82]. Monkeyface [HN84]. morar [MM85]. mordax [AMCD85]. morhua [DRV85, KT85, ØKF85]. Morone [Bod85, HD87, HD88, OGS^{+}83, Wal86a]. Morphoedaphic [JO82, Ryd82, SR82, YH82]. Morphogenesis [KW89]. Morphological [BB83]. Morphology [Bea84, BWMB88, BMW89, HFDL83, JH83, Mac89, MJL89, WS88, WW86a]. Morphometric [Fab87b, PC85, RPI89]. Morphometrics [HD88]. Morphometry [HD87, McG85, SJ81]. Morphs [GV83]. Mortalities [BO89]. Mortality [BP84, Dey81, EC86, FK88, FB83, Fra81, GW83, GT85, HPHF85, HG85, HG84, HH82b, HPG83, Jen84, Jen85, JM82, KLJ84, LN87, LM82, MSC86, MH84, Mat85, MZM89, NG86, NM82, OS86, PMB82, RH82, Ros88, SK86, TK85, TGC^{+}86, TSD81, War84, WT88]. Mosquito [TAS86]. Mosquitofish [Ang83, MTB81, WC83]. mossambica [Bow81]. Motility [Dup82]. Mount [BNW88]. Mountain [DA82, HPG83, LHW87, SB86a]. Mountains [LM85a]. Movement [TR89]. Movements [BSS^{+}88, BK85, CB88b, DCR89, DGK^{+}88, Enn84a, HSB^{+}83, HHN87, KT87, LF83, MCHL81, MK89, RW81b]. Mucus [CH86]. Muddy [LLBA86]. Mullet [DM81, Rad84]. multifiliis [WT88]. Multiple [BSS86, LNK87, SW88b]. Multiple-Use [LNK87]. Multispecies [SE87]. Mummichogs [DT84a, VHH86, WMRH86]. Muscle [FD89]. Museums [Smi89]. Muskellunge [DMH84, GSC81, MSC86, SCH81, TSC84]. Mussel [Mol83]. Musty [LLBA86]. Musty-Muddy [LLBA86]. MW [SK86]. Myomeres [BC84]. Mysid [BL83a]. Mysis [BL83a, RF81].

Names [Edi88]. Naphthalene [MBRK81, MR83]. Narcosis [Fab88]. Native [BMU87, CJ85, LM85a, Lem85, Wre84]. Natural [BS82, CLL86, EE89, Fra81, FH89, Hig87, Kru88, LN87, LD86, MW88, Ran86, SR82]. Naturalized [KM87a, KM87b]. Navigation [HBD88, HS83]. near [RW81b, SJ89b]. Nearshore [MWG82]. Nebish [Ser84]. Necrosis [MO86]. Need [Con89]. Needed [Bro87]. Neglected [DRH88]. Nematoda [WMRH86]. Neomysis [KO83]. Net [ED88b, LJL86, McG85, SHI84]. Net-Pen [SHI84]. Nets [ED88a, SCD81]. Neuroendocrine [Mei84]. Neustonic [Her85]. Neutralizing [Cip83, HM87b]. Nevada [Mar87a, Sny83]. Newfoundland [Bow89, Daw88, DB86, Nak87, PLC86]. Newly [TM85a]. Nickel [GGS82]. nilotica [DP85, KW89, LL82, TCS88]. Nineteenth [Cha86]. Ninth [EB86]. Nitrite [HA83b, LM86, SSGT84, ST83, WBH85]. Nitrite-Exposed [WBH85]. Nitrogen [HN84]. Nomenclature [Edi89]. Nonessentiality [LL82]. Nonmarket [Gre87]. Nonmonetary [Gre87]. Nonnatal [MR89]. Nonnative [CJ85]. Nonpiscivorous [HCG89]. Nonpolluted [Ang83]. Nonreporting [Mat81]. Nonstunted [HR87]. Normal [CEG88]. North [FCL^{+}83b, HHB81b, Lem85, Pow82, Sco88, WCM88, CLSRS89, DRV85, Edi88, MBM89, RVV85, RPI89, RH82, Ros88, Ros89]. North-Central [CLSRS89]. Northeast [War84]. Northeastern [NMC85, PLC86]. Northern [BS82, BH86, Car87, CB88b, Dia83, Dia87, Dup82, FLK^{+}87, HSG^{+}88, HH82a, INKP87, KWP88, MMI88, MM88a, PLC86, PCW83, WC86, Bar85, FP89, LM87, NM82, RW81b, TMF83]. Northwestern [GSMH88]. Norway [Jon85, VJ86]. Norwegian [HJ83, NJS86, ØKF85]. Notes [AAC88, CEG88, FD89, FH89, GC83, Mac89, MSB88, MHS88, Sta88]. Nova [Jes87]. Nuclear [PM81]. Nucleic [FD89]. Nucleotides [VHH86]. Number [Car83]. Numbers [BC84, Mes88]. Nurseries [MZM89]. Nursery [HHB81b, RVV85, RTV84, Tan85, VP88, WCM88]. Nutrient [PFT^{+}85, SP87]. Nutritional [MWMSH85, MMY^{+}82, PC85].

Observation [GAT82]. Observations [BSS^{+}88, FF86, Gal87, Gar83, Ogl82, WGR87]. Occurrence [Kir81, MM88b]. Occurring [SK85]. Ocean [NM82]. Oceanic [NS84b, SÍW89]. Odds [DP87]. Odor [BSM82]. Odors [BWQ84]. Off [BB82, Bow89, HG85, KN81, Mil81, NMC85, PLC86, War84]. Off-Flavor [BB82]. Offshore [FI88]. Ohio [HWGF82, HWD88, PWM81a]. Oil [BQW^{+}86, MMW85, MR83, WMM81, WRH^{+}85]. Oklahoma [MHS85]. Olfactory [HZ89, RJS85]. Omnivorous [DTLK84]. Oncorhynchus [TM85a, TM85b]. Oneida [HW81, MF81, MF83]. Ontario [Bed87, BO89, BR86, Bra86, BMM^{+}87, CP83, CP86, HPG83, KS86, KMKM89, Mar87b, Min89, OS86, OWNH88, SHCMS83, THV87, Ush87b, WM83].

Ontogenetic [Bra86, HS89, KW89, Lue86, Mab89]. Ontogeny [TM85b]. Oocytes [SS82b]. Open [SCD81]. Open-Water [SCD81]. Opeongo [Joh89]. Opioid [Kav84]. Optimum [KG83]. Oral [GRSD83]. Orally [MS88b]. Order [Lem85, Mab89]. Oregon [HG87, MWSR83, Rat81]. Organ [DT84b, HZ89, KW89]. Organ-Pancreas [DT84b]. Organic [JP81, PHD85, QB85]. Organisms [MBB^{+}84, WRH^{+}85]. Organochlorines [WOR85]. Organs [DS85, EB86]. Origin [MBM89, Sos82]. Original [HDLG88]. Oscillations [Mei84]. Osmerus [AMCD85]. Osmolality [MJL89]. Osmoregulation [RS83]. Osmotic [CMNS85]. Ossberger [GW83]. Other [MHS85, MJ83, Sco88, WMRH86]. Otolith [Bar85, Mac89, MS84a, RCB85]. Otoliths [DSM85, DT87, EC86, Het84, HSM85, MS82, NG86, Pow82, Rad84, RCB85, Sch89, TT82, Tho86, TY89, WBD87]. Otter [Nie83]. Ouananiche [RPI89]. Our [Lac89]. Outlet [Kay89]. Output [Hus87]. Ova [PMR82]. Overlap [Smi85, Wal81]. Overview [Ogl82, TL87]. Oxygen [AB82, Cob82, Cou85b, SB83, WH84]. Oxygen-Transfer [SB83]. Oxytetracycline [KLM86, TY89]. Ozark [JST^{+}81].

Pacific [Bea81, HC84, Han82, Hay86, McG85, MM88b, SRM83]. Pacific-Arctic [HC84]. Paddlefish [KMC87, MRTR82, RHU82, SH84, Wal86b]. Paddlewheel [AB82]. Pagrus [Tan85]. Palau [Mac82]. Pamlico [Pow82, WCM88]. Pamunkey [MO88]. Pancreas [DT84b]. Pancreatic [MO86]. Pandalus [PLC86]. Panulirus [GLC82, Mac82]. Paradigm [And87]. Parameter [RF88]. Parameters [ML84]. Parasite [FK88, WT88, YMH86]. Parasite-Induced [FK88]. Parasitic [HMSK86]. Parasitism [Mol83, MP84]. Parental [BH83, MPP82, WOR85]. Parr [HA83a, RJS85, RPI89]. Partial [TGRA82b]. Participation [Edw89]. Particle [DTLK84, DHVG87, JV87, MD86, SP85]. Particle-Grazing [DTLK84]. Partitioning [AMP82, FHT84, GV83, OWNH88]. Passage [CW88, WGR87]. Passed [SK86, TK85]. Passing [GW83]. Passive [HM87b]. past [MSBB88]. Pathology [HB86]. Pathways [MS86]. patronus [SWT^{+}85]. Patterns [Bow89, HD87, HS89, IN82, Jon85, Mar83, OGS^{+}83]. PCB [MMW85]. Pecos [EE89]. Peeler [BOY84]. Peelers [BOWY83]. Pelagic [Cou81, Tan85]. Pellet [GSC81]. Pellet-Reared [GSC81]. Pelvic [Sta88]. Pen [SHI84]. Penaeid [CC85]. Penaeus [HH82b, MZM89]. Pentachlorophenol [BJCM87]. People [BP87]. Peptides [Kav84]. Perca [Hin85]. Percent [ST83]. Perch [AAC88, BMM^{+}87, FS83, Gal87, HW81, HM87a, HR87, Hin85, KMC84, LB82, MJL89, MF81, MF83, MCR84, MCK86, MPSC89, MR82, MRN83, Nie83]. Perche [DCR89]. Percid [Car82]. Performance [DRFL89, Fui86, Rei83, STS87]. Performances [BP85, DSHB83]. Period [IM83, INKP87]. Periodic [LB84]. Periodicity [Bra86]. Periods [CB88b, KG89]. Peripheral [EMC^{+}85]. Periphytic [Bow81].

Permanganate [JP81]. Perspectives [Hig87]. Peru [WT88]. Peruvian [Lon88]. Petroleum [MBRK81, RTP85]. pH [ACG^{+}82, Dup82, HH82c, HPG83, JB87, KR85, PMR82, RW86, SL86, WSL86, WH82b]. Phases [BNF82, CP83]. Phenol [Ang83]. Photoperiod [DT84a, DT84b, SN84]. Photoperiod-Independent [DT84b]. Phototaxis [BM83c]. Physical [BS87, KM84, Ran86, SCSR86, SP88]. Physiological [BSS86, MSBB88, MS89, RSE87, WS87]. Physiology [PSBZ86]. Pickerel [Dup82]. Piedmont [Lem85]. Pike [BS82, Car87, CB88b, Dia83, Dia87, Dup82, HSG^{+}88, HH82a]. Pine [IM83]. Pineal [DT84b, Kav84]. Pines [Gal87]. Pink [BWMB88, GSMH88, Hau86, KT87, MR83, SBJD83, TS84, WS82]. Piscicidal [DHSA83]. Piscivores [JSC88]. Piscivorous [DKT88, HO83]. Piscivory [KMC84, TB87]. Pitfalls [GM87]. Plaice [DRV85, RVV85]. Plains [Bro86, BHN84]. Planktivory [KS86]. Plankton [DTLK84]. Planktonic [Mar83, Mar86]. Planning [Cun89, LNK87]. Plant [PAH89]. Plants [BGC81]. Plasma [CMNS85, CW88, DP86, SS84, SBJD83, ZB84]. Plata [QB85]. platessa [DRV85, RVV85]. Pleuronectes [DRV85, RVV85]. Ploidy [AS83, BB83]. Plume [RW81b]. Point [Gof86, PM83, RHU82, TSD81]. Polarized [KC89]. Policy [MI89, Ran87, Van87a]. Political [Yar87]. Pollock [Bea81, KG89]. Polluted [Ang83, HR88]. Polyculture [CPML85]. Polymorphisms [PZI89]. Polyploidy [TJS81]. Pond [Boy85b, CL82, FK88, LLBA86, ØKF85, SB83]. Pond-Raised [LLBA86]. Ponds [Boy81a, Boy81b, Boy82, Boy85a, BB82, BS82, CPML85, HB81, JBL88, LMW81, MB83, MRTR82, ST83]. Pool [HBD88, HS83]. Poor [Cou87]. Pop [LJL86]. Population [BHR85, BWQ84, Bro86, BMU87, Dol87, Ell86, FI88, GC84, Goo85b, HC84, HG84, Kwa81, LF83, LA84, LM87, MSB88, ML84, PZI89, RNT85, Ros88, SSS86, Ser82b, SWRM85, VAKT87, WH82b]. Population-Specific [BWQ84]. Populations [AM89, Ang83, Cob82, Dia87, Erm86, GA85, HR87, IM83, Kay89, KM87a, LM85a, LM85b, Lem85, New81, Rah81, RM87a, SRM83, Tyl88, WB85]. Portions [HE84]. Possibilities [Bay81]. Possible [Col84, CCP87, HH82a, KC89]. Postemergent [Sco85, SBJD83]. Postlarvae [CWCR89]. Postpreservation [McG85]. Potamodromous [MYA81]. Potassium [JP81]. Potential [Cou81, FK88, GD85b, HS83, MO88]. Potomac [BZP81, HZB^{+}89, MWMSH85, SHBM^{+}81]. Pots [BOY84]. Power [BGC81, PM81]. Powered [AB82]. Practice [KS81]. Practices [Han82, JO82]. Pragmatic [Ald85]. Prairie [LM85b, MS84b]. Precipitation [Hai81]. Precise [Lei85]. Predation [Bar82, BMM^{+}87, DRV85, GSC81, GC87, LM87, MO88, MF83, OT86, SCH81, TM85a, TSC84, WS88]. Predator [EB86, MCK86, SS82a, TSC84]. Predator-Prey [SS82a]. Predators [SKC81]. Predict [SJ81]. Predicted [JH82]. Predicting [HN84]. Prediction [KLJ84, SP85]. Preference [BHN84, BP83, SHS81, SZS82, TB81]. Preferences [BSS^{+}88, DB86, MSPS81, PAH89, SZ85, YMH83]. Preferenda [KWP88]. Preferred [KG83, PB83]. Preliminary [BJCM87]. Premolt [BOY84]. Prepared [LD86]. Prerecruit [BP84]. Presence [EG86]. Present [Tis89]. Preservation [Ker83]. Pressure [Gib84]. Prevalence [HMSK86]. Previously [GC81]. Prey [AMP82, Bar82, Bre88, HDLG88, HM87a, Hin85, JSC88, JTDM87, Kre86b, LO83, MCR84, MCK86, MPSC89, SS82a, SZ85, TSC84, WS88, Wet88]. Pricing [Bed87]. Prickleback [HN84]. Primary [Ger84, HFKR81, LMW81]. Private [HWD88]. Private-Boat [HWD88]. Probable [BSM82]. Problem [Van87a]. Problems [AW85, BP87]. Procedure [MJE89, New81]. Procedures [Ken89, PG87]. Proceedings [Cun89, TDG^{+}87]. Process [GA85]. Produced [KBH85]. Product [THV87]. Production [BNW88, BA85, BH88, BS82, CW82, Cou81, Dia83, FI88, Har86, HFKR81, Hil89, HDGE82, JO82, Jen84, LMW81, NP83, ØKF85, PN87, Rei83, Rim85, SB86a, Wat83, WS83]. Productivity [CPML85]. Products [Fos85, TGRA82a]. Progeny [HHEM86]. Programs [Boy81a, SSU86]. Progress [Tis89]. Prolarvae [HZB^{+}89]. Properties [BHL^{+}85, HMMJ81b, HMMJ81a]. Proportion [Koh81]. Protandrous [MWSR83]. Proteases [TGRA82b]. Protects [HM87b]. Protein [DP85, GA83, HN84, Mil82, Mil83]. Protein-Synthesis [GA83]. Proteins [Fab87a]. Protocol [MO86]. Protogynous [EG86]. Prototroctes [BSM82]. Protozoan [WT88]. Pterygiophore [HD87, OGS^{+}83]. Pterygiophore-Interdigitation [OGS^{+}83]. Public [DF87]. Publication [Ken81]. pulex [HW81, MF83]. Pulpwood [MENG86]. Pulses [KP87]. Pump [PV88]. Pumped [Fle85, VP88]. Pumpkinseed [Eva84, HR87]. Pumpkinseeds [FS83]. Pupfish [EE89]. Pure [DSHB83]. Pure-Strain [DSHB83]. Purification [TGRA82b]. Purified [SMLM83]. Purpose [DF87]. Pyramid [Sny83].

Quadrat [JSC88]. Quality [CTSD84b, CVC85, CPML85, HPHF85, HZB^{+}89, HG87, JST^{+}81, LNV83, MJE89, MMY^{+}82, MHEG89]. Quality-Induced [CTSD84b]. Quantify [HJ85]. Quantitative [KLM86, SHI84]. Quantity [GS82a]. Quebec [FP89]. Quota [BP87].

Rabbits [MJ83]. Raceways [CW88]. Radio [MW86, SK86]. Radio-Tagged [MW86]. Radiographic [WGR87]. Radioimmunoassay [HJ85]. Rain [THV87]. Rainbow [AMCD85, Arg82, BSDD86, BB84, CLS87, CJ85, CH85b, CG84, DBBF83, DA82, FJT89, FBCV82, FS85, GGS82, HC84, HFDL83, HPG83, IN82, KM87a, Kwa81, LM85a, LAPK87, LB84, LA84, Mar84, MMW85, MJ83, Rei83, Rim85, SLJ^{+}84, SFS83, SP85, TJS81, TR83, TRL^{+}84, TJ83, WD88, WSL86, WH84, WS83, WS87, ZSNC87]. Raised [LLBA86]. Rakers [DHVG87, WOL83]. Random [KWDS81]. Randomization [PH89]. Range [BHR85, MW86, NN82, SFW83]. Ranges [FS83]. Ranging [Bac84]. Rapid [DHSA83, HDS86, TCS88]. Rapidly [ZR81]. Rare [BP85]. Rate [DT87, GMW83, LN87, NG86, Rei83, SWM87, WD88]. Rates [CD83, DMO82, FB83, Gar81, SB83]. Ratio [HE84, OK86]. Ration [CC81, Ger84, WC83]. Rationality [Van87a]. Ray [Bea81]. Razorback [BP83, BMU87, Mar87a, Tyu87]. Really [Ret87]. Reared [BL83a, Car83, EWE^{+}81, GSC81, MBM89, WS87]. Rearing [FMS81, MR89, PSBZ86, PTMG89]. Recapture [Jen81, MS89]. Recaptured [Hil88, Mat81]. Reciprocal [FLK^{+}87, KWP88]. Recognition [CP81]. Recommendations [Tal87]. Recommending [OM82]. Reconsider [Con89]. Recreational [BSP87, Bed87, Edw89, EH86, Mat81, PG87, RTP85]. Recruit [DRH88, ED88b, MH84, RF88]. Recruitment [Car86, GCC85, GA85, HG84, Jen84, KT89, MI89, Pau82, PH89, RCH87, Ser82b]. Recurring [SP88]. Red [CLSRS89, CV89, Gal86, Gar83, HKA83, HA83b, LHA84, LO83, NM82, Tan85]. Redd [LF89, RW83, WM83]. Redd-Site [WM83]. Redds [BNF82, CWW^{+}86, Cha88, NB85]. Redeye [BM83a]. Reduced [CEG88, Cou87, GD85b, LAK85, MBRK81, NB85, Rim85, RJ85, SJR87]. Reduction [Ell86, HA83a]. Reductions [KR85]. Reference [Bed87, EB86, TC89]. Refinement [GA83]. Reflections [Bal84]. Refuges [Mos85]. Regeneration [MYA81]. Regimes [Wre84]. Region [WMRH86]. Regional [FKY84, Nak87]. Regulates [Mei84]. Regulating [KO83, MMI88, RCH87]. Rehabilitation [Esh87]. Related [FMS81, HM85, HS83, KT85, KTKD82, QB85, TM85a]. Relation [Cip83, Cob82, GMB^{+}87, HPHF85, PS81a, Rig82, Ser82b, SP85, ZB84]. Relations [LHW87, LM85b]. Relationship [BE82, DT84b, Goo85a, HFKR81, KG83, MM88a, MHH^{+}82, Pau82, ST83, SK81, WD88]. Relationships [AK84, BR86, CC85, Daw88, GSMH88, HG89, PM83, WJ87]. Relative [BSZF83, CB88a, GDH83, MPAR86, PV88]. Relatively [ML84]. Release [KLM86]. relicta [BL83a, RF81]. Remains [TB87]. Removal [DHVG87, Ell86]. Renibacterium [PEMM87]. Repeat [Lei85]. Replacement [GCC85, Wat83]. Replicated [Str82]. Reported [Goo85a]. Reporting [GMW83]. Reproduce [ATH86]. Reproduction [AV84, DT84a, GB86, Mac82, Man84, NN82, SWCB83, Tay84, Tyu87, Wal86b, WSL86, Wre84]. Reproductive [BS81, CLL86, Cou87, DMH84, DSHB83, EG86, Fos85, Gof86, GLC82, MDS84, MMR89]. Requirement [BM83b]. Requirements [HRN83, PS81b]. Reserves [TGC^{+}86]. Reservoir [CB88b, Cou87, Gri81, HG84, JO82, KM84, KWDS81, KN81, MMNW81, MHEG89, MENG86, MNT83, PM83, TSD81, VP88]. Reservoirs [BM83a, JH82, Mat85, Mos85, Ney81, WB81]. Residence [NG81, Pec82]. Residency [FGB^{+}88, Mes88]. Resident [FP89, Jon85]. Resistance [Cip83, RJ85]. Resource [CC84, Gre87, Hig87, LNK87, OWNH88, PRCM84, SJ89a]. Resources [BBW87, MRG87, TDG^{+}87]. Respiration [BHN84]. Respiratory [PWM81b]. Response [CRF83, DP86, ED86, GS82a, MS88a, MK89, MG88, Ran86, SCSR86, SW88b, SMLM83, SJR87, TM85b, TCS88, WC86].

Responses [BSS86, BWQ84, CH86, FCL^{+}83a, GS82b, KP87, MW81, RC87, SSGT84, WM84, WS87]. Restricted [Car86]. Results [Rah81]. Retained [Joh83b]. Retention [JV87, WOL83]. Retorted [WRH^{+}85]. Return [Hil88, TS84]. Returning [BWQ84]. Reversal [BSYW84, JSYW83]. Review [Cha88, Cra84, Hai81, HM86, LM86, NS84b]. Reviews [Ano81a, Ano81b, Ano81c, Ano81d, Ano81e, Ano82a, Ano82b, Ano82c, Ano82d, Ano82e, Ano82f, Ano83a, Ano83b, Ano83c, Ano83d, Ano83e, Ano84a, Ano84b, Ano84c, Ano84d, Ano84e, Ano84f, Ano85a, Ano85b, Ano85c, Ano85d, Ano85e, Ano85f, Ano86a, Ano86b, Ano86c, Ano86d, Ano86e, Ano87a, Ano87b, Ano87c, Ano87d, Kre86a, Ush87a]. Revised [LN87]. Rheotaxis [Kay89]. Rhode [Fab87a, SB86b]. Rhythmic [Gib84, Kav84]. Rhythmicity [MDS84, SK84]. Rhythms [DSP84]. Richness [Min89, RW86]. Riffle [Fis87]. Rigs [RTP85]. Ring [RCB85]. Rings [MS82, MS84a]. Risk [CWSM86, Cou85b, Fle85]. Risks [Cou85a]. River [AB83, BSP87, BSS^{+}88, CJ85, CWW^{+}86, Dey81, GC83, GD85b, HHB81b, HG87, LF89, MWMSH85, MO88, MR89, MS84b, NG86, OIWD81, PCC85, PTMG89, QB85, SHS81, Sny83, Tyu87, Uph89, VJ86, Wal86b, WO86, AV84, Bay81, BMW89, BP85, BC84, Bro86, BK85, Cha86, CVC85, Cob82, Col84, CB83, DCR89, Fab87a, FK88, HZB^{+}89, HSR^{+}86, HE84, HWGF82, HBB84, HBD88, HS83, Hol86, HHN87, IN82, MMNW81, MSBB88, MCHL81, MBC85, NJS86, OK86, PCC85, PB83, RHU82, Rug86, SB86a, SWCB83, Sco85, SH84, WFM^{+}89]. Riverine [SÍW89]. Rivers [BHL^{+}85, HG89, RPI89, SD83, DB86, FB88, KZ83]. Rock [FC87, MJL89, PN87, PRCM84, SHCMS83]. Rockfish [Car86, Fra81, LN87]. Rockling [Her85]. Rocky [LHW87, Mar86, SB86a]. Role [BH83, ED86, Ker82, KC89, MCH87, PG87, Smi89, WS88]. rosenbergii [CPML85]. Rotenone [DHSA83, JSC88, MWG82, RG86, SCD81]. Run [CLL86, DPR89, Jon85].

Sac [HPG83, JB87, KMC87]. Sacramento [KO83, SKMK85, Col84, JV87]. Sacramento-San [KO83, SKMK85]. Safe [CGT82]. SAFR [TDG^{+}87, Tal87, TL87]. Sagehen [Erm86]. Saline [HA83a, PHD85]. Salinities [DP85]. Salinity [MK89, MP84, MRC81, MR82, RS83, ZG89]. Salmo [RJ85]. Salmon [BSS86, Bea84, BWMB88, BMW89, BST89, BNF82, BSN88, BNW88, BL83b, BWQ84, BQW^{+}86, BH88, BE82, CWW^{+}86, Cha86, CW88, CCMD81, Cun89, Dee81, DB86, Dol86, Dol87, EH83, Eri84, FMS81, FW86, FV82, GSMH88, Han82, HA83a, Har86, HCFE81, Hau86, HHEM86, Hil89, HGP87, HDGE82, Joh83b, KMTU84, KT87, KLM86, LF89, MPAR86, MSBB88, MHS88, Mer86, MS88b, MBRK81, MR83, Mol83, MM88b, MR89, NG81, NG86, NB85, Ors87, PEMM87, PSBZ86, PS81a, PMR82, PMR86, PTMG89, Rat81, RS83, RSE87, RJS85, RPI89, RM87a, SHI84, SÍW89, SSU86, SW88b, SBE84, Sos82, SS82b, SS84, Sti89, SK86, SBJD83, TM85a, TM85b, Tay88, TGC^{+}86, TS84, Tis89, WS82, WM86, WQB82, WB85, YMH86]. salmonicida [CRF83, Cip83, CH86, DS85]. Salmonid [DS85, RW83, SKC81]. Salmonids [Cha88, Cip83, CH86, GW83, OWNH88, Rug86]. salmoninarum [PEMM87]. Salmonines [JTDM87]. salpa [Ger84]. Salt [Cou85a, RTV84, TAS86]. Salt-Marsh [RTV84]. Salts [SL86, SS84]. Saltwater [BCNW87, Mat81, ØKF85]. Salvelinus [KR85]. Sample [Jen81]. Sampler [Fis87]. Sampling [AAC88, BLD89, DRFL89, GC83, HM85, HCG89, KWDS81, Kus81, LJL86, MWG82, Ors87, PV88, RMC85, SCD81, TCKW83, Tho86]. San [KO83, SKMK85]. Sarpa [Ger84]. Sashimi [Tho86]. Sashimi-Grade [Tho86]. Satisfactory [Hoe87]. Saturation [HMJ87a]. Savannah [AV84]. saxatilis [Bod85]. Scale [AAC88, BCNW87, Car82, Car83, Enn84a, GW83, MYA81, RC82]. Scale-Number [Car83]. Scale-Sampling [AAC88]. Scales [GA83, Joh89, PD88]. Scallops [Gar83]. Schedule [SN84]. School [Car86]. Schools [Kol85]. Science [Con89, Ker82, Lof87, MCH87]. Scientific [Ken81]. Scombrid [Tho86]. Scophthalmus [Qua85]. Scotia [Jes87]. Screens [SJ81]. Sculpin [Bra86, MJM^{+}83]. Sea [BE82, CFCSI84, Fab88, Gar83, HKA83, Jon85, SSGT84, Tan85, Cra84, DRV85, HC84, RVV85]. Sea-Grass [HKA83]. Sea-Run [Jon85]. Search [TL87]. Season [GMB^{+}87, MHEG89]. Seasonal [Col84, CCP87, DT84a, Dol87, Enn84a, FD85, FCBTSR83, HS89, HN84, Joh83b, Kir81, Mei84, Vet82]. Seasonality [GDH83]. Seasonally [AMP82, BH88]. Seasons [BH86]. Seatrout [MHLL87, Vet82, VO87]. Seaward [HCFE81, OK86]. Seaward-Migrating [OK86]. Seawater [MPAR86, MR83]. Seber [HG84]. Sections [Bar85, Bea81]. Sediment [MRN83, MHA81, PTMG89]. Sedimented [HGP87]. Sediments [CCMD81]. Seepage [SP85]. Segregation [HJ83]. Seine [BLD89]. Selected [HDS86, LNV83]. Selection [AAC88, CB88b, CZCP84, GC87, HDS86, HW81, HGP87, JBL88, JTDM87, Kre86b, MHS88, MCR84, MCK86, Mos85, MW88, Rah81, Ran86, RMC85, SCSR86, Str82, TCS88, WM83]. Selective [Bar82, MD86, RVV85, WS88]. Selectivity [ED88a, Gar81, Wet88]. Selenium [CV89, GB86]. Selenium-Contaminated [CV89]. Seminatural [Rim85]. Sense [EB86, KW89]. Sensitivity [PHD85, PC85, RJS85, WFM^{+}89]. Sequences [Mab89]. Sequential [Hil88, IN82]. Sera [HM87b]. Serum [CH86, DM81, Hon84]. Seven [LAPK87]. Seventh [Sny84]. Several [DS85, FG83]. Sewage [OIWD81]. Sex [BSYW84, GRSD83, HE84, HJ83, HBB84, JSYW83, OK86, SK81, SGK^{+}89, ZB84]. Sexual [Bow89, Sta88]. Shad [CB83, DSM85, DMO82, HWGF82, IZ83, KM84, Mat84, MW88, OK86, PV88, PWM81a, PWM81b, SWM87, Sto86, TK85, VP88]. Shale [WRH^{+}85]. Shallow [MS84b, Pec82]. Shape [RC82, Sta88]. Sharks [Smi84, WGR87]. shasta [BSZF83, HHEM86, Rat81]. Sheep [OIWD81]. Shelbyville [Sto86]. Shelf [SWT^{+}85]. Shelikof [KG89]. Shifts [Bra86, CC84, IZ83]. Shiner [SHS81]. Shiners [BB84, CV89, Gal86, Kir81, PM81]. Shock [TJS81]. Shore [SHCMS83, RPI89]. Shoreline [PCC85]. Short [Daw88, Heg88].

Short-Finned [Daw88]. Short-Term [Heg88]. Shortnose [BK85]. Shovelnose [HHN87]. Shrimp [BP84, BL83a, CC85, HH82b, KO83, MZM89, MP84, PLC86, WCM88]. Shrimps [Pau82]. sign [MPP82, STS87, YMH83]. Significance [IN82, SZS82, WOL83]. Silver [KMCB84, RA86]. Silverside [Kol85, MDS84]. Similarities [MSPS81]. Simple [RF88]. Simplified [Car87]. Simplifying [SD87]. Simulated [CG84, Goo85b, SS82a, SCSR86]. Simulation [Bor83, Dia87, ED88b, KT89, OT86]. Single [Jen81]. Site [WM83]. Sited [BGC81]. Sites [AAC88, GW83]. Sixes [NG86]. Sixth [CE83]. Size [AR82, Bar82, BWMB88, BW89, BHR85, BE82, DC82, ED88b, FI88, GMB^{+}87, GA85, HM87a, HBB84, HG84, JBL88, KT85, MCK86, MPSC89, MD86, SFW83, SZ85, SK81, SP85, TM85a, WD88, WDS85]. Size-Related [TM85a]. Size-Selective [Bar82]. Size-Structured [ED88b]. Sizes [Jen81, LF83]. Skate [War84]. Slimy [Bra86]. Small [AK84, Boy82, BS82, Dol87, EE89, Enn84a, GW83, HW81, Heg88, Hol86, LHW87, LO86, LM87, MG88, NP83, Ors87, SJ81, SCD81]. Small-Mesh [SJ81]. Small-Scale [Enn84a, GW83]. Smallmouth [CD83, Gof86, GO86, PRCM84, Ran86, SCSR86, Ser82b, Ser84, SWRM85, TR89, WDS85, Wre84]. Smelt [AMCD85, Arg82, FS85, HC84, LA84, SLJ^{+}84, TJ83]. Smolt [RJS85]. Smoltification [SHI84]. Smolting [Joh83b, PSBZ86]. Smolts [CW88, MPAR86, MHS88, SK86]. Snake [WFM^{+}89]. Snapper [Joh83a]. Snappers [NM82]. Snowy [MH84, ML84]. Social [Gal87, HA86, MCH87, TDG^{+}87, Ush87b, Van87a, Van87b]. Society [Noa88]. Socioeconomic [Esh87, WH82a]. Sockeye [BST89, KLM86, LF89, WB85]. Soft [CLS87, MJL89]. Soft-Water [CLS87]. Sole [FC87, KT89, PD88]. Soles [KTKD82]. Solids [PB83, RSE87]. Solubility [Boy81b]. Soluble [DGS^{+}85]. Some [Bal84, CMNS85, Car82, SZS82]. Sound [Pow82, WCM88]. Source [GS82a]. Sources [MBC85, PS81b, Qua85]. South [MH84, QB85, SHCMS83, BM83a, BOWY83, EG86, FB88, FS85, GLC82, HE84, Joh83a, OWNH88, PTMG89, Rec86, RH82, SW85]. South-Central [OWNH88]. Southeast [Dol86, Dol87]. Southeastern [BO89, FD85, HMSK86, MM88b, YHW89]. Southern [NN82, Pre89, SW85, Wre84, Bea84, Bow89, BR86, CLS87, CFCSI84, LM85a, Mil81, Nob81, PLC86]. Southwest [Res89]. Southwestern [KM84, MHEG89, WM83]. Spanish [ED88a, ED88b]. Spatial [KYFS87, Mar83, SHBM^{+}81]. Spawn [Kay89]. Spawner [GB82, NG81]. Spawners [Gal83, GD85a, Gal86, Lei85]. Spawning [BS82, CP83, CP86, DMH84, GB82, Gal83, GD85a, Gal86, GO86, Hay86, INKP87, LF89, MJE89, MW86, MS84a, NG81, PTMG89, SD83, SP85, YHW89]. Special [Bed87]. Species [CG84, DP86, DC82, HD87, JST82, JTDM87, KG83, LM81, LKF82, MPP82, Mil81, Min89, OGS^{+}83, Pin82, RW86, RW81b, Sta88, TAS86, WB81, Wre84]. Specific [BWQ84, WD88]. Speckled [MH84]. Speculative [Cou85b, PFT^{+}85]. Speed [CWCR89, TCKW83]. Sperm [BSDD86, Dup82, Ker83]. Spermatozoa [KBH85]. Sphyrion [Enn86]. Spider [RC87]. Spiny [GLC82, Mac82, NMC85]. Sport [Bro87, Mar87b, PM83, PS81a, SJR87]. Sport-Fisheries [Bro87]. Sport-Fishing [PM83, PS81a]. Sportfish [JH82]. Sportfishing [HWD88, THV87]. Spot [HH82a, PC85, VO87]. Spotfin [Kir81]. Spots [HHB81b, WO86]. Spottail [PM81]. Spotted [CL82, MHLL87, VO87]. spp [Her85]. Spring [BE82, DPR89, HCFE81, KM84, MMNW81, MPAR86]. Spring- [DPR89]. Spring-Summer [KM84]. Springs [EE89]. Springtime [RTV84]. Squawfish [AM89, WRH^{+}85]. Squid [Daw88]. St [BNW88, RPI89]. Stability [STS87]. Stage [PCC85]. Stages [CLH^{+}86, MO88, RJS85, Smi89, WW86a]. Stamina [BHL^{+}85]. Standard [Car82, Eva84, MO86]. Standing [Arg82, LHW87]. Starry [Cam84]. Startle [TM85b]. Starvation [PC85]. State [FCL^{+}83b, Mab89, MWMSH85, Edw89]. States [MH84, BHL^{+}85, Mat85, NMC85, Ney81, Nob81, War84, WB81]. Static [MPAR86, PAH89, SFS83]. Static-Water [SFS83]. Station [PM81]. Statistical [Smi85]. Steelhead [BS87, BSN88, BSZF83, CB88a, CB89, Cha86, CLL86, CW88, DPR89, Lei85, PT88, RSE87, RJ85, SW88a]. Steelheads [CRF83, SBE84]. Steering [TDG^{+}87]. Sterile [HDGE82]. Steroid [GRSD83]. Steroids [DM81, ZB84]. Stochastic [GC84, MI89]. Stock [BA83, HH87, Koh81, KM87b, LHW87, LB82, LA84, MBM89, ML84, Pau82, Pra88, SE87, SLJ^{+}84, SSU86]. Stock-Recruitment [Pau82]. Stocked [BNW88, GMB^{+}87, JSC88, KMKM89, SCH81]. Stocking [BS82, DA82, GMB^{+}87, MMI88, MSC86, MNT83]. Stocks [Arg82, BP84, BA85, BMU87, CM83, DSHB83, EE89, Fab87a, Fab87b, HHEM86, INKP87, RC82, SÍW89, SW88a]. Stokely [Kwa81]. Stomachs [MM88b]. Stone [Res89, SH86]. Stonerollers [Car83]. Storage [FD89, VP88]. Stores [WM84]. Strain [DSHB83]. Strains [BS81, BSZF83, KMKM89]. Strait [FC87, KG89]. Strategies [Noa88]. Stratification [MHS85]. Stream [AK84, Bac84, BVBM87, CHW87, CP86, CG84, Dol86, Dol87, FKY84, FW86, FBCV82, FHT84, HR88, HCFE81, Heg88, HGP87, LHW87, LM85a, Man84, Mer86, Mes88, MG88, MB85, MHA81, NP83, New81, PRCM84, Ran86, RB86, SCSR86, Tay88, TR89, Wat83, WS83, YHW89]. Stream-Dwelling [PRCM84, TR89]. Streamer [HH82b]. Streamflow [GO86]. Streams [BLD89, BW89, BSN88, BNW88, BR86, CLS87, CHW87, CCMD81, CCP87, JST^{+}81, JB87, KYFS87, Kay89, LHW87, LM85b, Lem85, LO86, PN87, SB86a, WM83]. Strength [Ser86]. Stress [BSS86, BS87, CTSD84a, CTSD84b, DP86, FMS81, HM86, HMJ87b, JH83, MPAR86, RS83, TGC^{+}86, VHH86, WC86, WS87]. Stress-Related [FMS81]. Stresses [SW88b]. Stressors [MSC86]. Striped [AV84, AW85, BL86, BCNW87, Bod85, Bor83, BA85, BZP81, Bre88, BHL^{+}85, CVC85, Col84, CM81, CP81, CV89, CZCP84, Cou85a, Cou85b, Cou87, CC81, Dey81, DM81, EWE^{+}81, Fab87a, Fab87b, FG89, FB88, GCC85, Goo85a, Goo85b, HPHF85, HZB^{+}89, HB86, KLJ84, Ker83, KBH85, KHS81, Koh81, MBM86, MWMSH85, MHS85, Mat85, MHEG89, MO88, MCHL81, MHH^{+}82, Mil82, Mil83, MRC81, MRN83, MWSR83, Mos85, PHD85, PFT^{+}85, Rad84, RW81a, RNT85, RM87b, SRM83, SHBM^{+}81, SKMK85, Uph89, WOR85]. Structural [HZ89, Rad84]. Structure [And87, Bro86, Car86, Erm86, Hud84, JBL88, Kol85, Kru88, LA84, MG88, SLJ^{+}84]. Structured [ED88b]. Structures [LJL86]. Studies [AAC88, Ald85, HZB^{+}89, ML82, RW81a, Smi89]. Study [MS88a, SHM83]. Studying [FCL^{+}83a]. Stunted [HR87, KS86]. Stunting [Dia87]. Sturgeon [BC89, LD86]. Sturgeons [BK85, HHN87]. Subject [KR85]. Sublethal [Kol85, SSGT84, WMM81]. Submerged [MENG86]. Submersed [SS82a]. Submersible [GAT82]. Subsample [Kre86b, Str82]. Subsamples [Str82]. Subsistence [MRG87]. Subspecies [LAPK87, Mac89, PCW83]. Substock [HG85]. Substrate [YHW89]. Substrates [SP85]. Subyearling [CB89]. Success [CLL86, Cou87, DMH84, Gof86, KT85, MMI88, MCR84, MNT83, RW83]. Successive [MHS88]. Sucker [BMU87, CP83, Tyu87]. Suckers [BP83, CP86, Mar87a]. suckleyi [MP84]. Sudden [KR85]. Suitability [PN87]. Suitable [EH83]. Summer [BNW88, BSZF83, CLL86, Cou87, HGP87, JH82, KM84, MMNW81, MHS85, Mat85, Mos85, OWNH88, Pow82, Ser84]. Summer-Run [CLL86]. Sunapee [JH83]. Sunfish [Lem85, SJ82]. Superior [DPR89, KM87a, KM87b, Kwa81, LA84, SLJ^{+}84]. Superposed [PH89]. Supersaturation [Bou82, CW82, Col84]. Supply [MWMSH85]. Support [Lac89]. Suppression [Lem85]. Surf [SK85]. Surgically [MS86, SM84]. Surplus [Jen84]. Surplus-Production [Jen84]. Survival [BH83, CB89, CH85a, CEG88, Dee81, GC84, GMB^{+}87, Hin85, Hol86, HA83b, JB87, KBH85, LD86, MM85, MJL89, MHS88, RCH87, Ser84, SH86, SFS83, SP85, Uph89, WM86, WOR85, WJ87]. Susceptibility [BSZF83, CH86, Har87, HHEM86, MBM86]. Suspended [RSE87, WQB82]. Sustainability [MI89]. Swim [MMB81]. Swimming [BP85, BHL^{+}85, CWCR89, Eri84, Fui86, TM85a]. Switch [LD86]. Swordfish [GDH83]. Sympatric [CP86, GV83, Vet82, WS83]. Symposium [Cun89, Mar85, Noa88, Ogl82, Tal87, TL87]. Synchronization [Tay84]. Synergism [Mei84]. Synthesis [GA83, LL82, Res89, SB86c, TL87]. Synthetic [MBB^{+}84]. System [BSS^{+}88, FCL^{+}83a, GGS82, HSR^{+}86]. Systems [ACG^{+}82, Fle85, GSC81, Lac89, LM81, Wal86b].

Tag [Enn86, GMW83, Hil88, VO87]. Tag-Reporting [GMW83]. Tagged [Mat81, MW86, TS84]. Tags [EH83, Hil88, HH82b, MNT83]. Tailwaters [Mos85]. Taku [LF89]. Tamiahua [FCBTSR83]. Tanana [HG89]. Taneycomo [WH82a, WH84]. Tanganyika [Cou81, HFKR81]. Tanks [CW88, SZ85]. tarda [HB86]. Taxonomic [Edi88]. Technical [ZSNC87]. Technique [HMJ87a, JSC88, Tho86]. Techniques [BM87, BOWY83, BC89, FG83, GM87, HM85, PEMM87]. Telangiectasis [HM85]. Telemetry [HSB^{+}83, SK86]. Teleost [ED86, EB86, KC89, ZB84]. Teleosts [HZ89]. Temperature [BVBM87, Bor83, BHN84, BP83, CD83, CZCP84, Cou85b, CC81, DT84a, Eva84, Fui86, GO86, MSPS81, MP84, MRC81, MR82, NB85, Pre89, RW81a, SWM87, Ser82b, SHS81, SZS82, Tay88, TGC^{+}86, WD88, WBH85, Wre84, WC83]. Temperature-Dependent [CD83]. Temperatures [BL83a, Car83, CEG88, KG83, MW81, STP82]. Temporal [BC84, KYFS87, Mei84, MS84a, Nak87, SHBM^{+}81]. Temporary [LM88]. Tennessee [Wal86b, CVC85]. Teratological [MJ83]. Term [Erm86, Heg88, Rec86]. Terminos [CFCSI84]. Terms [Jen84]. Territorial [Enn84b]. Tessellated [GD85a]. Test [EG86, HR87, MJE89, MPAR86, PH89]. Tested [JST^{+}81]. Testing [LO86]. Tetracycline [Het84, Smi84]. Tetracycline-Injected [Smi84]. Texas [MHS85, FG89, FH89, MHLL87, MMI88, Mat81, Pre89]. Texoma [MHS85]. Thailand [BP84]. Their [FLK^{+}87, Kru88, KC89, KR85, MM88a, Mac89, SKC81]. Theory [KS81, LNK87, Yar87]. Thermal [BLR81, FLK^{+}87, KLJ84, KWP88, MB82, Mos85, RW81b, SWRM85, ZR81]. Thiobencarb [FF86]. Thiocyanate [HTMZ85, MS88b]. Third [Cun89]. Threadfin [IZ83, PV88, VP88]. Threatened [SB86a]. Three [BM87, BP85, BH86, Dia83, FB88, HHEM86, JST82, KYFS87, Lon88, PT88, PMR86, WS88]. Thymallus [Sco85]. Thyroid [SBJD83]. Thyroid-Hormone [SBJD83]. Thyroxine [SS84, SN84]. Tidal [BHR85, DGK^{+}88, HSB^{+}83, RVV85, WO86]. Tiger [GSC81, MSC86, STS87, SCH81, TSC84]. Tilapia [Pin82, Bow81, DP85, DHVG87, KW89, LL82, TCS88]. Tilapias [SZ85, ZG89]. Tilefish [EG86, GAT82, HG85, RH82]. Timber [VP88]. Time [BE82, CB88a, GO86, NG81, NS84a, SJ82, SW88b]. Times [CH85a]. Timing [Lei85, Smi84]. Tissue [FD89, GB86]. Tissues [CWSM86, MS88b, RG86]. Titicaca [WT88]. Tofino [SJ89b]. Tolerance [Ang83, BNF82, MB82, MP84, NB85, Pre89, Rah81, SZS82, WBH85, ZG89]. Toluene [MBRK81]. Total [BBW87, PB83, Ran87, TL87, Van87a]. Tow [CC85]. Towing [TCKW83]. Toxaphene [HMMJ81a]. Toxic [Goo85b]. Toxicant [CGT82]. Toxicities [FV82]. Toxicity [CH85a, CLH^{+}86, CV89, FB88, HA83a, HTMZ85, LM86, LGP83, MTB81, MTBB81, MMW85, MBB^{+}84, SL86, TR83, TRP83, TRL^{+}84, TRM^{+}86, WRH^{+}85]. Toxins [DKT86, DKT88]. Trace [RB86]. Trace-Metal [RB86]. Tract [Mar87a]. Tractor [AB82]. Tractor-Powered [AB82]. Tracts [WGR87]. Traffic [Hol86]. Transactions [Ken89]. Transfer [AB82, DEAT87, HM87b, Lof87, SB83]. Transformer [MMW85]. Transintestinal [SM84]. Transmitters [CH85b, MS86, SM84]. Transport [GT85, Hau86, LVNM82, MS84b, NS84b, RVV85, SWT^{+}85, SHM83]. Transportation [ACG^{+}82, DP86, JM82]. Transporting [MSBB88]. Traps [Kus81, RC87]. Travel [SD87, THV87]. Trawl [CC85, Nie83]. Trawls [DRFL89]. Treatment [JP81]. Trematodes [MM88b]. Trends [CM83, Goo85b, HH82c, Rec86]. Tributaries [BH88, DPR89, FW86, KM87a, KM87b, MR89, SW88a, WS82]. Tributary [DCR89, HPHF85, Man84]. Trihalomethanes [MTBB81]. Triploid [ATH86, BSDD86, PAH89, STS87, WW86b]. Triploidy [WLC81, WLC82]. Trolling [Ors87]. Trout [Bac84, BSDD86, BSN88, BR86, BB84, CLS87, CJ85, CH85b, CHW87, CLH^{+}86, CG84, CCP87, DBBF83, DA82, Edi88, Esh87, FW86, FJT89, FBCV82, Fos85, FP89, GGS82, GS82b, GMB^{+}87, HMJ87a, HMJ87b, HJ83, Heg88, HTMZ85, HFDL83, HRN83, HPG83, IN82, JH83, Jon85, JB87, KS86, KM87a, KM87b, KMKM89, Kwa81, KR85, LHW87, LM85a, LAK85, LAPK87, LB84, Lue86, Mar84, MMY^{+}82, MMW85, MO86, MS89, Mes88, MJ83, MG88, MYA81, Pec82, PZI89, Rei83, Rim85, STS87, SD83, SFS83, Sos82, SP85, STP82, SJR87, TJS81, TR83, TRL^{+}84, Tro89, WD88, Wat83, WSL86, WH82a, WH84, WS83, WM83, WMM81, WS87, WFM^{+}89, YHW89, ZSNC87]. Trouts [SB86a]. Truckee [SWCB83, Sny83]. True [RCB85]. Trypsin [HJ85]. Trypsinogen [HJ85]. Tubercular [SRM83]. Tui [GV83]. Tumor [BSP87]. Turbid [Mat84]. Turbidity [Bre88, Gar81, GS82b, SBE84]. Turbine [SK86, TK85]. Turbines [GW83]. Turbot [Qua85]. Turbulence [KMC87]. Two [BM83a, BL83a, DRFL89, DB86, DPR89, GV83, GC83, GC81, Har86, Mac89, MW86, MBRK81, MM88b, MW88, PN87, RM87a, WRH^{+}85]. Typology [Bro87].

ui [SWCB83, Sco88]. Ultrasonic [HSB^{+}83]. Ultrastructural [MS88a]. Uncertainty [RF88]. Underlying [Mab89]. Unexploited [ML84]. Unfished [GC81]. Unit [BA83]. United [BHL^{+}85, NMC85, Nob81, WB81, MH84, Mat85, Ney81, War84]. Upper [HZB^{+}89, HBD88, HS83, Hol86, HHN87, MMR89, PCC85, RM87b, SH84]. Upstream [Lei85]. Uptake [AQM^{+}82, GA83, KLM86, PMR82, RJ85]. Upwelling [KTKD82, Lon88]. Urban [SSS86]. Urophycis [Her85]. Use [Bea81, BSN88, CGT82, CHW87, Con89, Cra84, CC84, DP85, DB86, Dol87, Fis87, GM87, HDLG88, Heg88, HH87, HSM85, Hus87, JST82, JSC88, LNK87, Mab89, MHEG89, MB85, MNT83, MRG87, PRCM84, Qua85, RC87, RTV84, Ryd82, SJ81, SSU86, SH84, SK86, TR89, Tyu87]. Used [Cha88, ED88a, EH86, HB81]. Using [BA83, HG84, Yar87]. Utah [Tyu87]. Utility [DRH88]. Utility-per-Recruit [DRH88]. Utilization [HJ83, Mil83, SJ82].

Vaccinated [Dee81, WM86]. Validation [BM83b, FC87, LN87, TY89]. Valuation [BBW87, Fed87, GM87, Hoe87, MCH87, RTP85, TC87]. Value [Fed87, Hig87, HWD88, Ran87, TL87, Van87a, Wal86a, WH82a]. Values [Dup82]. Varden [Dol86, Ell86]. Variability [KYFS87, RNT85]. Variable [CL82]. Variables [Cha88, LM85b, LNV83]. Variance [HCG89, RMC85]. Variation [BWMB88, BMW89, BC84, FGB^{+}88, GC84, Goo85b, KMTU84, KTKD82, LAK85, MM88a, MBC85, MBM89, Nie83, RPI89, SA82, TCKW83]. Variations [FS83, HBD88, Nak87, SÍW89, SN84]. Various [Dup82, Ser82a]. Vegetated [SCD81]. Vegetation [GC87, LNV83, SS82a, TSC84]. Velocity [LVNM82, Tay88]. Ventilation [CGT82]. Verification [TT82]. Vermont [FL81, LF83]. versicolor [Mac82]. versus [BCNW87, BS82, KWDS81, Ors87, PAH89]. Vertebral [Smi84]. Vertebral-Band [Smi84]. Vertical [Mat84]. Vessel [GAT82]. Viability [Hay86]. Vibrio [Dee81]. Vibriosis [WM86]. Vicinity [KHS81]. View [Ald85]. Virginia [JB87, MO88, PN87, WO86]. Virus [HM87b, MO86]. Visual [MCR84]. Vitamin [HMMJ81a, LL82]. Vitro [FCL^{+}83a, SS82b]. Volcanic [WQB82].

Walleye [Bea81, BM83c, HSG^{+}88, HPG83, KG89, LM87, MSB88, ML82, MD81, RC82, Ser86]. Walleyes [CP86, KMC84, LM82, MNT83, SK81, Ser82a]. Warmwater [AK84, DP86, GC81, RG86]. Washington [BW89, CJ85, SSS86]. Waste [MMW85]. Water [BWQ84, CLS87, CTSD84b, CVC85, CPML85, Cou85a, DGS^{+}85, DHSA83, Eri84, FB88, GCCL81, HPHF85, HZB^{+}89, HE84, HA83a, HG87, Joh83b, JST^{+}81, LNV83, MMNW81, MHEG89, MW81, MJL89, MB83, MBRK81, Pec82, PMR82, PAH89, RJ85, SWM87, SJ81, SL86, SCD81, SFS83, Tay88]. Water-Quality [LNV83]. Water-Soluble [DGS^{+}85]. Waters [Ang83, BM87, CA83, Cra84, Fab87a, GT85, HC84, HPG83, LM88, MM88b, Nak87, PHD85, SWT^{+}85]. Weight [Daw88, MM88a, McG85]. Weight-to-Length [MM88a]. Weighted [Smi85]. West [JB87, NM82, PM83, TSD81]. West-Central [NM82]. Western [BW89, DPR89, FCBTSR83, KMC84, KT87, LA84, RC82, SLJ^{+}84, Sco88, SK85, SWT^{+}85, WB81]. White [AV84, BC89, CP83, CP86, FD89, FH89, LD86, MR82, MRN83, SHM83]. Whitefish [NJS86, Rec86, TB81]. Whitney [FG89]. Whole [BQW^{+}86]. Wild [Bac84, CLL86, Han82, HER81, Hic82, Rah81, SW88a, Sos82, WS87]. Will [ATH86]. Willamette [HG87]. Winter [BO89, CHW87, CCP87, GS82a, HGP87, KMCB84]. Wires [TS84]. Wisconsin [Cob82, GC81, IM83, LM87, SK81, Ser82a, Ser82b, Ser84, Ser86, TMF83, WH82b]. Witch [Bow89]. within [Car86]. Woods [Ush87b]. Woody [AK84, BW89]. Wound [NM88]. Wyoming [CHW87, WMM81, YHW89].

X [WGR87]. X-Radiographic [WGR87].

Yakima [CJ85]. Year [Car86, CP81, DSM85, DT87, Dey81, Fos85, GV83, GSMH88, Har87, HBD88, IM83, INKP87, KE85, KBDP85, MRTR82, MS82, MS84a, MG88, NG86, Pec82, Ser82b, Ser86, WJ87]. Year-0 [KE85]. Year-Class [Car86, CP81, KBDP85, Ser86]. Yearling [JH83, RS83].

Yearlings [HZB^{+}89]. Yearly [BK85]. Years [BNW88, Wat83]. Yellow [AAC88, BMM^{+}87, FS83, FH89, HW81, HM87a, HR87, Hin85, KMC84, LB82, MJL89, MF81, MF83, MCR84, MCK86, MPSC89, Nie83]. Yellowtail [Car86, Fra81, Joh83a, LN87]. Yield [Bay81, Cou81, ED88b, FI88, Har86, MH84, MI89, RF88, SÍW89]. Yield-per-Recruit [ED88b, MH84, RF88]. Yields [Bay81, LMW81, SR82]. Yolk [HPG83, JB87, KMC87]. Yolk-Sac [HPG83, JB87, KMC87]. York [SP88, HW81, MF81, MF83, WO86]. Young [AMCD85, BS82, DSM85, DP85, DT87, Dey81, DRRP82, Fos85, GV83, GS82a, GC84, Goo85a, GC87, Har87, Hau86, HHB81b, HBD88, HKA83, Kay89, KG83, MRTR82, MS82, MS84a, MF81, MF83, MCR84, MCK86, MZM89, MG88, PHD85, Pec82, Pow82, SL86, SMLM83, TM85b]. Young-of-the-Year [DSM85, DT87, Dey81, Fos85, Har87, HBD88, MRTR82, MS82, MG88, Pec82]. Young-of-Year [MS84a]. Yukon [BMW89].

Zealand [GD85b]. Zebra [Fui86]. Zinc [FV82, Rah81]. Zone [LM87]. Zones [SK85]. Zooplanktivory [HS89]. Zooplankton [KN81, MMNW81, MCK86, TB81, WOL83].

References

Al-Absy:1988:NCS

[AAC88] Ahmad H. Al-Absy and Kenneth D. Carlander. Notes: Criteria for selection of scale-sampling sites in growth studies of yellow perch. Transactions of the American Fisheries Society, 117(2): 209–212, March 1988. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Armstrong:1982:OTC

[AB82] Martin S. Armstrong and Claude E. Boyd. Oxygen transfer calculations for a tractor-powered paddlewheel aerator. Transactions of the American Fisheries Society, 111(3):361–366, May 1982. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Armstrong:1983:DDF

[AB83] Michael L. Armstrong and Arthur V. Brown. Diel drift and feeding of channel catfish alevins in the Illinois River, Arkansas. Transactions of the American Fisheries Society, 112(2):302–307, March 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Amend:1982:TFC

[ACG^{+}82] Donald F. Amend, Thomas R. Croy, Berverly A. Goven, Keith A. Johnson, and Donald H. McCarthy. Transportation of fish in closed systems: Methods to control ammonia, carbon dioxide, pH, and bacterial growth. Transactions of the American Fisheries Society, 111(5):603–611, September 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Amend:1982:EED

[AGE82] Donald F. Amend, Beverly A. Goven, and Diane G. Elliot. Etomidate: Effective dosages for a new fish anesthetic. Transactions of the American Fisheries Society, 111(3):337–341, May 1982. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Angermeier:1984:RBW

[AK84] Paul L. Angermeier and James R. Karr. Relationships between woody debris and fish habitat in a small warmwater stream. Transactions of the American Fisheries Society, 113(6):716–726, November 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Alderdice:1985:PVE

[Ald85] D. F. Alderdice. A pragmatic view of early life history studies of fishes. Transactions of the American Fisheries Society, 114(4): 445–451, July 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ammerman:1989:BGE

[AM89] Loren K. Ammerman and Donald C. Morizot. Biochemical genetics of endangered Colorado squawfish populations. Transactions of the American Fisheries Society, 118(4):435–440, July 1989. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Akielaszek:1985:ECY

[AMCD85] John J. Akielaszek, John R. Moring, Samuel R. Chapman, and John H. Dearborn. Experimental culture of young rainbow smelt Osmerus mordax. Transactions of the American Fisheries Society, 114(4):596–603, July 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Adams:1982:EPL

[AMP82] S. Marshall Adams, R. B. McLean, and J. A. Parrotta. Energy partitioning in largemouth bass under conditions of seasonally fluctuating prey availability. Transactions of the American Fisheries Society, 111(5):549–558, September 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anderson:1987:EFM

[And87] Lee G. Anderson. Expansion of the fisheries management paradigm to include institutional structure and function. Transactions of the American Fisheries Society, 116(3):396–404, May 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Angus:1983:PTP

[Ang83] Robert A. Angus. Phenol tolerance in populations of mosquitofish from polluted and nonpolluted waters. Transactions of the American Fisheries Society, 112(6):794–799, November 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1981:BRa

[Ano81a] Anonymous. Book reviews. Transactions of the American Fisheries Society, 110(1):188–190, January 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1981:BRb

[Ano81b] Anonymous. Book reviews. Transactions of the American Fisheries Society, 110(3):465–468, May 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1981:BRc

[Ano81c] Anonymous. Book reviews. Transactions of the American Fisheries Society, 110(4):569–572, July 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1981:BRd

[Ano81d] Anonymous. Book reviews. Transactions of the American Fisheries Society, 110(5):665–668, September 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1981:BRe

[Ano81e] Anonymous. Book reviews. Transactions of the American Fisheries Society, 110(6):783–787, November 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1981:E

[Ano81f] Anonymous. Errata. Transactions of the American Fisheries Society, 110(6):787, November 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1982:BRa

[Ano82a] Anonymous. Book reviews. Transactions of the American Fisheries Society, 111(1):110–113, January 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1982:BRb

[Ano82b] Anonymous. Book reviews. Transactions of the American Fisheries Society, 111(2):250–253, March 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1982:BRc

[Ano82c] Anonymous. Book reviews. Transactions of the American Fisheries Society, 111(3):409–411, May 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1982:BRd

[Ano82d] Anonymous. Book reviews. Transactions of the American Fisheries Society, 111(4):538–542, July 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1982:BRe

[Ano82e] Anonymous. Book reviews. Transactions of the American Fisheries Society, 111(5):656–663, September 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1982:BRf

[Ano82f] Anonymous. Book reviews. Transactions of the American Fisheries Society, 111(6):787–792, November 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1983:BRa

[Ano83a] Anonymous. Book reviews. Transactions of the American Fisheries Society, 112(1):123–128, January 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1983:BRb

[Ano83b] Anonymous. Book reviews. Transactions of the American Fisheries Society, 112(3):452–456, May 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1983:BRc

[Ano83c] Anonymous. Book reviews. Transactions of the American Fisheries Society, 112(4):570–575, July 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1983:BRd

[Ano83d] Anonymous. Book reviews. Transactions of the American Fisheries Society, 112(5):729–733, September 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1983:BRe

[Ano83e] Anonymous. Book reviews. Transactions of the American Fisheries Society, 112(6):838–844, November 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1984:BRa

[Ano84a] Anonymous. Book reviews. Transactions of the American Fisheries Society, 113(1):95–100, January 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1984:BRb

[Ano84b] Anonymous. Book reviews. Transactions of the American Fisheries Society, 113(2):260–262, March 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1984:BRc

[Ano84c] Anonymous. Book reviews. Transactions of the American Fisheries Society, 113(3):407–409, May 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1984:BRd

[Ano84d] Anonymous. Book reviews. Transactions of the American Fisheries Society, 113(4):553–556, July 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1984:BRe

[Ano84e] Anonymous. Book reviews. Transactions of the American Fisheries Society, 113(5):672–675, September 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1984:BRf

[Ano84f] Anonymous. Book reviews. Transactions of the American Fisheries Society, 113(6):805–806, November 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1984:Ea

[Ano84g] Anonymous. Errata. Transactions of the American Fisheries Society, 113(2):263, March 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1984:Eb

[Ano84h] Anonymous. Erratum. Transactions of the American Fisheries Society, 113(3):410, May 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1985:BRa

[Ano85a] Anonymous. Book reviews. Transactions of the American Fisheries Society, 114(1):152–157, January 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1985:BRb

[Ano85b] Anonymous. Book reviews. Transactions of the American Fisheries Society, 114(2):310–316, March 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1985:BRc

[Ano85c] Anonymous. Book reviews. Transactions of the American Fisheries Society, 114(3):441–442, May 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1985:BRd

[Ano85d] Anonymous. Book reviews. Transactions of the American Fisheries Society, 114(4):622–629, July 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1985:BRe

[Ano85e] Anonymous. Book reviews. Transactions of the American Fisheries Society, 114(5):768–770, September 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1985:BRf

[Ano85f] Anonymous. Book reviews. Transactions of the American Fisheries Society, 114(6):914–918, November 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1985:E

[Ano85g] Anonymous. Errata. Transactions of the American Fisheries Society, 114(5):770, September 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1986:BRa

[Ano86a] Anonymous. Book reviews. Transactions of the American Fisheries Society, 115(1):172–175, January 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1986:BRb

[Ano86b] Anonymous. Book reviews. Transactions of the American Fisheries Society, 115(2):359–362, March 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1986:BRc

[Ano86c] Anonymous. Book reviews. Transactions of the American Fisheries Society, 115(4):641–642, July 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1986:BRd

[Ano86d] Anonymous. Book reviews. Transactions of the American Fisheries Society, 115(5):784–786, September 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1986:BRe

[Ano86e] Anonymous. Book reviews. Transactions of the American Fisheries Society, 115(6):918–922, November 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1986:Ea

[Ano86f] Anonymous. Errata. Transactions of the American Fisheries Society, 115(1):175, January 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1986:Eb

[Ano86g] Anonymous. Errata. Transactions of the American Fisheries Society, 115(3):499, May 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1987:BRa

[Ano87a] Anonymous. Book reviews. Transactions of the American Fisheries Society, 116(1):123–132, January 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1987:BRb

[Ano87b] Anonymous. Book reviews. Transactions of the American Fisheries Society, 116(2):286–288, March 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1987:BRc

[Ano87c] Anonymous. Book reviews. Transactions of the American Fisheries Society, 116(4):674–676, July 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1987:BRd

[Ano87d] Anonymous. Book reviews. Transactions of the American Fisheries Society, 116(6):891–893, November 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1987:Ea

[Ano87e] Anonymous. Errata. Transactions of the American Fisheries Society, 116(5):778, September 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1987:Eb

[Ano87f] Anonymous. Errata. Transactions of the American Fisheries Society, 116(6):893, November 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Alexander:1982:DBU

[AQM^{+}82] H. C. Alexander, J. A. Quick, Jr., R. J. Moolenaar, E. A. Bartlett, C. G. Mendoza, and G. U. Boggs. Dynamics of bromide uptake and depuration by fathead minnows. Transactions of the American Fisheries Society, 111(6):755–760, November 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Acheson:1982:BEE

[AR82] James M. Acheson and Robert Reidman. Biological and economic effects of increasing the minimum legal size of American lobster in Maine. Transactions of the American Fisheries Society, 111(1): 1–12, January 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Argyle:1982:ARS

[Arg82] Ray L. Argyle. Alewives and rainbow smelt in Lake Huron: Midwater and bottom aggregations and estimates of standing stocks. Transactions of the American Fisheries Society, 111(3):267–285, May 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Allen:1983:PHG

[AS83] Standish K. Allen, Jr. and Jon G. Stanley. Ploidy of hybrid grass carp × bighead carp determined by flow cytometry. Transactions of the American Fisheries Society, 112(3):431–435, May 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Allen:1986:CEL

[ATH86] Standish K. Allen, Jr., Richard G. Thiery, and Neal T. Hagstrom. Cytological evaluation of the likelihood that triploid grass carp will reproduce. Transactions of the American Fisheries Society, 115(6):841–848, November 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Avise:1984:GAR

[AV84] John C. Avise and Michael J. Van Den Avyle. Genetic analysis of reproduction of hybrid white bass × striped bass in the Savannah River. Transactions of the American Fisheries Society, 113(5): 563–570, September 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Axon:1985:SBM

[AW85] James R. Axon and David K. Whitehurst. Striped bass management in lakes with emphasis on management problems. Transactions of the American Fisheries Society, 114(1):8–11, January 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bannerot:1983:UFD

[BA83] Scott P. Bannerot and C. Bruce Austin. Using frequency distributions of catch per unit effort to measure fish-stock abundance. Transactions of the American Fisheries Society, 112(5): 608–617, September 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Boreman:1985:PHA

[BA85] John Boreman and Herbert M. Austin. Production and harvest of anadromous striped bass stocks along the Atlantic Coast. Transactions of the American Fisheries Society, 114(1):3–7, January 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bachman:1984:FBF

[Bac84] Robert A. Bachman. Foraging behavior of free-ranging wild and hatchery brown trout in a stream. Transactions of the American Fisheries Society, 113(1):1–32, January 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Balon:1984:RSD

[Bal84] Eugene K. Balon. Reflections on some decisive events in the early life of fishes. Transactions of the American Fisheries Society, 113 (2):178–185, March 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bartell:1982:IPA

[Bar82] Steven M. Bartell. Influence of prey abundance on size-selective predation by bluegills. Transactions of the American Fisheries Society, 111(4):453–461, July 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Barger:1985:AGA

[Bar85] Lyman E. Barger. Age and growth of Atlantic croakers in the Northern Gulf of Mexico, based on otolith sections. Transactions of the American Fisheries Society, 114(6):847–850, November 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bayley:1981:FYA

[Bay81] Peter B. Bayley. Fish yield from the Amazon in Brazil: Comparison with African River yields and management possibilities. Transactions of the American Fisheries Society, 110(3):351–359, May 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brown:1982:FCC

[BB82] Steven W. Brown and Claude E. Boyd. Off-flavor in channel catfish from commercial ponds. Transactions of the American Fisheries Society, 111(3):379–383, May 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beck:1983:PHB

[BB83] Melvin L. Beck and Charles J. Biggers. Ploidy of hybrids between grass carp and bighead carp determined by morphological analysis. Transactions of the American Fisheries Society, 112(6): 808–811, November 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brooks:1984:EFC

[BB84] Arthur S. Brooks and James M. Bartos. Effects of free and combined chlorine and exposure duration on rainbow trout, channel catfish, and emerald shiners. Transactions of the American Fisheries Society, 113(6):786–793, November 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bishop:1987:TTE

[BBW87] Richard C. Bishop, Kevin J. Boyle, and Michael P. Welsh. Toward total economic valuation of great lakes fishery resources. Transactions of the American Fisheries Society, 116(3):339–345, May 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bosley:1984:GTV

[BC84] Timothy R. Bosley and John V. Conner. Geographic and temporal variation in numbers of myomeres in fish larvae from the Lower Mississippi River. Transactions of the American Fisheries Society, 113(2):238–242, March 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brennan:1989:CAD

[BC89] James S. Brennan and Gregor M. Cailliet. Comparative agedetermination techniques for white sturgeon in California. Transactions of the American Fisheries Society, 118(3):296–310, May 1989. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Belanger:1987:DFV

[BCNW87] Scott E. Belanger, Donald S. Cherry, John J. Ney, and David K. Whitehurst. Differentiation of freshwater versus saltwater striped bass by elemental scale analysis. Transactions of the American Fisheries Society, 116(4):594–600, July 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Buckman:1982:RBS

[BE82] M. Buckman and R. D. Ewing. Relationship between size and time of entry into the sea and gill (Na+K)-ATPase activity for juvenile spring Chinook salmon. Transactions of the American Fisheries Society, 111(6):681–687, November 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beamish:1981:UFR

[Bea81] Richard J. Beamish. Use of fin-ray sections to age walleye pollock, Pacific cod, and albacore, and the importance of this method. Transactions of the American Fisheries Society, 110(2):287–299, March 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beacham:1984:AMC

[Bea84] Terry D. Beacham. Age and morphology of chum salmon in Southern British Columbia. Transactions of the American Fisheries Society, 113(6):727–736, November 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bedi:1987:PRF

[Bed87] Nilam Bedi. Pricing of recreational fishing access — a discussion of major issues, with special reference to Ontario. Transactions of the American Fisheries Society, 116(3):390–395, May 1987. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Boreman:1981:EME

[BGC81] John Boreman, C. Phillip Goodyear, and Sigurd W. Christensen. An empirical methodology for estimating entrainment losses at power plants sited on estuaries. Transactions of the American Fisheries Society, 110(2):253–260, March 1981. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bain:1983:RMP

[BH83] Mark B. Bain and Louis A. Helfrich. Role of male parental care in survival of larval bluegills. Transactions of the American Fisheries Society, 112(1):47–52, January 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Buck:1986:CGN

[BH86] D. Homer Buck and Michael L. Hooe. Comparative growth of northern largemouth bass and F_{1} hybrid largemouth bass through three growing seasons. Transactions of the American Fisheries Society, 115(2):296–304, March 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brown:1988:CSF

[BH88] Thomas G. Brown and Gordon F. Hartman. Contribution of seasonally flooded lands and minor tributaries to the production of Coho salmon in Carnation Creek, British Columbia. Transactions of the American Fisheries Society, 117(6):546–551, November 1988. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Buckley:1985:CSS

[BHL^{+}85] Lawrence J. Buckley, Thomas A. Halavik, Geoffrey C. Laurence, Steven J. Hamilton, and Paul Yevich. Comparative swimming stamina, biochemical composition, backbone mechanical properties, and histopathology of juvenile striped bass from rivers and hatcheries of the Eastern United States. Transactions of the American Fisheries Society, 114(1):114–124, January 1985. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bryan:1984:IAT

[BHN84] James D. Bryan, Loren G. Hill, and William H. Neill. Interdependence of acute temperature preference and respiration in the plains minnow. Transactions of the American Fisheries Society, 113(5):557–562, September 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bozeman:1985:PSH

[BHR85] Earl L. Bozeman, Gene S. Helfman, and Thelma Richardson. Population size and home range of American eels in a Georgia tidal creek. Transactions of the American Fisheries Society, 114(6): 821–825, November 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brown:1987:IEF

[BJCM87] Joseph A. Brown, Peter H. Johansen, Patrick W. Colgan, and R. Alastair Mathers. Impairment of early feeding behavior of largemouth bass by pentachlorophenol exposure: a preliminary assessment. Transactions of the American Fisheries Society, 116 (1):71–78, January 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Buckley:1985:YMS

[BK85] Jack Buckley and Boyd Kynard. Yearly movements of shortnose sturgeons in the Connecticut River. Transactions of the American Fisheries Society, 114(6):813–820, November 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Berrill:1983:LCF

[BL83a] Michael Berrill and David C. Lasenby. Life cycles of the freshwater mysid shrimp Mysis relicta reared at two temperatures. Transactions of the American Fisheries Society, 112(4):551–553, July 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Black:1983:CSD

[BL83b] E. A. Black and C. J. Low. Ctenophores in salmon diets. Transactions of the American Fisheries Society, 112(5):728, September 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Baragi:1986:DEA

[BL86] V. Baragi and Richard T. Lovell. Digestive enzyme activities in striped bass from first feeding through larva development. Transactions of the American Fisheries Society, 115(3):478–484, May 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bayley:1989:ESF

[BLD89] Peter B. Bayley, R. Weldon Larimore, and David C. Dowling. Electric seine as a fish-sampling gear in streams. Transactions of the American Fisheries Society, 118(4):447–453, July 1989. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bonin:1981:MTL

[BLR81] John D. Bonin, Raymond M. Lee, and John N. Rinne. Measuring thermal limits of fish. Transactions of the American Fisheries Society, 110(5):662–664, September 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Barwick:1983:AGR

[BM83a] D. Hugh Barwick and Philip R. Moore. Abundance and growth of redeye bass in two South Carolina reservoirs. Transactions of the American Fisheries Society, 112(2):216–219, March 1983. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beamish:1983:FRA

[BM83b] R. J. Beamish and G. A. McFarlane. The forgotten requirement for age validation in fisheries biology. Transactions of the American Fisheries Society, 112(6):735–743, November 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bulkowski:1983:CPD

[BM83c] Loretta Bulkowski and James W. Meade. Changes in phototaxis during early development of walleye. Transactions of the American Fisheries Society, 112(3):445–447, May 1983. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Barber:1987:ETT

[BM87] Willard E. Barber and Gordon A. McFarlane. Evaluation of three techniques to age Arctic char from Alaskan and Canadian waters. Transactions of the American Fisheries Society, 116(6):874–881, November 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beamish:1988:ADL

[BM88] F. W. H. Beamish and T. E. Medland. Age determination for lampreys. Transactions of the American Fisheries Society, 117 (1):63–71, January 1988. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Buckler:1981:ACE

[BMHW81] Denny R. Buckler, Foster L. Mayer, James N. Huckins, and Arthur Witt, Jr. Acute and chronic effects of Kepone and mirex on the fathead minnow. Transactions of the American Fisheries Society, 110(2):270–280, March 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brandt:1987:PAL

[BMM^{+}87] Stephen B. Brandt, Doran M. Mason, David B. Macneill, Thomas Coates, and John E. Gannon. Predation by alewives on larvae of yellow perch in Lake Ontario. Transactions of the American Fisheries Society, 116(4):641–645, July 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Buth:1987:PDI

[BMU87] Donald G. Buth, Robert W. Murphy, and Linda Ulmer. Population differentiation and introgressive hybridization of the flannelmouth sucker and of hatchery and native stocks of the razorback sucker. Transactions of the American Fisheries Society, 116 (1):103–110, January 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beacham:1989:AMB

[BMW89] Terry D. Beacham, Clyde B. Murray, and Ruth E. Withler. Age, morphology, and biochemical genetic variation of Yukon River Chinook salmon. Transactions of the American Fisheries Society, 118(1):46–63, January 1989. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Becker:1982:EDC

[BNF82] C. Dale Becker, Duane A. Neitzel, and Duane H. Fickeisen. Effects of dewatering on Chinook salmon redds: Tolerance of four developmental phases to daily dewaterings. Transactions of the American Fisheries Society, 111(5):624–637, September 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bisson:1988:SPC

[BNW88] Peter A. Bisson, Jennifer L. Nielsen, and James W. Ward. Summer production of Coho salmon stocked in Mount St. Helens streams 3–6 years after the 1980 eruption. Transactions of the American Fisheries Society, 117(4):322–335, July 1988. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bergstedt:1989:DAS

[BO89] Roger A. Bergstedt and Robert O’Gorman. Distribution of alewives in Southeastern Lake Ontario in autumn and winter: a clue to winter mortalities. Transactions of the American Fisheries Society, 118(6):687–692, November 1989. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bodammer:1985:CDL

[Bod85] Joel E. Bodammer. Corneal damage in larvae of striped bass Morone saxatilis exposed to copper. Transactions of the American Fisheries Society, 114(4):577–583, July 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Boreman:1983:SSB

[Bor83] John Boreman. Simulation of striped bass egg and larva development based on temperature. Transactions of the American Fisheries Society, 112(2):286–292, March 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bouck:1982:GID

[Bou82] Gerald R. Bouck. Gasometer: an inexpensive device for continuous monitoring of dissolved gases and supersaturation. Transactions of the American Fisheries Society, 111(4):505–516, July 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bowen:1981:DAP

[Bow81] Stephen H. Bowen. Digestion and assimilation of periphytic detrital aggregate by Tilapia mossambica. Transactions of the American Fisheries Society, 110(2):239–245, March 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bowering:1989:WFD

[Bow89] W. R. Bowering. Witch flounder distribution off Southern Newfoundland, and changes in age, growth, and sexual maturity patterns with commercial exploitation. Transactions of the American Fisheries Society, 118(6):659–669, November 1989. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bishop:1983:CBC

[BOWY83] James M. Bishop, Eugene J. Olmi III, J. David Whitaker, and G. Michael Yianopoulos. Capture of blue crab peelers in South Carolina: an analysis of techniques. Transactions of the American Fisheries Society, 112(1):60–70, January 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Boyd:1981:CFF

[Boy81a] Claude E. Boyd. Comparison of five fertilization programs for fish ponds. Transactions of the American Fisheries Society, 110(4): 541–545, July 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Boyd:1981:SGI

[Boy81b] Claude E. Boyd. Solubility of granular inorganic fertilizers for fish ponds. Transactions of the American Fisheries Society, 110(3): 451–454, May 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Boyd:1982:HSE

[Boy82] Claude E. Boyd. Hydrology of small experimental fish ponds at Auburn, Alabama. Transactions of the American Fisheries Society, 111(5):638–644, September 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bishop:1984:EPP

[BOY84] James M. Bishop, Eugene J. Olmi III, and G. Michael Yianopoulos. Efficacy of peeler pots and experimental habitat, pots for capture of premolt blue crabs. Transactions of the American Fisheries Society, 113(5):642–654, September 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Boyd:1985:CBC

[Boy85a] Claude E. Boyd. Chemical budgets for channel catfish ponds. Transactions of the American Fisheries Society, 114(2):291–298, March 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Boyd:1985:PE

[Boy85b] Claude E. Boyd. Pond evaporation. Transactions of the American Fisheries Society, 114(2):299–303, March 1985. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bulkley:1983:TPA

[BP83] R. V. Bulkley and R. Pimentel. Temperature preference and avoidance by adult razorback suckers. Transactions of the American Fisheries Society, 112(5):601–607, September 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bayley:1984:PMG

[BP84] Peter B. Bayley and Daniel Pauly. Prerecruit mortality in Gulf of Thailand shrimp stocks. Transactions of the American Fisheries Society, 113(3):403–406, May 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Berry:1985:SPT

[BP85] Charles R. Berry, Jr. and Richard Pimentel. Swimming performances of three rare Colorado River fishes. Transactions of the American Fisheries Society, 114(3):397–402, May 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Berkes:1987:QMP

[BP87] Fikret Berkes and Dorothy Pocock. Quota management and “people problems”: a case history of Canadian Lake Erie fisheries. Transactions of the American Fisheries Society, 116(3):494–502, May 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brannon:1986:HAC

[BQW^{+}86] E. L. Brannon, T. P. Quinn, R. P. Whitman, A. E. Nevissi, R. E. Nakatani, and Clayton D. McAuliffe. Homing of adult Chinook salmon after brief exposure to whole and dispersed crude oil. Transactions of the American Fisheries Society, 115(6):823– 827, November 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bowlby:1986:TBH

[BR86] James N. Bowlby and John C. Roff. Trout biomass and habitat relationships in Southern Ontario streams. Transactions of the American Fisheries Society, 115(4):503–514, July 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brandt:1986:OSH

[Bra86] Stephen B. Brandt. Ontogenetic shifts in habitat, diet, and dielfeeding periodicity of slimy sculpin in Lake Ontario. Transactions of the American Fisheries Society, 115(5):711–715, September 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Breitburg:1988:ETP

[Bre88] Denise L. Breitburg. Effects of turbidity on prey consumption by striped bass larvae. Transactions of the American Fisheries Society, 117(1):72–77, January 1988. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brownell:1985:LAC

[Bro85] Charles L. Brownell. Laboratory analysis of cannibalism by larvae of the Cape anchovy Engraulis capensis. Transactions of the American Fisheries Society, 114(4):512–518, July 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brown:1986:PDG

[Bro86] Karen L. Brown. Population demographic and genetic structure of plains killifish from the Kansas and Arkansas River basins in Kansas. Transactions of the American Fisheries Society, 115(4): 568–576, July 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brown:1987:THD

[Bro87] Tommy L. Brown. Typology of human dimensions information needed for great lakes sport-fisheries management. Transactions of the American Fisheries Society, 116(3):320–324, May 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Broussard:1981:ERC

[BS81] Meryl C. Broussard, Jr. and Robert R. Stickney. Evaluation of reproductive characters for four strains of channel catfish. Transactions of the American Fisheries Society, 110(4):502–506, July 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bry:1982:PYN

[BS82] C. Bry and Y. Souchon. Production of young northern pike families in small ponds:natural spawning versus fry stocking. Transactions of the American Fisheries Society, 111(4):476–480, July 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Barton:1987:MCA

[BS87] Bruce A. Barton and Carl B. Schreck. Metabolic cost of acute physical stress in juvenile steelhead. Transactions of the American Fisheries Society, 116(2):257–263, March 1987. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Benfey:1986:FCC

[BSDD86] Tillmann J. Benfey, Igor I. Solar, Gary De Jong, and Edward M. Donaldson. Flow-cytometric confirmation of aneuploidy in sperm from triploid rainbow trout. Transactions of the American Fisheries Society, 115(6):838–840, November 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Berra:1982:PIC

[BSM82] Tim M. Berra, J. F. Smith, and J. D. Morrison. Probable identification of the cucumber odor of the Australian grayling Prototroctes maraena. Transactions of the American Fisheries Society, 111(1):78–82, January 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bisson:1988:CHH

[BSN88] Peter A. Bisson, Kathleen Sullivan, and Jennifer L. Nielsen. Channel hydraulics, habitat use, and body form of juvenile Coho salmon, steelhead, and cutthroat trout in streams. Transactions of the American Fisheries Society, 117(3):262–273, May 1988. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Baumann:1987:TFC

[BSP87] Paul C. Baumann, W. David Smith, and William K. Parland. Tumor frequencies and contaminant concentrations in brown bullheads from an industrialized river and a recreational lake. Transactions of the American Fisheries Society, 116(1):79–86, January 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Barton:1986:MAD

[BSS86] Bruce A. Barton, Carl B. Schreck, and Linda A. Sigismondi. Multiple acute disturbances evoke cumulative physiological stress responses in juvenile Chinook salmon. Transactions of the American Fisheries Society, 115(2):245–251, March 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Breeser:1988:OMH

[BSS^{+}88] Steven W. Breeser, F. David Stearns, Michael W. Smith, Robin L. West, and James B. Reynolds. Observations of movements and habitat preferences of burbot in an Alaskan glacial river system. Transactions of the American Fisheries Society, 117(5):506–509, September 1988. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beauchamp:1989:CBM

[BST89] David A. Beauchamp, Donald J. Stewart, and G. L. Thomas. Corroboration of a bioenergetics model for sockeye salmon. Transactions of the American Fisheries Society, 118(6):597–607, November 1989. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Boney:1984:SRB

[BSYW84] Stephen E. Boney, William L. Shelton, Shiu-Liu Yang, and L. O. Wilken. Sex reversal and breeding of grass carp. Transactions of the American Fisheries Society, 113(3):348–353, May 1984. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Buchanan:1983:RSF

[BSZF83] D. V. Buchanan, J. E. Sanders, J. L. Zinn, and J. L. Fryer. Relative susceptibility of four strains of summer steelhead to infection by Ceratomyxa shasta. Transactions of the American Fisheries Society, 112(4):541–543, July 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Baltz:1987:ITM

[BVBM87] Donald M. Baltz, Bruce Vondracek, Larry R. Brown, and Peter B. Moyle. Influence of temperature on microhabitat choice by fishes in a California stream. Transactions of the American Fisheries Society, 116(1):12–20, January 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bilby:1989:CCF

[BW89] Robert E. Bilby and James W. Ward. Changes in characteristics and function of woody debris with increasing size of streams in Western Washington. Transactions of the American Fisheries Society, 118(4):368–378, July 1989. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beacham:1988:VBS

[BWMB88] Terry D. Beacham, Ruth E. Withler, Clyde B. Murray, and L. W. Barner. Variation in body size, morphology, egg size, and biochemical genetics of pink salmon in British Columbia. Transactions of the American Fisheries Society, 117(2):109–126, March 1988. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brannon:1984:RRA

[BWQ84] E. L. Brannon, R. P. Whitman, and T. P. Quinn. Responses of returning adult Coho salmon to home water and populationspecific odors. Transactions of the American Fisheries Society, 113(3):374–377, May 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Boynton:1981:IJS

[BZP81] Walter R. Boynton, H. Howard Zion, and Tibor T. Polgar. Importance of juvenile striped bass food habits in the Potomac estuary. Transactions of the American Fisheries Society, 110(1): 56–63, January 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Colton:1983:FEB

[CA83] Douglas E. Colton and William S. Alevizon. Feeding ecology of bonefish in Bahamian waters. Transactions of the American Fisheries Society, 112(2):178–184, March 1983. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Campana:1984:CAD

[Cam84] Steven E. Campana. Comparison of age determination methods for the starry flounder. Transactions of the American Fisheries Society, 113(3):365–369, May 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Carlander:1982:SIC

[Car82] Kenneth D. Carlander. Standard intercepts for calculating lengths from scale measurements for some centrarchid and percid fishes. Transactions of the American Fisheries Society, 111(3):332–336, May 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Carmichael:1983:SND

[Car83] Gary J. Carmichael. Scale-number differences of central stonerollers incubated and reared at different temperatures. Transactions of the American Fisheries Society, 112(3):441–444, May 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Carlson:1986:RYC

[Car86] H. Richard Carlson. Restricted year-class structure and recruitment lag within a discrete school of yellowtail rockfish. Transactions of the American Fisheries Society, 115(3):490–492, May 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Carline:1987:SMB

[Car87] Robert F. Carline. Simplified method based on bioenergetics modeling to estimate food consumption by largemouth bass and northern pike. Transactions of the American Fisheries Society, 116(2):224–231, March 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Crecco:1983:FEC

[CB83] Victor A. Crecco and Mark M. Blake. Feeding ecology of coexisting larvae of American shad and blueback herring in the Connecticut River. Transactions of the American Fisheries Society, 112(4):498–507, July 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Chandler:1988:AGI

[CB88a] G. L. Chandler and T. C. Bjornn. Abundance, growth, and interactions of juvenile steelhead relative to time of emergence. Transactions of the American Fisheries Society, 117(5):432–443, September 1988. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cook:1988:MHS

[CB88b] Mark F. Cook and Eric P. Bergersen. Movements, habitat selection, and activity periods of northern pike in Eleven Mile Reservoir, Colorado. Transactions of the American Fisheries Society, 117(5):495–502, September 1988. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Chandler:1989:SGS

[CB89] G. L. Chandler and T. C. Bjornn. Survival and growth of subyearling steelhead homozygous for alternative alleles at the dipeptidase locus. Transactions of the American Fisheries Society, 118 (1):74–77, January 1989. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cox:1981:GDJ

[CC81] David K. Cox and Charles C. Coutant. Growth dynamics of juvenile striped bass as functions of temperature and ration. Transactions of the American Fisheries Society, 110(2):226–238, March 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Crowder:1984:ESR

[CC84] Larry B. Crowder and Helen L. Crawford. Ecological shifts in resource use by bloaters in Lake Michigan. Transactions of the American Fisheries Society, 113(6):694–700, November 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Carothers:1985:RBT

[CC85] Paul E. Carothers and Mark E. Chittenden, Jr. Relationships between trawl catch and tow duration for penaeid shrimp. Transactions of the American Fisheries Society, 114(6):851–856, November 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Crouse:1981:EFS

[CCMD81] M. R. Crouse, C. A. Callahan, K. W. Malueg, and S. E. Dominguez. Effects of fine sediments on growth of juvenile Coho salmon in laboratory streams. Transactions of the American Fisheries Society, 110(2):281–286, March 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cunjak:1987:SEB

[CCP87] Richard A. Cunjak, Allen Curry, and Geoffrey Power. Seasonal energy budget of brook trout in streams: Implications of a possible deficit in early winter. Transactions of the American Fisheries Society, 116(6):817–828, November 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Coutant:1983:CTD

[CD83] Charles C. Coutant and Donald L. Deangelis. Comparative temperature-dependent growth rates of largemouth and smallmouth bass fry. Transactions of the American Fisheries Society, 112(3):416–423, May 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Chairman:1983:SLF

[CE83] F. Douglas Martin Conference Chairman and Editor. The sixth larval fish conference. Transactions of the American Fisheries Society, 112(2):233–234, March 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cloud:1988:NSC

[CEG88] J. G. Cloud, A. L. Erdahl, and E. F. Graham. Notes: Survival and continued normal development of fish embryos after incubation at reduced temperatures. Transactions of the American Fisheries Society, 117(5):503–506, September 1988. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Chavance:1984:ELH

[CFCSI84] P. Chavance, C. Flores-Coto, and A. Sanchez-Iturbe. Early life history and adult biomass of sea bream in the terminos lagoon, Southern Gulf of Mexico. Transactions of the American Fisheries Society, 113(2):166–177, March 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cunjak:1984:SDB

[CG84] Richard A. Cunjak and John M. Green. Species dominance by brook trout and rainbow trout in a simulated stream environment. Transactions of the American Fisheries Society, 113(6):737–743, November 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cairns:1982:UFV

[CGT82] Michael A. Cairns, Ronald R. Garton, and Richard A. Tubb. Use of fish ventilation frequency to estimate chronically safe toxicant concentrations. Transactions of the American Fisheries Society, 111(1):70–77, January 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Chew:1985:TCE

[CH85a] Robert D. Chew and Martin A. Hamilton. Toxicity curve estimation: Fitting a compartment model to median survival times. Transactions of the American Fisheries Society, 114(3):403–412, May 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Chisholm:1985:EDT

[CH85b] Ian M. Chisholm and Wayne A. Hubert. Expulsion of dummy transmitters by rainbow trout. Transactions of the American Fisheries Society, 114(5):766–767, September 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cipriano:1986:SSF

[CH86] Rocco C. Cipriano and Charles M. Heartwell III. Susceptibility of salmonids to furunculosis: Differences between serum and mucus responses against Aeromonas salmonicida. Transactions of the American Fisheries Society, 115(1):83–88, January 1986. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Chapman:1986:SSA

[Cha86] D. W. Chapman. Salmon and steelhead abundance in the Columbia River in the Nineteenth Century. Transactions of the American Fisheries Society, 115(5):662–670, September 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Chapman:1988:CRV

[Cha88] D. W. Chapman. Critical review of variables used to define effects of fines in redds of large salmonids. Transactions of the American Fisheries Society, 117(1):1–21, January 1988. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Chisholm:1987:WSC

[CHW87] Ian M. Chisholm, Wayne A. Hubert, and Thomas A. Wesche. Winter stream conditions and use of habitat by brook trout in high-elevation Wyoming streams. Transactions of the American Fisheries Society, 116(2):176–184, March 1987. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cipriano:1983:RSA

[Cip83] Rocco C. Cipriano. Resistance of salmonids to Aeromonas salmonicida: Relation between agglutinins and neutralizing activities. Transactions of the American Fisheries Society, 112(1): 95–99, January 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Campton:1985:EEG

[CJ85] Donald E. Campton and James M. Johnston. Electrophoretic evidence for a genetic admixture of native and nonnative rainbow trout in the Yakima River, Washington. Transactions of the American Fisheries Society, 114(6):782–793, November 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Clady:1982:CBS

[CL82] Michael D. Clady and George W. Luker. Competition between spotted and largemouth bass in variable pond ecosystems. Transactions of the American Fisheries Society, 111(6):665–674, November 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cleveland:1986:ITA

[CLH^{+}86] Laverne Cleveland, Edward E. Little, Steven J. Hamilton, Denny R. Buckler, and Joseph B. Hunn. Interactive toxicity of aluminum and acidity to early life stages of brook trout. Transactions of the American Fisheries Society, 115(4):610–620, July 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Chilcote:1986:DRS

[CLL86] Mark W. Chilcote, Steven A. Leider, and John J. Loch. Differential reproductive success of hatchery and wild summer-run steelhead under natural conditions. Transactions of the American Fisheries Society, 115(5):726–735, September 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cada:1987:EFL

[CLS87] Glenn F. Cada, James M. Loar, and Michael J. Sale. Evidence of food limitation of rainbow and brown trout in Southern Appalachian soft-water streams. Transactions of the American Fisheries Society, 116(5):692–702, September 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Comyns:1989:AGR

[CLSRS89] Bruce H. Comyns, Joanne Lyczkowski-Shultz, Chester F. Rakocinski, and John P. Steen, Jr. Age and growth of red drum larvae in the North-Central Gulf of Mexico. Transactions of the American Fisheries Society, 118(2):159–167, March 1989. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cooper:1981:BMS

[CM81] Jon C. Cooper and John V. Merriner. Biology and management of striped bass introduction. Transactions of the American Fisheries Society, 110(1):93–94, January 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Campbell:1983:DAL

[CM83] A. Campbell and R. K. Mohn. Definition of American lobster stocks for the Canadian maritimes by analysis of fishery-landing trends. Transactions of the American Fisheries Society, 112(6): 744–759, November 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Canfield:1985:POE

[CMNS85] Daniel E. Canfield, Jr., Michael J. Maceina, Frank G. Nordlie, and Jerome V. Shireman. Plasma osmotic and electrolyte concentrations of largemouth bass from some acidic Florida lakes. Transactions of the American Fisheries Society, 114(3):423–429, May 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Coble:1982:FPR

[Cob82] Daniel W. Coble. Fish populations in relation to dissolved oxygen in the Wisconsin River. Transactions of the American Fisheries Society, 111(5):612–623, September 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Colt:1984:SCD

[Col84] John Colt. Seasonal changes in dissolved-gas supersaturation in the Sacramento River and possible effects on striped bass. Transactions of the American Fisheries Society, 113(5):655–665, September 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cone:1989:NRU

[Con89] R. Scott Cone. The need to reconsider the use of condition indices in fishery science. Transactions of the American Fisheries Society, 118(5):510–514, September 1989. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Coulter:1981:BPP

[Cou81] George W. Coulter. Biomass, production, and potential yield of the Lake Tanganyika pelagic fish community. Transactions of the American Fisheries Society, 110(3):325–335, May 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Coutant:1985:SBE

[Cou85a] Charles C. Coutant. Striped bass: Environmental risks in fresh and salt water. Transactions of the American Fisheries Society, 114(1):1–2, January 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Coutant:1985:SBT

[Cou85b] Charles C. Coutant. Striped bass, temperature, and dissolved oxygen: a speculative hypothesis for environmental risk. Transactions of the American Fisheries Society, 114(1):31–61, January 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Coutant:1987:PRS

[Cou87] Charles C. Coutant. Poor reproductive success of striped bass from a reservoir with reduced summer habitat. Transactions of the American Fisheries Society, 116(2):154–160, March 1987. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cooper:1981:RYC

[CP81] Jon C. Cooper and Tibor T. Polgar. Recognition of year-class dominance in striped bass management. Transactions of the American Fisheries Society, 110(1):180–187, January 1981. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Corbett:1983:SEL

[CP83] B. Corbett and P. M. Powles. Spawning and early-life ecological phases of the white sucker in Jack lake, Ontario. Transactions of the American Fisheries Society, 112(2):308–313, March 1983. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Corbett:1986:SLD

[CP86] B. W. Corbett and P. M. Powles. Spawning and larva drift of sympatric walleyes and white suckers in an Ontario stream. Transactions of the American Fisheries Society, 115(1):41–46, January 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Costa-Pierce:1985:EPM

[CPML85] Barry A. Costa-Pierce, Spencer R. Malecha, and Edward A. Laws. Effects of polyculture and manure fertilization on water quality and heterotrophic productivity in Macrobrachium rosenbergii ponds. Transactions of the American Fisheries Society, 114(6): 826–836, November 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Craig:1984:FUC

[Cra84] P. C. Craig. Fish use of coastal waters of the Alaskan Beaufort Sea: a review. Transactions of the American Fisheries Society, 113(3):265–282, May 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Chen:1983:EMK

[CRF83] Martin F. Chen, J. S. Rohovec, and J. L. Fryer. Effect of methotrexate, kenalog, and cyclophosphamide on the antibody response of steelheads to Aeromonas salmonicida. Transactions of the American Fisheries Society, 112(4):561–564, July 1983. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Carmichael:1984:CAS

[CTSD84a] G. J. Carmichael, J. R. Tomasso, B. A. Simco, and K. B. Davis. Characterization and alleviation of stress associated with hauling largemouth bass. Transactions of the American Fisheries Society, 113(6):778–785, November 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Carmichael:1984:CWQ

[CTSD84b] G. J. Carmichael, J. R. Tomasso, B. A. Simco, and K. B. Davis. Confinement and water quality-induced stress in largemouth bass. Transactions of the American Fisheries Society, 113(6):767–777, November 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cunjak:1989:ASP

[Cun89] Richard A. Cunjak. Atlantic salmon: Planning for the future. proceedings of the third international Atlantic salmon symposium. Transactions of the American Fisheries Society, 118(4): 454–460, July 1989. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Coughlan:1989:DTS

[CV89] David J. Coughlan and John S. Velte. Dietary toxicity of selenium-contaminated red shiners to striped bass. Transactions of the American Fisheries Society, 118(4):400–408, July 1989. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cheek:1985:IWQ

[CVC85] T. E. Cheek, M. J. Van Den Avyle, and C. C. Coutant. Influences of water quality on distribution of striped bass in a Tennessee River impoundment. Transactions of the American Fisheries Society, 114(1):67–76, January 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Colt:1982:PGS

[CW82] John Colt and Harry Westers. Production of gas supersaturation by aeration. Transactions of the American Fisheries Society, 111(3):342–360, May 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Congleton:1988:ELI

[CW88] James L. Congleton and Eric J. Wagner. Effects of light intensity on plasma cortisol concentrations in migrating smolts of Chinook salmon and steelhead held in tanks or raceways and after passage through experimental flumes. Transactions of the American Fisheries Society, 117(4):385–393, July 1988. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cobb:1989:SDS

[CWCR89] J. S. Cobb, D. Wang, D. B. Campbell, and P. Rooney. Speed and direction of swimming by postlarvae of the American lobster. Transactions of the American Fisheries Society, 118(1):82– 86, January 1989. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Carmichael:1986:GMI

[CWSM86] Gary J. Carmichael, J. Holt Williamson, Maureen E. Schmidt, and Donald C. Morizot. Genetic marker identification in largemouth bass with electrophoresis of low-risk tissues. Transactions of the American Fisheries Society, 115(3):455–459, May 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Chapman:1986:ERF

[CWW^{+}86] D. W. Chapman, D. E. Weitkamp, T. L. Welsh, M. B. Dell, and T. H. Schadt. Effects of river flow on the distribution of Chinook salmon redds. Transactions of the American Fisheries Society, 115(4):537–547, July 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Coutant:1984:TSJ

[CZCP84] Charles C. Coutant, Kenneth L. Zachman, David K. Cox, and Bradley L. Pearman. Temperature selection by juvenile striped bass in laboratory and field. Transactions of the American Fisheries Society, 113(5):666–671, September 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Donald:1982:IES

[DA82] D. B. Donald and R. S. Anderson. Importance of environment and stocking density for growth of rainbow trout in mountain lakes. Transactions of the American Fisheries Society, 111(6):

675–680, November 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dabrowski:1989:ALF

[Dab89] Konrad Dabrowski. 12th annual larval fish conference formulation of a bioenergetic model for coregonine early life history. Transactions of the American Fisheries Society, 118(2):138–150, March 1989. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dawe:1988:LWR

[Daw88] Earl G. Dawe. Length–weight relationships for short-finned squid in Newfoundland and the effect of diet on condition and growth. Transactions of the American Fisheries Society, 117(6):591–599, November 1988. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Degraaf:1986:HUP

[DB86] D. A. Degraaf and L. H. Bain. Habitat use by and preferences of juvenile Atlantic salmon in two Newfoundland Rivers. Transactions of the American Fisheries Society, 115(5):671–681, September 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dauble:1983:CEC

[DBBF83] Dennis D. Dauble, Stephen A. Barraclough, Roger M. Bean, and William E. Fallon. Chronic effects of coal-liquid dispersions on fathead minnows and rainbow trout. Transactions of the American Fisheries Society, 112(5):712–719, September 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Deangelis:1982:GBS

[DC82] D. L. Deangelis and C. C. Coutant. Genesis of bimodal size distributions in species cohorts. Transactions of the American Fisheries Society, 111(3):384–388, May 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dames:1989:MCF

[DCR89] H. Ross Dames, Thomas G. Coon, and John W. Robinson. Movements of channel and flathead catfish between the Missouri River and a tributary, Perche Creek. Transactions of the American Fisheries Society, 118(6):670–679, November 1989. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dunham:1987:TMH

[DEAT87] Rex A. Dunham, Jill Eash, Janine Askins, and Tim M. Townes. Transfer of the metallothionein-human growth hormone fusion gene into channel catfish. Transactions of the American Fisheries Society, 116(1):87–91, January 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Deegan:1981:IES

[Dee81] Linda A. Deegan. Increased estuarine survival of juvenile Coho salmon vaccinated against Vibrio anguillarum. Transactions of the American Fisheries Society, 110(5):656–659, September 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dey:1981:MGY

[Dey81] William P. Dey. Mortality and growth of young-of-the-year striped bass in the Hudson River Estuary. Transactions of the American Fisheries Society, 110(1):151–157, January 1981. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dochoda:1987:PPG

[DF87] Margaret Ross Dochoda and Carlos M. Fetterolf, Jr. Public purpose of great lakes fishery management: Lessons from the management experience. Transactions of the American Fisheries Society, 116(3):302–308, May 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dutil:1988:TIM

[DGK^{+}88] J.-D. Dutil, A. Giroux, A. Kemp, G. Lavoie, and J.-P. Dallaire. Tidal influence on movements and on daily cycle of activity of American eels. Transactions of the American Fisheries Society, 117(5):488–494, September 1988. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dauble:1985:AWS

[DGS^{+}85] Dennis D. Dauble, Robert H. Gray, J. R. Skalski, E. W. Lusty, and M. A. Simmons. Avoidance of a water-soluble fraction of coal liquid by fathead minnows. Transactions of the American Fisheries Society, 114(5):754–760, September 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dawson:1983:RMM

[DHSA83] V. K. Dawson, P. D. Harman, D. P. Schultz, and J. L. Allen. Rapid method for measuring rotenone in water at piscicidal concentrations. Transactions of the American Fisheries Society, 112 (5):725–727, September 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Drenner:1987:PIT

[DHVG87] Ray W. Drenner, K. David Hambright, Gary L. Vinyard, and Moshe Gophen. Particle ingestion by Tilapia galilaea is not affected by removal of gill rakers and microbranchiospines. Transactions of the American Fisheries Society, 116(2):272–276, March 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Diana:1983:GMP

[Dia83] James S. Diana. Growth, maturation, and production of northern pike in three Michigan lakes. Transactions of the American Fisheries Society, 112(1):38–46, January 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Diana:1987:SMC

[Dia87] James S. Diana. Simulation of mechanisms causing stunting in northern pike populations. Transactions of the American Fisheries Society, 116(4):612–617, July 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Davin:1986:ECT

[DKT86] William T. Davin, Jr., Christopher C. Kohler, and Donald R. Tindall. Effects of ciguatera toxins on the bluehead. Transactions of the American Fisheries Society, 115(6):908–912, November 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Davin:1988:CTA

[DKT88] William T. Davin, Jr., Christopher C. Kohler, and Donald R. Tindall. Ciguatera toxins adversely affect piscivorous fishes. Transactions of the American Fisheries Society, 117(4):374–384, July 1988. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dindo:1981:ACS

[DM81] J. J. Dindo and R. Macgregor III. Annual cycle of serum gonadal steroids and serum lipids in striped mullet. Transactions of the American Fisheries Society, 110(3):403–409, May 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dombeck:1984:MSH

[DMH84] Michael P. Dombeck, Bruce W. Menzel, and Paul N. Hinz. Muskellunge spawning habitat and reproductive success. Transactions of the American Fisheries Society, 113(2):205–216, March 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Drenner:1982:FFR

[DMO82] Ray W. Drenner, John R. Mummert, and W. John O’Brien. Filter-feeding rates of gizzard shad. Transactions of the American Fisheries Society, 111(2):210–215, March 1982. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dolloff:1986:ESC

[Dol86] C. Andrew Dolloff. Effects of stream cleaning on juvenile Coho salmon and Dolly Varden in Southeast Alaska. Transactions of the American Fisheries Society, 115(5):743–755, September 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dolloff:1987:SPC

[Dol87] C. Andrew Dolloff. Seasonal population characteristics and habitat use by juvenile Coho salmon in a small Southeast Alaska stream. Transactions of the American Fisheries Society, 116(6): 829–838, November 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

DeSilva:1985:EDP

[DP85] Sena S. De Silva and Mala K. Perera. Effects of dietary protein level on growth, food conversion, and protein use in young Tilapia nilotica at four salinities. Transactions of the American Fisheries Society, 114(4):584–589, July 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Davis:1986:PCS

[DP86] Kenneth B. Davis and Nick C. Parker. Plasma corticosteroid stress response of fourteen species of warmwater fish to transportation. Transactions of the American Fisheries Society, 115 (3):495–499, May 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Deriso:1987:OCF

[DP87] Richard B. Deriso and Ana M. Parma. On the odds of catching fish with angling gear. Transactions of the American Fisheries Society, 116(2):244–256, March 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dubois:1989:FSF

[DPR89] Robert B. Dubois, Scott D. Plaster, and Paul W. Rasmussen. Fecundity of spring- and fall-run steelhead from two Western Lake Superior tributaries. Transactions of the American Fisheries Society, 118(3):311–316, May 1989. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dealteris:1989:CPT

[DRFL89] Joseph T. Dealteris, Conrad W. Recksiek, Abdesslam Fahfouhi, and Xu Liuxiong. Comparison of the performance of two bottomsampling trawls. Transactions of the American Fisheries Society, 118(2):119–130, March 1989. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Die:1988:UPR

[DRH88] David J. Die, Victor R. Restrepo, and John M. Hoenig. Utilityper-recruit modeling: a neglected concept. Transactions of the American Fisheries Society, 117(3):274–281, May 1988. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dorsa:1982:EDC

[DRRP82] Warren J. Dorsa, H. Randall Robinette, Edwin H. Robinson, and W. E. Poe. Effects of dietary cottonseed meal and gossypol on growth of young channel catfish. Transactions of the American Fisheries Society, 111(5):651–655, September 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Daan:1985:PNS

[DRV85] N. Daan, A. D. Rijnsdorp, and G. R. Van Overbeeke. Predation by North Sea herring Clupea harengus on eggs of plaice Pleuronectes platessa and cod Gadus morhua. Transactions of the American Fisheries Society, 114(4):499–506, July 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Daly:1985:ICS

[DS85] J. G. Daly and R. M. W. Stevenson. Importance of culturing several organs to detect Aeromonas salmonicida in salmonid fish. Transactions of the American Fisheries Society, 114(6):909–910, November 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dunham:1983:RPC

[DSHB83] Rex A. Dunham, R. Oneal Smitherman, Jamie L. Horn, and Timothy O. Bice. Reproductive performances of crossbred and purestrain channel catfish brood stocks. Transactions of the American Fisheries Society, 112(3):436–440, May 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Davis:1985:DGI

[DSM85] Robert D. Davis, Ted W. Storck, and Steven J. Miller. Daily growth increments in the otoliths of young-of-the-year gizzard shad. Transactions of the American Fisheries Society, 114(2): 304–306, March 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Davis:1984:BAI

[DSP84] Kenneth B. Davis, Mary Anna Suttle, and Nick C. Parker. Biotic and abiotic influences on corticosteroid hormone rhythms in channel catfish. Transactions of the American Fisheries Society, 113(4):414–421, July 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Day:1984:PTI

[DT84a] Jonathan R. Day and Malcolm H. Taylor. Photoperiod and temperature interaction in the seasonal reproduction of female mummichogs. Transactions of the American Fisheries Society, 113(4): 452–457, July 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Delahunty:1984:PIA

[DT84b] George Delahunty and Michele Tomlinson. Photoperiodindependent actions of the pineal organ: Aspects of a pineal organ-pancreas relationship. Transactions of the American Fisheries Society, 113(4):439–443, July 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Deegan:1987:GRL

[DT87] Linda A. Deegan and Bruce A. Thompson. Growth rate and life history events of young-of-the-year Gulf menhaden as determined from otoliths. Transactions of the American Fisheries Society, 116(4):663–667, July 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Drenner:1984:PGP

[DTLK84] Ray W. Drenner, Scott B. Taylor, Xavier Lazzaro, and Dean Kettle. Particle-grazing and plankton community impact of an omnivorous cichlid. Transactions of the American Fisheries Society, 113(3):397–402, May 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Duplinsky:1982:SMN

[Dup82] Paul D. Duplinsky. Sperm motility of northern pike and chain pickerel at various pH values. Transactions of the American Fisheries Society, 111(6):768–771, November 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Editor:1986:NLF

[EB86] Clark Hubbs Special Editor and J. H. S. Blaxter. Ninth larval fish conference: Development of sense organs and behaviour of teleost larvae with special reference to feeding and predator avoidance. Transactions of the American Fisheries Society, 115(1):98–114, January 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Essig:1986:MEL

[EC86] Ronald J. Essig and Charles F. Cole. Methods of estimating larval fish mortality from daily increments in otoliths. Transactions of the American Fisheries Society, 115(1):34–40, January 1986. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Eaton:1986:RTE

[ED86] Robert C. Eaton and Randolf Didomenico. Role of the teleost escape response during development. Transactions of the American Fisheries Society, 115(1):128–142, January 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ehrhardt:1988:SGN

[ED88a] Nelson M. Ehrhardt and David J. Die. Selectivity of gill nets used in the commercial Spanish mackerel fishery of Florida. Transactions of the American Fisheries Society, 117(6):574–580, November 1988. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ehrhardt:1988:SSY

[ED88b] Nelson M. Ehrhardt and David J. Die. Size-structured yieldper-recruit simulation for the Florida gill-net fishery for Spanish mackerel. Transactions of the American Fisheries Society, 117 (6):581–590, November 1988. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Editor:1988:TCN

[Edi88] Robert L. Kendall Managing Editor. Taxonomic changes in North American trout names. Transactions of the American Fisheries Society, 117(4):321, July 1988. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Editor:1989:FGN

[Edi89] Robert L. Kendall Managing Editor. Fish genetic nomenclature. Transactions of the American Fisheries Society, 118(2):218–227, March 1989. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Edwards:1989:FSP

[Edw89] Steven F. Edwards. Forecasts of in-state participation in marine recreational fishing. Transactions of the American Fisheries Society, 118(5):564–572, September 1989. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Edds:1989:GCH

[EE89] David R. Edds and Anthony A. Echelle. Genetic comparisons of hatchery and natural stocks of small endangered fishes: Leon springs pupfish, comanche springs pupfish, and pecos Gambusia.

Transactions of the American Fisheries Society, 118(4):441–446, July 1989. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Erickson:1986:RDT

[EG86] Daniel L. Erickson and Gary D. Grossman. Reproductive demography of tilefish from the South Atlantic bight with a test for the presence of protogynous hermaphroditism. Transactions of the American Fisheries Society, 115(2):279–285, March 1986. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Eames:1983:EFT

[EH83] Michael J. Eames and Mark K. Hino. An evaluation of four tags suitable for marking juvenile Chinook salmon. Transactions of the American Fisheries Society, 112(4):464–468, July 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Engstrom-Heg:1986:IAC

[EH86] Robert Engstrom-Heg. Interaction of area with catchability indices used in analyzing inland recreational fisheries. Transactions of the American Fisheries Society, 115(6):818–822, November 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Elliott:1986:RDV

[Ell86] S. T. Elliott. Reduction of a Dolly Varden population and macrobenthos after removal of logging debris. Transactions of the American Fisheries Society, 115(3):392–400, May 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ellsaesser:1985:ACC

[EMC^{+}85] Catherine F. Ellsaesser, Norman W. Miller, Marvin A. Cuchens, Craig J. Lobb, and L. William Clem. Analysis of channel catfish peripheral blood leucocytes by bright-field microscopy and flow cytometry. Transactions of the American Fisheries Society, 114 (2):279–285, March 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ennis:1984:SSS

[Enn84a] G. P. Ennis. Small-scale seasonal movements of the American lobster Homarus americanus. Transactions of the American Fisheries Society, 113(3):336–338, May 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ennis:1984:TBA

[Enn84b] G. P. Ennis. Territorial behavior of the American lobster Homarus americanus. Transactions of the American Fisheries Society, 113(3):330–335, May 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ennis:1986:STL

[Enn86] G. P. Ennis. Sphyrion tag loss from the American lobster Homarus americanus. Transactions of the American Fisheries Society, 115(6):914–917, November 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Eriksson:1984:AAC

[Eri84] Torleif Eriksson. Adjustments in annual cycles of swimming behaviour in juvenile Baltic salmon in fresh and brackish water. Transactions of the American Fisheries Society, 113(4):467–471, July 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Erman:1986:LTS

[Erm86] Don C. Erman. Long-term structure of fish populations in Sagehen Creek, California. Transactions of the American Fisheries Society, 115(5):682–692, September 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Eshenroder:1987:SAL

[Esh87] Randy L. Eshenroder. Socioeconomic aspects of lake trout rehabilitation in the great lakes. Transactions of the American Fisheries Society, 116(3):309–313, May 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Evans:1984:TIA

[Eva84] David O. Evans. Temperature independence of the annual cycle of standard metabolism in the pumpkinseed. Transactions of the American Fisheries Society, 113(4):494–512, July 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Eldridge:1981:EFF

[EWE^{+}81] Maxwell B. Eldridge, Jeannette A. Whipple, Dana Eng, Michael J. Bowers, and Brian M. Jarvis. Effects of food and feeding factors on laboratory-reared striped bass larvae. Transactions of the American Fisheries Society, 110(1):111–120, January 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fabrizio:1987:CCB

[Fab87a] Mary C. Fabrizio. Contribution of Chesapeake Bay and Hudson River stocks of striped bass to rhode Island coastal waters as estimated by isoelectric focusing of eye lens proteins. Transactions of the American Fisheries Society, 116(4):588–593, July 1987. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fabrizio:1987:GID

[Fab87b] Mary C. Fabrizio. Growth-invariant discrimination and classification of striped bass stocks by morphometric and electrophoretic methods. Transactions of the American Fisheries Society, 116 (5):728–736, September 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fabrizio:1988:LNL

[Fab88] Mary C. Fabrizio. Luminescence, narcosis, and life in the deep sea. Transactions of the American Fisheries Society, 117(6):603– 606, November 1988. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fournier:1983:EAM

[FB83] D. A. Fournier and P. A. Breen. Estimation of abalone mortality rates with growth analysis. Transactions of the American Fisheries Society, 112(3):403–411, May 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Finger:1988:TWT

[FB88] Susan E. Finger and James S. Bulak. Toxicity of water from three South Carolina Rivers to larval striped bass. Transactions of the American Fisheries Society, 117(6):521–528, November 1988. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fisher:1982:GRT

[FBCV82] P. W. Fisher, D. Browne, D. G. Cameron, and E. R. Vyse. Genetics of rainbow trout in a geothermally heated stream. Transactions of the American Fisheries Society, 111(3):312–316, May 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fargo:1987:AVR

[FC87] J. Fargo and D. E. Chilton. Age validation for rock sole in Hecate Strait, British Columbia. Transactions of the American Fisheries Society, 116(5):776–778, September 1987. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Flores-Coto:1983:SDA

[FCBTSR83] César Flores-Coto, Francisco Barba-Torres, and Jesús Sánchez-Robles. Seasonal diversity, abundance, and distribution of ichthyoplankton in tamiahua lagoon, Western Gulf of Mexico. Transactions of the American Fisheries Society, 112(2):247–256, March 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Faulmann:1983:ECS

[FCL^{+}83a] E. Faulmann, M. A. Cuchens, C. J. Lobb, N. W. Miller, and L. W. Clem. An effective culture system for studying in vitro mitogenic responses of channel catfish lymphocytes. Transactions of the American Fisheries Society, 112(5):673–679, September 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fuiman:1983:SAI

[FCL^{+}83b] Lee A. Fuiman, John V. Conner, Barbara F. Lathrop, Gerard L. Buynak, Darrel E. Snyder, and Jules J. Loos. State of the art of identification for cyprinid fish larvae from Eastern North America. Transactions of the American Fisheries Society, 112(2):319–332, March 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Flath:1985:SED

[FD85] Loren E. Flath and James S. Diana. Seasonal energy dynamics of the alewife in Southeastern Lake Michigan. Transactions of the American Fisheries Society, 114(3):328–337, May 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ferguson:1989:NET

[FD89] Moira M. Ferguson and Lenore R. Drahushchak. Notes: Effects of tissue collection and storage methods on nucleic acid determinations in white muscle of fishes. Transactions of the American Fisheries Society, 118(6):709–713, November 1989. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fedkiw:1987:CBM

[Fed87] John Fedkiw. Coming back to market value and valuation for the great lakes fisheries. Transactions of the American Fisheries Society, 116(3):346–351, May 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Finlayson:1986:CLF

[FF86] B. J. Finlayson and G. A. Faggella. Comparison of laboratory and field observations of fish exposed to the herbicides molinate and thiobencarb. Transactions of the American Fisheries Society, 115 (6):882–890, November 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Faber:1983:SDT

[FG83] Daniel J. Faber and Sally Gadd. Several drawing techniques to illustrate larval fishes. Transactions of the American Fisheries Society, 112(2):349–353, March 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Farquhar:1989:DMA

[FG89] Bobby W. Farquhar and Steve Gutreuter. Distribution and migration of adult striped bass in Lake Whitney, Texas. Transactions of the American Fisheries Society, 118(5):523–532, September 1989. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Friedland:1988:IVD

[FGB^{+}88] K. D. Friedland, G. C. Garman, A. J. Bejda, A. L. Studholme, and B. Olla. Interannual variation in diet and condition in juvenile bluefish during estuarine residency. Transactions of the American Fisheries Society, 117(5):474–479, September 1988. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fries:1989:NNH

[FH89] L. T. Fries and W. D. Harvey. Notes: Natural hybridization of white bass with yellow bass in Texas. Transactions of the American Fisheries Society, 118(1):87–89, January 1989. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Floyd:1984:CAH

[FHT84] Keith B. Floyd, Robert D. Hoyt, and Shirley Timbrook. Chronology of appearance and habitat partitioning by stream larval fishes.

Transactions of the American Fisheries Society, 113(2):217–223, March 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fogarty:1988:AYE

[FI88] Michael J. Fogarty and Josef S. Idoine. Application of a yield and egg production model based on size to an offshore American lobster population. Transactions of the American Fisheries Society, 117(4):350–362, July 1988. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fisher:1987:BFS

[Fis87] William L. Fisher. Benthic fish sampler for use in riffle habitats. Transactions of the American Fisheries Society, 116(5): 768–772, September 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fields:1989:DFR

[FJT89] Robert D. Fields, Kenneth R. Johnson, and Gary H. Thorgaard. DNA fingerprints in rainbow trout detected by hybridization with DNA of bacteriophage M13. Transactions of the American Fisheries Society, 118(1):78–81, January 1989. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fischer:1988:PPI

[FK88] Steven A. Fischer and William E. Kelso. Potential parasiteinduced mortality in age-0 bluegills in a floodplain pond of the Lower Mississippi River. Transactions of the American Fisheries Society, 117(6):565–573, November 1988. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fausch:1984:RAI

[FKY84] Kurt D. Fausch, James R. Karr, and Philip R. Yant. Regional application of an index of biotic integrity based on stream fish communities. Transactions of the American Fisheries Society, 113(1):39–55, January 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Facey:1981:BAE

[FL81] Douglas E. Facey and George W. Labar. Biology of American eels in Lake Champlain, Vermont. Transactions of the American Fisheries Society, 110(3):396–402, May 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fletcher:1985:RAF

[Fle85] R. Ian Fletcher. Risk analysis for fish diversion experiments: Pumped intake systems. Transactions of the American Fisheries Society, 114(5):652–694, September 1985. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Flebbe:1989:B

[Fle89] Patricia A. Flebbe. Biodiversity. Transactions of the American Fisheries Society, 118(2):228, March 1989. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fields:1987:CCT

[FLK^{+}87] Robert Fields, Shirley S. Lowe, Christine Kaminski, Gregory S. Whitt, and David P. Philipp. Critical and chronic thermal maxima of northern and Florida largemouth bass and their reciprocal F_{1} and F_{2} hybrids. Transactions of the American Fisheries Society, 116(6):856–863, November 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fagerlund:1981:SRE

[FMS81] U. H. M. Fagerlund, J. R. McBride, and E. T. Stone. Stressrelated effects of hatchery rearing density on Coho salmon. Transactions of the American Fisheries Society, 110(5):644–649, September 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Foster:1985:LTR

[Fos85] Neal R. Foster. Lake trout reproductive behavior: Influence of chemosensory cues from young-of-the-year by-products. Transactions of the American Fisheries Society, 114(6):794–803, November 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fraser:1989:ILT

[FP89] Noel C. Fraser and Geoffrey Power. Influences of lake trout on lake-resident Arctic char in Northern Quebec, Canada. Transactions of the American Fisheries Society, 118(1):36–45, January 1989. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fraidenburg:1981:FEN

[Fra81] Michael E. Fraidenburg. First estimates of natural mortality for yellowtail rockfish. Transactions of the American Fisheries Society, 110(4):551–553, July 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fish:1983:VHR

[FS83] Pamela A. Fish and J. Savitz. Variations in home ranges of largemouth bass, yellow perch, bluegills, and pumpkinseeds in an Illinois lake. Transactions of the American Fisheries Society, 112(2): 147–153, March 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Frie:1985:DRS

[FS85] Richard V. Frie and George R. Spangler. Dynamics of rainbow smelt during and after exploitation in South Bay, Lake Huron. Transactions of the American Fisheries Society, 114(5):713–724, September 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fuiman:1986:BSP

[Fui86] Lee A. Fuiman. Burst-swimming performance of larval zebra danios and the effects of diel temperature fluctuations. Transactions of the American Fisheries Society, 115(1):143–148, January 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Finlayson:1982:TCZ

[FV82] B. J. Finlayson and K. M. Verrue. Toxicities of copper, zinc, and cadmium mixtures to juvenile Chinook salmon. Transactions of the American Fisheries Society, 111(5):645–650, September 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fausch:1986:CAJ

[FW86] Kurt D. Fausch and Ray J. White. Competition among juveniles of Coho salmon, brook trout, and brown trout in a laboratory stream, and implications for great lakes tributaries. Transactions of the American Fisheries Society, 115(3):363–381, May 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Goolish:1983:GUF

[GA83] Edward M. Goolish and Ira R. Adelman. ^{14}C-glycine uptake by fish scales: Refinement of a growth index and effects of a proteinsynthesis inhibitor. Transactions of the American Fisheries Society, 112(5):647–652, September 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gutreuter:1985:IBS

[GA85] Steven J. Gutreuter and Richard O. Anderson. Importance of body size to the recruitment process in largemouth bass populations. Transactions of the American Fisheries Society, 114(3): 317–327, May 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gale:1983:FSF

[Gal83] William F. Gale. Fecundity and spawning frequency of caged bluntnose minnows — fractional spawners. Transactions of the American Fisheries Society, 112(3):398–402, May 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gale:1986:IFS

[Gal86] William F. Gale. Indeterminate fecundity and spawning behavior of captive red shiners — fractional, crevice spawners. Transactions of the American Fisheries Society, 115(3):429–437, May 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gale:1987:SAP

[Gal87] Richard P. Gale. Social assessment from pines to perch: Comparative observations of fisheries and forestry. Transactions of the American Fisheries Society, 116(3):486–493, May 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gardner:1981:ETF

[Gar81] Mark B. Gardner. Effects of turbidity on feeding rates and selectivity of bluegills. Transactions of the American Fisheries Society, 110(3):446–450, May 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Garman:1983:OJR

[Gar83] Greg C. Garman. Observations on juvenile red hake associated with sea scallops in Frenchman Bay, Maine. Transactions of the American Fisheries Society, 112(2):212–215, March 1983. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Grimes:1982:DOS

[GAT82] C. B. Grimes, K. W. Able, and S. C. Turner. Direct observation from a submersible vessel of commercial longlines for tilefish. Transactions of the American Fisheries Society, 111(1):94– 98, January 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gale:1982:FSF

[GB82] William F. Gale and Gerard L. Buynak. Fecundity and spawning frequency of the fathead minnow — a fractional spawner. Transactions of the American Fisheries Society, 111(1):35–40, January 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gillespie:1986:EHT

[GB86] Robert B. Gillespie and Paul C. Baumann. Effects of high tissue concentrations of selenium on reproduction by bluegills. Transactions of the American Fisheries Society, 115(2):208–213, March 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Goedde:1981:EAP

[GC81] Larry E. Goedde and Daniel W. Coble. Effects of angling on a previously fished and an unfished warmwater fish community in two Wisconsin lakes. Transactions of the American Fisheries Society, 110(5):594–603, September 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gallagher:1983:CTI

[GC83] Robert P. Gallagher and John V. Conner. Comparison of two ichthyoplankton sampling gears with notes on microdistribution of fish larvae in a large river. Transactions of the American Fisheries Society, 112(2):280–285, March 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Goodyear:1984:BEF

[GC84] C. Phillip Goodyear and Sigurd W. Christensen. Bias-elimination in fish population models with stochastic variation in survival of the young. Transactions of the American Fisheries Society, 113 (5):627–632, September 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gotceitas:1987:SBD

[GC87] Vytenis Gotceitas and Patrick Colgan. Selection between densities of artificial vegetation by young bluegills avoiding predation. Transactions of the American Fisheries Society, 116(1): 40–49, January 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Goodyear:1985:MSB

[GCC85] C. Phillip Goodyear, Joel E. Cohen, and Sigurd W. Christensen. Maryland striped bass: Recruitment declining below replacement. Transactions of the American Fisheries Society, 114(1):146–151, January 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Giattina:1981:CLF

[GCCL81] J. D. Giattina, D. S. Cherry, John Cairns, Jr., and S. R. Larrick. Comparison of laboratory and field avoidance behavior of fish in heated chlorinated water. Transactions of the American Fisheries Society, 110(4):526–535, July 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gale:1985:FSF

[GD85a] William F. Gale and William G. Deutsch. Fecundity and spawning frequency of captive tessellated darters — fractional spawners. Transactions of the American Fisheries Society, 114(2):220–229, March 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Glova:1985:PER

[GD85b] Gordon J. Glova and Maurice J. Duncan. Potential effects of reduced flows on fish habitats in a large braided river, New Zealand. Transactions of the American Fisheries Society, 114(2):165–181, March 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Grall:1983:DRA

[GDH83] Charlene Grall, Donald P. De Sylva, and Edward D. Houde. Distribution, relative abundance, and seasonality of swordfish larvae. Transactions of the American Fisheries Society, 112(2):235–246, March 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gerking:1984:FPC

[Ger84] Shelby D. Gerking. Fish as primary consumers: Assimilation and maintenance ration of an herbivorous fish, Sarpa salpa, feeding on a green alga. Transactions of the American Fisheries Society, 113(3):378–387, May 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Giattina:1982:ACN

[GGS82] James D. Giattina, Ronald R. Garton, and Donald G. Stevens. Avoidance of copper and nickel by rainbow trout as monitored by a computer-based data acquisition system. Transactions of the American Fisheries Society, 111(4):491–504, July 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gibson:1984:HPR

[Gib84] R. N. Gibson. Hydrostatic pressure and the rhythmic behaviour of intertidal marine fishes. Transactions of the American Fisheries Society, 113(4):479–483, July 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gregory:1982:RDS

[GLC82] Douglas R. Gregory, Jr., Ronald F. Labisky, and Christopher L. Combs. Reproductive dynamics of the spiny lobster Panulirus argus in South Florida. Transactions of the American Fisheries Society, 111(5):575–584, September 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Glass:1987:PLU

[GM87] Ronald J. Glass and Robert M. Muth. Pitfalls and limitations in the use of fishery valuation techniques. Transactions of the American Fisheries Society, 116(3):381–389, May 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gunn:1987:SGS

[GMB^{+}87] J. M. Gunn, M. J. McMurtry, J. N. Bowlby, J. M. Casselman, and V. A. Liimatainen. Survival and growth of stocked lake trout in relation to body size, stocking season, lake acidity, and biomass of competitors. Transactions of the American Fisheries Society, 116(4):618–627, July 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Green:1983:MDE

[GMW83] Albert W. Green, Gary C. Matlock, and James E. Weaver. A method for directly estimating the tag-reporting rate of anglers. Transactions of the American Fisheries Society, 112(3):412–415, May 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Graham:1986:ETS

[GO86] Robert J. Graham and Donald J. Orth. Effects of temperature and streamflow on time and duration of spawning by smallmouth bass. Transactions of the American Fisheries Society, 115(5):693– 702, September 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Goff:1986:RSM

[Gof86] Gregory P. Goff. Reproductive success of male smallmouth bass in Long Point Bay, Lake Erie. Transactions of the American Fisheries Society, 115(3):415–423, May 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Goodyear:1985:RBR

[Goo85a] C. Phillip Goodyear. Relationship between reported commercial landings and abundance of young striped bass in Chesapeake Bay, Maryland. Transactions of the American Fisheries Society, 114 (1):92–96, January 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Goodyear:1985:TMF

[Goo85b] C. Phillip Goodyear. Toxic materials, fishing, and environmental variation: Simulated effects on striped bass population trends. Transactions of the American Fisheries Society, 114(1):107–113, January 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gregory:1987:NMN

[Gre87] Robin Gregory. Nonmonetary measures of nonmarket fishery resource benefits. Transactions of the American Fisheries Society, 116(3):374–380, May 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Grizzle:1981:EHD

[Gri81] John M. Grizzle. Effects of hypolimnetic discharge on fish health below a reservoir. Transactions of the American Fisheries Society, 110(1):29–43, January 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Goudie:1983:FCC

[GRSD83] Cheryl A. Goudie, Barry D. Redner, Bill A. Simco, and Kenneth B. Davis. Feminization of channel catfish by oral administration of steroid sex hormones. Transactions of the American Fisheries Society, 112(5):670–672, September 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gatlin:1982:FWG

[GS82a] Delbert M. Gatlin III and Robert R. Stickney. Fall-winter growth of young channel catfish in response to quantity and source of dietary lipid. Transactions of the American Fisheries Society, 111(1):90–93, January 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gradall:1982:RBT

[GS82b] Kenneth S. Gradall and William A. Swenson. Responses of brook trout and creek chubs to turbidity. Transactions of the American Fisheries Society, 111(3):392–395, May 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gillen:1981:PPR

[GSC81] Alan L. Gillen, Roy A. Stein, and Robert F. Carline. Predation by pellet-reared tiger muskellunge on minnows and bluegills in experimental systems. Transactions of the American Fisheries Society, 110(2):197–209, March 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gharrett:1988:GRE

[GSMH88] A. J. Gharrett, Cameo Smoot, A. J. McGregor, and P. B. Holmes. Genetic relationships of even-year Northwestern Alaskan pink salmon. Transactions of the American Fisheries Society, 117(6): 536–545, November 1988. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Graham:1985:MGT

[GT85] Joseph J. Graham and David W. Townsend. Mortality, growth, and transport of larval Atlantic herring Clupea harengus in Maine coastal waters. Transactions of the American Fisheries Society, 114(4):490–498, July 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Galat:1983:FPB

[GV83] David L. Galat and Nancy Vucinich. Food partitioning between young of the year of two sympatric tui chub morphs. Transactions of the American Fisheries Society, 112(4):486–497, July 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Graham:1988:DCA

[GV88] John H. Graham and Robert C. Vrijenhoek. Detrended correspondence analysis of dietary data. Transactions of the American Fisheries Society, 117(1):29–36, January 1988. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gloss:1983:MJS

[GW83] Steven P. Gloss and James R. Wahl. Mortality of juvenile salmonids passing through ossberger crossflow turbines at smallscale hydroelectric sites. Transactions of the American Fisheries Society, 112(2):194–200, March 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Harader:1983:ATC

[HA83a] Rolf R. Harader, Jr. and George H. Allen. Ammonia toxicity to Chinook salmon parr: Reduction in saline water. Transactions of the American Fisheries Society, 112(6):834–837, November 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Holt:1983:EAN

[HA83b] G. Joan Holt and C. R. Arnold. Effects of ammonia and nitrite on growth and survival of red drum eggs and larvae. Transactions of the American Fisheries Society, 112(2):314–318, March 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Henry:1986:BCS

[HA86] Mary G. Henry and Gary J. Atchison. Behavioral changes in social groups of bluegills exposed to copper. Transactions of the American Fisheries Society, 115(4):590–595, July 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Haines:1981:APC

[Hai81] Terry A. Haines. Acidic precipitation and its consequences for aquatic ecosystems: a review. Transactions of the American Fisheries Society, 110(6):669–707, November 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hankin:1982:EEP

[Han82] David G. Hankin. Estimating escapement of Pacific salmon: Marking practices to discriminate wild and hatchery fish. Transactions of the American Fisheries Society, 111(3):286–298, May 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hard:1986:PYJ

[Har86] Jeffrey J. Hard. Production and yield of juvenile Chinook salmon in two Alaskan lakes. Transactions of the American Fisheries Society, 115(2):305–313, March 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Harvey:1987:SYY

[Har87] Bret C. Harvey. Susceptibility of young-of-the-year fishes to downstream displacement by flooding. Transactions of the American Fisheries Society, 116(6):851–855, November 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hauck:1986:GBD

[Hau86] A. K. Hauck. Gas bubble disease due to helicopter transport of young pink salmon. Transactions of the American Fisheries Society, 115(4):630–635, July 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hay:1986:EDS

[Hay86] D. E. Hay. Effects of delayed spawning on viability of eggs and larvae of Pacific herring. Transactions of the American Fisheries Society, 115(1):155–161, January 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hunt:1981:ALA

[HB81] David Hunt and Claude E. Boyd. Alkalinity losses from ammonium fertilizers used in fish ponds. Transactions of the American Fisheries Society, 110(1):81–85, January 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Herman:1986:PCB

[HB86] Roger Lee Herman and G. L. Bullock. Pathology caused by the bacterium Edwardsiella tarda in striped bass. Transactions of the American Fisheries Society, 115(2):232–235, March 1986. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Helfman:1984:SAS

[HBB84] Gene S. Helfman, Earl L. Bozeman, and Edward B. Brothers. Size, age, and sex of American eels in a Georgia River. Transactions of the American Fisheries Society, 113(2):132–141, March 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Holland-Bartels:1988:VAY

[HBD88] Leslie E. Holland-Bartels and Michael C. Duval. Variations in abundance of young-of-the-year channel catfish in a navigation pool of the Upper Mississippi River. Transactions of the American Fisheries Society, 117(2):202–208, March 1988. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Haldorson:1984:LHE

[HC84] Lewis Haldorson and Peter Craig. Life history and ecology of a Pacific-Arctic population of rainbow smelt in coastal waters of the Beaufort Sea. Transactions of the American Fisheries Society, 113(1):33–38, January 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hart:1981:SMG

[HCFE81] C. E. Hart, Greg Concannon, C. A. Fustish, and R. D. Ewing. Seaward migration and gill (Na + K)-ATPase activity of spring Chinook salmon in an artificial stream. Transactions of the American Fisheries Society, 110(1):44–50, January 1981. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hodgson:1989:ESF

[HCG89] James R. Hodgson, Stephen R. Carpenter, and Alan P. Gripentrog. Effect of sampling frequency on intersample variance and food consumption estimates of nonpiscivorous largemouth bass. Transactions of the American Fisheries Society, 118(1):11–19, January 1989. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Harrell:1987:PIP

[HD87] Reginal M. Harrell and John Mark Dean. Pterygiophore interdigitation patterns and morphometry of larval hybrids of Morone species. Transactions of the American Fisheries Society, 116(5): 719–727, September 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Harrell:1988:IJH

[HD88] Reginal M. Harrell and John Mark Dean. Identification of juvenile hybrids of Morone based on meristics and morphometrics. Transactions of the American Fisheries Society, 117(6):529–535, November 1988. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hunter:1982:PAF

[HDGE82] George A. Hunter, Edward M. Donaldson, Fredrick W. Goetz, and Philip R. Edgell. Production of all-female and sterile Coho salmon, and experimental evidence for male heterogamety. Transactions of the American Fisheries Society, 111(3):367–372, May 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hansel:1988:UDB

[HDLG88] Hal C. Hansel, Stephen D. Duke, Peter T. Lofy, and Gerard A. Gray. Use of diagnostic bones to identify and estimate original lengths of ingested prey fishes. Transactions of the American Fisheries Society, 117(1):55–62, January 1988. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hallerman:1986:SDI

[HDS86] Eric M. Hallerman, R. A. Dunham, and R. O. Smitherman. Selection or drift–isozyme allele frequency changes among channel catfish selected for rapid growth. Transactions of the American Fisheries Society, 115(1):60–68, January 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hansen:1984:AGS

[HE84] Richard A. Hansen and Arnold G. Eversole. Age, growth, and sex ratio of American eels in brackish-water portions of a South Carolina River. Transactions of the American Fisheries Society, 113(6):744–749, November 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Heggenes:1988:EST

[Heg88] Jan Heggenes. Effects of short-term flow fluctuations on displacement of, and habitat use by, brown trout in a small stream. Transactions of the American Fisheries Society, 117(4):336–344, July 1988. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hazen:1981:AAA

[HER81] Terry C. Hazen, Gerald W. Esch, and Mark L. Raker. Agglutinating antibody to Aeromonas hydrophila in wild largemouth bass. Transactions of the American Fisheries Society, 110(4):514–518, July 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hermes:1985:DNL

[Her85] Rudolf Hermes. Distribution of neustonic larvae of hakes Urophycis spp. and fourbeard rockling Enchelyopus cimbrius in the Georges Bank Area. Transactions of the American Fisheries Society, 114(4):604–608, July 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hettler:1984:MOI

[Het84] William F. Hettler. Marking otoliths by immersion of marine fish larvae in tetracycline. Transactions of the American Fisheries Society, 113(3):370–373, May 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Henderson:1983:EHF

[HFDL83] V. Henderson, J. W. Fisher, R. D’Allessandris, and J. M. Livingston. Effects of hydrazine on functional morphology of rainbow trout embryos and larvae. Transactions of the American Fisheries Society, 112(1):100–104, January 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hecky:1981:RBP

[HFKR81] R. E. Hecky, E. J. Fee, H. J. Kling, and J. W. M. Rudd. Relationship between primary production and fish production in Lake Tanganyika. Transactions of the American Fisheries Society, 110(3):336–345, May 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hightower:1984:UJS

[HG84] Joseph E. Hightower and Ronnie J. Gilbert. Using the jolly-seber model to estimate population size, mortality, and recruitment for a reservoir fish population. Transactions of the American Fisheries Society, 113(5):633–641, September 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Harris:1985:GMA

[HG85] Michael J. Harris and Gary D. Grossman. Growth, mortality, and age composition of a lightly exploited tilefish substock off Georgia. Transactions of the American Fisheries Society, 114(6): 837–846, November 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hughes:1987:LCF

[HG87] Robert M. Hughes and James R. Gammon. Longitudinal changes in fish assemblages and water quality in the Willamette River, Oregon. Transactions of the American Fisheries Society, 116(2): 196–209, March 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hop:1989:GRA

[HG89] Haakon Hop and A. J. Gharrett. Genetic relationships of Arctic grayling in the koyukuk and tanana rivers, Alaska. Transactions of the American Fisheries Society, 118(3):290–295, May 1989. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hillman:1987:SWH

[HGP87] T. W. Hillman, J. S. Griffith, and W. S. Platts. Summer and winter habitat selection by juvenile Chinook salmon in a highly sedimented Idaho stream. Transactions of the American Fisheries Society, 116(2):185–195, March 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Harrison:1982:PEB

[HH82a] Edward J. Harrison and Wayne F. Hadley. Possible effects of black-spot disease on northern pike. Transactions of the American Fisheries Society, 111(1):106–109, January 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Howe:1982:MJB

[HH82b] Nathan R. Howe and Peter R. Hoyt. Mortality of juvenile brown shrimp Penaeus aztecus associated with streamer tags. Transactions of the American Fisheries Society, 111(3):317–325, May 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Howells:1982:IAP

[HH82c] Gwyneth Howells and Terry A. Haines. Interpretation of aquatic pH trends. Transactions of the American Fisheries Society, 111 (6):779–786, November 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hoenig:1987:ULL

[HH87] John Maurice Hoenig and Dennis M. Heisey. Use of a log-linear model with the EM algorithm to correct estimates of stock composition and to convert length to age. Transactions of the American Fisheries Society, 116(2):232–243, March 1987. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hatch:1981:EAB

[HHB81a] Richard W. Hatch, Paul M. Haack, and Edward H. Brown, Jr. Estimation of alewife biomass in Lake Michigan, 1967–1978. Transactions of the American Fisheries Society, 110(5):575–584, September 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hodson:1981:FHY

[HHB81b] Ronald G. Hodson, John O. Hackman, and Charles R. Bennett. Food habits of young spots in nursery areas of the Cape Fear River Estuary, North Carolina. Transactions of the American Fisheries Society, 110(4):495–501, July 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hemmingsen:1986:SPC

[HHEM86] A. R. Hemmingsen, R. A. Holt, R. D. Ewing, and J. D. McIntyre. Susceptibility of progeny from crosses among three stocks of Coho salmon to infection by Ceratomyxa shasta. Transactions of the American Fisheries Society, 115(3):492–495, May 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hurley:1987:HMS

[HHN87] Stephen T. Hurley, Wayne A. Hubert, and John G. Nickum. Habitats and movements of shovelnose sturgeons in the Upper Mississippi River. Transactions of the American Fisheries Society, 116(4):655–662, July 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hickey:1982:CHW

[Hic82] Clarence R. Hickey, Jr. Comparative hematology of wild and captive cunners. Transactions of the American Fisheries Society, 111(2):242–249, March 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Higgs:1987:CVP

[Hig87] Eric S. Higgs. Changing value perspectives in natural resource allocation: From market to ecosystem. Transactions of the American Fisheries Society, 116(3):525–531, May 1987. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hilborn:1988:DTR

[Hil88] Ray Hilborn. Determination of tag return from recaptured fish by sequential examination for tags. Transactions of the American Fisheries Society, 117(5):510–514, September 1988. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hilborn:1989:SPM

[Hil89] Ray Hilborn. Salmon production, management, and allocation. Transactions of the American Fisheries Society, 118(1):90–93, January 1989. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hinshaw:1985:EIP

[Hin85] Jeffrey M. Hinshaw. Effects of illumination and prey contrast on survival and growth of larval yellow perch Perca flavescens. Transactions of the American Fisheries Society, 114(4):540–545, July 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Haraldstad:1983:ASS

[HJ83] ØRnulf Haraldstad and Bror Jonsson. Age and sex segregation in habitat utilization by brown trout in a Norwegian lake. Transactions of the American Fisheries Society, 112(1):27–37, January 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hjelmeland:1985:ERM

[HJ85] Knut Hjelmeland and Trond Jørgensen. Evaluation of radioimmunoassay as a method to quantify trypsin and trypsinogen in fish. Transactions of the American Fisheries Society, 114(4):619– 621, July 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Holt:1983:DYR

[HKA83] Scott A. Holt, Christopher L. Kitting, and Connie R. Arnold. Distribution of young red drums among different sea-grass meadows. Transactions of the American Fisheries Society, 112(2):267–271, March 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Herman:1985:GLD

[HM85] Roger L. Herman and J. W. Meade. Gill lamellar dilations (Telangiectasis) related to sampling techniques. Transactions of the American Fisheries Society, 114(6):911–913, November 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hamilton:1986:MFR

[HM86] Steven J. Hamilton and Paul M. Mehrle. Metallothionein in fish: Review of its importance in assessing stress from metal contaminants. Transactions of the American Fisheries Society, 115(4): 596–609, July 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hayward:1987:EEP

[HM87a] Robert S. Hayward and F. Joseph Margraf. Eutrophication effects on prey size and food available to yellow perch in Lake Erie. Transactions of the American Fisheries Society, 116(2):210–223, March 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hedrick:1987:PTS

[HM87b] R. P. Hedrick and T. McDowell. Passive transfer of sera with antivirus neutralizing activity from adult channel catfish protects juveniles from channel catfish virus disease. Transactions of the American Fisheries Society, 116(2):277–281, March 1987. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hamilton:1987:CST

[HMJ87a] Steven J. Hamilton, Paul M. Mehrle, and John R. Jones. Cadmium-saturation technique for measuring metallothionein in brook trout. Transactions of the American Fisheries Society, 116(4):541–550, July 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hamilton:1987:EMM

[HMJ87b] Steven J. Hamilton, Paul M. Mehrle, and John R. Jones. Evaluation of metallothionein measurement as a biological indicator of stress from cadmium in brook trout. Transactions of the American Fisheries Society, 116(4):551–560, July 1987. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hamilton:1981:MPB

[HMMJ81a] Steven J. Hamilton, Paul M. Mehrle, Foster L. Mayer, and John R. Jones. Mechanical properties of bone in channel catfish as affected by vitamin C and toxaphene. Transactions of the American Fisheries Society, 110(6):718–724, November 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hamilton:1981:MEM

[HMMJ81b] Steven J. Hamilton, Paul M. Mehrle, Foster L. Mayer, and John R. Jones. Method to evaluate mechanical properties of bone in fish. Transactions of the American Fisheries Society, 110(6): 708–717, November 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hawkes:1986:PPB

[HMSK86] Clayton R. Hawkes, Theodore R. Meyers, Thomas C. Shirley, and Timothy M. Koeneman. Prevalence of the parasitic barnacle Briarosaccus callosus on king crabs of Southeastern Alaska.

Transactions of the American Fisheries Society, 115(2):252–257, March 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Horn:1984:PNA

[HN84] Michael H. Horn and Margaret A. Neighbors. Protein and nitrogen assimilation as a factor in predicting the seasonal macroalgal diet of the monkeyface prickleback. Transactions of the American Fisheries Society, 113(3):388–396, May 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Howick:1983:PFB

[HO83] Gregory L. Howick and W. John O’Brien. Piscivorous feeding behavior of largemouth bass: an experimental analysis. Transactions of the American Fisheries Society, 112(4):508–516, July 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hoehn:1987:CVF

[Hoe87] John P. Hoehn. Contingent valuation in fisheries management: The design of satisfactory contingent valuation formats. Transactions of the American Fisheries Society, 116(3):412–419, May 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Holland:1986:EBT

[Hol86] Leslie E. Holland. Effects of barge traffic on distribution and survival of ichthyoplankton and small fishes in the Upper Mississippi River. Transactions of the American Fisheries Society, 115 (1):162–165, January 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hontela:1984:DCS

[Hon84] Alice Hontela. Daily cycles of serum gonadotropin hormone in fish. Transactions of the American Fisheries Society, 113(4):458– 466, July 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hulsman:1983:MWE

[HPG83] Peter F. Hulsman, Perce M. Powles, and John M. Gunn. Mortality of walleye eggs and rainbow trout yolk-sac larvae in LowpH waters of the LaCloche mountain area, Ontario. Transactions of the American Fisheries Society, 112(5):680–688, September 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hall:1985:MSB

[HPHF85] Lenwood W. Hall, Jr., Alfred E. Pinkney, Larry O. Horseman, and Susan E. Finger. Mortality of striped bass larvae in relation to contaminants and water quality in a Chesapeake Bay tributary. Transactions of the American Fisheries Society, 114(6):861–868, November 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Heath:1987:TGD

[HR87] Daniel Heath and Derek A. Roff. Test of genetic differentiation in growth of stunted and nonstunted populations of yellow perch and pumpkinseed. Transactions of the American Fisheries Society, 116(1):98–102, January 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hamilton:1988:BDC

[HR88] Steven J. Hamilton and Rob J. Reash. Bone development in creek chub from a stream chronically polluted with heavy metals. Transactions of the American Fisheries Society, 117(1):48– 54, January 1988. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hughes:1983:DRE

[HRN83] Steven G. Hughes, Gary L. Rumsey, and Malden C. Nesheim. Dietary requirements for essential branched-chain amino acids by lake trout. Transactions of the American Fisheries Society, 112 (6):812–817, November 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Holland:1983:DLF

[HS83] L. E. Holland and J. R. Sylvester. Distribution of larval fishes related to potential navigation impacts on the Upper Mississippi River, pool 7. Transactions of the American Fisheries Society, 112(2):293–301, March 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hewett:1989:ZAL

[HS89] Steven W. Hewett and Donald J. Stewart. Zooplanktivory by alewives in Lake Michigan: Ontogenetic, seasonal, and historical patterns. Transactions of the American Fisheries Society, 118 (6):581–596, November 1989. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Helfman:1983:UTA

[HSB^{+}83] G. S. Helfman, D. L. Stoneburner, E. L. Bozeman, P. A. Christian, and R. Whalen. Ultrasonic telemetry of American eel movements in a tidal creek. Transactions of the American Fisheries Society, 112(1):105–110, January 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hallerman:1988:EDG

[HSG^{+}88] Eric M. Hallerman, John F. Schneider, Mark L. Gross, Anthony J. Faras, Perry B. Hackett, Kevin S. Guise, and Anne R. Kapuscinski. Enzymatic dechorionation of goldfish, walleye, and northern pike eggs. Transactions of the American Fisheries Society, 117 (5):456–460, September 1988. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hoyer:1985:UOD

[HSM85] Mark V. Hoyer, Jerome V. Shireman, and Michael J. Maceina. Use of otoliths to determine age and growth of largemouth bass in Florida. Transactions of the American Fisheries Society, 114 (2):307–309, March 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hallerman:1986:BGL

[HSR^{+}86] Eric M. Hallerman, R. O. Smitherman, Russell B. Reed, William H. Tucker, and Rex A. Dunham. Biochemical genetics of largemouth bass in mesosaline and freshwater areas of the Alabama River system. Transactions of the American Fisheries Society, 115(1):15–20, January 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Heming:1985:ATT

[HTMZ85] Thomas A. Heming, Robert V. Thurston, Elizabeth L. Meyn, and Richard K. Zajdel. Acute toxicity of thiocyanate to trout. Transactions of the American Fisheries Society, 114(6):895–905, November 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hudgins:1984:SAE

[Hud84] Michael D. Hudgins. Structure of the angling experience. Transactions of the American Fisheries Society, 113(6):750–759, November 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hushak:1987:UIO

[Hus87] Leroy J. Hushak. Use of input-output analysis in fisheries assessment. Transactions of the American Fisheries Society, 116(3): 441–449, May 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hansen:1981:SSD

[HW81] Michael J. Hansen and David H. Wahl. Selection of small Daphnia pulex by yellow perch fry in Oneida Lake, New York. Transactions of the American Fisheries Society, 110(1):64–71, January 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hushak:1988:EVG

[HWD88] Leroy J. Hushak, Jane M. Winslow, and Nilima Dutta. Economic value of great lakes sportfishing: The case of private-boat fishing in Ohio’s Lake Erie. Transactions of the American Fisheries Society, 117(4):363–373, July 1988. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hatfield:1982:EAG

[HWGF82] Jeff S. Hatfield, Thomas E. Wissing, Sheldon I. Guttman, and Michael P. Farrell. Electrophoretic analysis of gizzard shad from the Lower Mississippi River and Ohio. Transactions of the American Fisheries Society, 111(6):742–748, November 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hara:1989:SFD

[HZ89] Toshiaki J. Hara and Barbara Zielinski. Structural and functional development of the olfactory organ in teleosts. Transactions of the American Fisheries Society, 118(2):183–194, March 1989. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hall:1989:SCW

[HZB^{+}89] Lenwood W. Hall, Jr., Michael C. Ziegenfuss, Steven J. Bushong, Michael A. Unger, and Roger L. Herman. Studies of contaminant and water quality effects on striped bass prolarvae and yearlings in the Potomac River and Upper Chesapeake Bay in 1988. Transactions of the American Fisheries Society, 118(6):619–629, November 1989. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Inskip:1983:CFP

[IM83] Peter D. Inskip and John J. Magnuson. Changes in fish populations over an 80-year period: Big Pine Lake, Wisconsin. Transactions of the American Fisheries Society, 112(3):378–389, May 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Irvine:1982:SSF

[IN82] J. R. Irvine and T. G. Northcote. Significance of sequential feeding patterns of juvenile rainbow trout in a large lake-Fed River. Transactions of the American Fisheries Society, 111(4):446–452, July 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Isely:1987:SPF

[INKP87] J. Jeffery Isely, Richard L. Noble, Jeffrey B. Koppelman, and David P. Philipp. Spawning period and first-year growth of northern, Florida, and intergrade stocks of largemouth bass. Transactions of the American Fisheries Society, 116(5):757–762, September 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ingram:1983:DSB

[IZ83] William Ingram and Charles D. Ziebell. Diet shifts to benthic feeding by threadfin shad. Transactions of the American Fisheries Society, 112(4):554–556, July 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jordahl:1987:ELP

[JB87] David M. Jordahl and Arnold Benson. Effect of low pH on survival of brook trout embryos and yolk-sac larvae in West Virginia streams. Transactions of the American Fisheries Society, 116(6): 807–816, November 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Johnson:1988:SHS

[JBL88] David L. Johnson, Raymond A. Beaumier, and William E. Lynch, Jr. Selection of habitat structure interstice size by bluegills and largemouth bass in ponds. Transactions of the American Fisheries Society, 117(2):171–179, March 1988. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jensen:1981:SSS

[Jen81] A. L. Jensen. Sample sizes for single mark and single recapture experiments. Transactions of the American Fisheries Society, 110(3):455–458, May 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jensen:1984:LSP

[Jen84] A. L. Jensen. Logistic surplus-production model with explicit terms for growth, mortality, and recruitment. Transactions of the American Fisheries Society, 113(5):617–626, September 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jensen:1985:CCC

[Jen85] A. L. Jensen. Comparison of catch-curve methods for estimation of mortality. Transactions of the American Fisheries Society, 114 (5):743–747, September 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jessop:1987:MAE

[Jes87] B. M. Jessop. Migrating American eels in Nova Scotia. Transactions of the American Fisheries Society, 116(2):161–170, March 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jones:1982:SHP

[JH82] John R. Jones and Mark V. Hoyer. Sportfish harvest predicted by summer Chlorophyll-α concentration in Midwestern lakes and reservoirs. Transactions of the American Fisheries Society, 111 (2):176–179, March 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jagoe:1983:AGE

[JH83] Charles H. Jagoe and Terry A. Haines. Alterations in gill epithelial morphology of yearling Sunapee trout exposed to acute acid stress. Transactions of the American Fisheries Society, 112(5): 689–695, September 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Johnson:1982:CCF

[JM82] David L. Johnson and Michael T. Metcalf. Causes and controls of freshwater drum mortality during transportation. Transactions of the American Fisheries Society, 111(1):58–62, January 1982. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jenkins:1982:MIC

[JO82] Robert M. Jenkins and Ray T. Oglesby. The morphoedaphic index — concepts and practices: The morphoedaphic index and reservoir fish production. Transactions of the American Fisheries Society, 111(2):133–140, March 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Johnson:1983:AGY

[Joh83a] Allyn G. Johnson. Age and growth of yellowtail snapper from South Florida. Transactions of the American Fisheries Society, 112(2):173–177, March 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Johnston:1983:SCG

[Joh83b] C. E. Johnston. Seasonal changes in gill (Na+K)-ATPase activity in Atlantic salmon retained in fresh water after smolting. Transactions of the American Fisheries Society, 112(5):720–724, September 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Johnson:1989:MFS

[Joh89] Murray G. Johnson. Metals in fish scales collected in Lake Opeongo, Canada, from 1939 to 1979. Transactions of the American Fisheries Society, 118(3):331–335, May 1989. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jonsson:1985:LHP

[Jon85] Bror Jonsson. Life history patterns of freshwater resident and searun migrant brown trout in Norway. Transactions of the American Fisheries Society, 114(2):182–194, March 1985. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jee:1981:EOL

[JP81] Lau Kui Jee and John A. Plumb. Effects of organic load on potassium permanganate as a treatment for Flexibacter columnaris. Transactions of the American Fisheries Society, 110(1): 86–89, January 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Johnson:1988:UQR

[JSC88] Brett M. Johnson, Roy A. Stein, and Robert F. Carline. Use of a quadrat rotenone technique and bioenergetics modeling to evaluate prey availability to stocked piscivores. Transactions of the American Fisheries Society, 117(2):127–141, March 1988. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jones:1981:BIT

[JST^{+}81] John R. Jones, Miles M. Smart, Bryn H. Tracy, Donald H. Hazelwood, and Jeanne L. Sebaugh. Biotic index tested for ability to assess water quality of Missouri Ozark streams. Transactions of the American Fisheries Society, 110(5):627–637, September 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jensen:1982:UCF

[JST82] A. L. Jensen, S. A. Spigarelli, and M. M. Thommes. Use of conventional fishery models to assess entrainment and impingement of three Lake Michigan fish species. Transactions of the American Fisheries Society, 111(1):21–34, January 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jensen:1983:SRG

[JSYW83] Gary L. Jensen, William L. Shelton, Shiu-Lin Yang, and L. O. Wilken. Sex reversal of gynogenetic grass carp by implantation of methyltestosterone. Transactions of the American Fisheries Society, 112(1):79–85, January 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jude:1987:DSM

[JTDM87] David J. Jude, Frank J. Tesar, Scott F. Deboe, and Timothy J. Miller. Diet and selection of major prey species by Lake Michigan salmonines, 1973–1982. Transactions of the American Fisheries Society, 116(5):677–691, September 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Johnson:1987:FFB

[JV87] Philip C. Johnson and Gary L. Vinyard. Filter-feeding behavior and particle retention efficiency of Sacramento blackfish. Transactions of the American Fisheries Society, 116(4):634–640, July 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kavaliers:1984:OPP

[Kav84] Martin Kavaliers. Opioid peptides, the pineal gland, and rhythmic behavior in fish. Transactions of the American Fisheries Society, 113(4):432–438, July 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kaya:1989:RYA

[Kay89] Calvin M. Kaya. Rheotaxis of young Arctic grayling from populations that spawn in inlet or outlet streams of a lake. Transactions of the American Fisheries Society, 118(5):474–481, September 1989. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Koslow:1985:AAY

[KBDP85] J. A. Koslow, S. Brault, J. Dugas, and F. Page. Anatomy of an apparent year-class failure: The early life history of the 1983 browns bank haddock Melanogrammus aeglefinus. Transactions of the American Fisheries Society, 114(4):478–489, July 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kerby:1985:GSH

[KBH85] Jerome Howard Kerby, Jack D. Bayless, and Reginal M. Harrell. Growth, survival, and harvest of striped bass produced with cryopreserved spermatozoa. Transactions of the American Fisheries Society, 114(5):761–765, September 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kunz:1989:EFT

[KC89] Yvette W. Kunz and E. Callaghan. Embryonic fissures in teleost eyes and their possible role in detection of polarized light. Transactions of the American Fisheries Society, 118(2):195–202, March 1989. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Keast:1985:GDY

[KE85] Allen Keast and John McA. Eadie. Growth depensation in year-0 largemouth bass: The influence of diet. Transactions of the American Fisheries Society, 114(2):204–213, March 1985. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kendall:1981:DPS

[Ken81] Robert L. Kendall. Dual publication of scientific information. Transactions of the American Fisheries Society, 110(5):573–574, September 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kendall:1989:NEP

[Ken89] Robert L. Kendall. New editorial procedures for Transactions. Transactions of the American Fisheries Society, 118(5):461–462, September 1989. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kerr:1982:REA

[Ker82] S. R. Kerr. The role of external analysis in fisheries science. Transactions of the American Fisheries Society, 111(2):165–170, March 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kerby:1983:CPS

[Ker83] Jerome Howard Kerby. Cryogenic preservation of sperm from striped bass. Transactions of the American Fisheries Society, 112(1):86–94, January 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kellogg:1983:RBO

[KG83] Robert L. Kellogg and James J. Gift. Relationship between optimum temperatures for growth and preferred temperatures for the young of four fish species. Transactions of the American Fisheries Society, 112(3):424–430, May 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kim:1989:CDW

[KG89] Suam Kim and Donald R. Gunderson. Cohort dynamics of walleye Pollock in Shelikof Strait, Gulf of Alaska, during the egg and larval periods. Transactions of the American Fisheries Society, 118(3):264–273, May 1989. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kernehan:1981:ELH

[KHS81] Ronnie J. Kernehan, Michael R. Headrick, and Reinald E. Smith. Early life history of striped bass in the Chesapeake and Delaware Canal and vicinity. Transactions of the American Fisheries Society, 110(1):137–150, January 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kirby:1981:SOE

[Kir81] John M. Kirby. Seasonal occurrence of the ectoparasite Gyrodactylus atratuli on spotfin shiners. Transactions of the American Fisheries Society, 110(3):462–464, May 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kellogg:1984:TMP

[KLJ84] Robert L. Kellogg, Rebecca J. Ligotino, and Steven M. Jinks. Thermal mortality prediction equations for entrainable striped bass. Transactions of the American Fisheries Society, 113(6): 794–802, November 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Koenings:1986:QDO

[KLM86] J. P. Koenings, J. Lipton, and P. McKay. Quantitative determination of oxytetracycline uptake and release by juvenile sockeye salmon. Transactions of the American Fisheries Society, 115(4): 621–629, July 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kashuba:1984:PCL

[KM84] Stephen A. Kashuba and William J. Matthews. Physical condition of larval shad during spring-summer in a southwestern reservoir. Transactions of the American Fisheries Society, 113 (2):199–204, March 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Krueger:1987:GCN

[KM87a] Charles C. Krueger and Bernie May. Genetic comparison of naturalized rainbow trout populations among Lake Superior tributaries: Differentiation based on allozyme data. Transactions of the American Fisheries Society, 116(6):795–806, November 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Krueger:1987:SIN

[KM87b] Charles C. Krueger and Bernie May. Stock identification of naturalized brown trout in Lake Superior tributaries: Differentiation based on allozyme data. Transactions of the American Fisheries Society, 116(6):785–794, November 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Knight:1984:PWY

[KMC84] Roger L. Knight, F. Joseph Margraf, and Robert F. Carline. Piscivory by walleyes and yellow perch in Western Lake Erie. Transactions of the American Fisheries Society, 113(6):677–693, November 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Killgore:1987:ETY

[KMC87] K. Jack Killgore, Andrew C. Miller, and Kenneth C. Conley. Effects of turbulence on yolk-sac larvae of paddlefish. Transactions of the American Fisheries Society, 116(4):670–673, July 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Klein-Macphee:1984:ESE

[KMCB84] Grace Klein-Macphee, John A. Cardin, and Walter J. Berry. Effects of silver on eggs and larvae of the winter flounder. Transactions of the American Fisheries Society, 113(2):247–251, March 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Krueger:1989:GDA

[KMKM89] Charles C. Krueger, J. Ellen Marsden, Harold L. Kincaid, and Bernie May. Genetic differentiation among lake trout strains stocked into Lake Ontario. Transactions of the American Fisheries Society, 118(3):317–330, May 1989. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kobayashi:1984:GBE

[KMTU84] Tokimasa Kobayashi, George B. Milner, David J. Teel, and Fred M. Utter. Genetic basis for electrophoretic variation of adenosine deaminase in Chinook salmon. Transactions of the American Fisheries Society, 113(1):86–89, January 1984. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kohler:1981:CAF

[KN81] Christopher C. Kohler and John J. Ney. Consequences of an alewife die-off to fish and zooplankton in a reservoir. Transactions of the American Fisheries Society, 110(3):360–369, May 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Knutson:1983:FRA

[KO83] Arthur C. Knutson, Jr. and James J. Orsi. Factors regulating abundance and distribution of the shrimp Neomysis mercedis in the sacramento-san joaquin estuary. Transactions of the American Fisheries Society, 112(4):476–485, July 1983. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kohlenstein:1981:PCB

[Koh81] Lawrence C. Kohlenstein. On the proportion of the Chesapeake Bay stock of striped bass that migrates into the coastal fishery. Transactions of the American Fisheries Society, 110(1):168–179, January 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Koltes:1985:ESC

[Kol85] Karen H. Koltes. Effects of sublethal copper concentrations on the structure and activity of Atlantic silverside schools. Transactions of the American Fisheries Society, 114(3):413–422, May 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Klauda:1987:RBH

[KP87] Ronald J. Klauda and Robert E. Palmer. Responses of blueback herring eggs and larvae to pulses of acid and aluminum. Transactions of the American Fisheries Society, 116(4):561–569, July 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kwain:1985:GBT

[KR85] W. Kwain and G. A. Rose. Growth of brook trout Salvelinus fontinalis subject to sudden reductions of pH during their early life history. Transactions of the American Fisheries Society, 114(4):

564–570, July 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kremers:1986:BR

[Kre86a] Walter K. Kremers. Book reviews. Transactions of the American Fisheries Society, 115(3):500–502, May 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kremers:1986:SHP

[Kre86b] Walter K. Kremers. Subsample heterogeneity and prey selection indices. Transactions of the American Fisheries Society, 115(2): 356–358, March 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Krueger:1988:FFL

[Kru88] William H. Krueger. Fishes: a field and laboratory manual on their structure, identification, and natural history. Transactions of the American Fisheries Society, 117(2):218–220, March 1988. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kircheis:1981:TPF

[KS81] Frederick W. Kircheis and Jon G. Stanley. Theory and practice of forage-fish management in New England. Transactions of the American Fisheries Society, 110(6):729–737, November 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Konkle:1986:PSL

[KS86] Blake R. Konkle and W. Gary Sprules. Planktivory by stunted lake trout in an Ontario lake. Transactions of the American Fisheries Society, 115(4):515–521, July 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Knutsen:1985:GDF

[KT85] G. M. Knutsen and S. Tilseth. Growth, development, and feeding success of Atlantic cod larvae Gadus morhua related to egg size. Transactions of the American Fisheries Society, 114(4):507–511, July 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kocik:1987:DMA

[KT87] John F. Kocik and William W. Taylor. Diet and movements of age-1+ pink salmon in Western Lake Huron. Transactions of the American Fisheries Society, 116(4):628–633, July 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kruse:1989:ESE

[KT89] Gordon H. Kruse and Albert V. Tyler. Exploratory simulation of English sole recruitment mechanisms. Transactions of the American Fisheries Society, 118(2):101–118, March 1989. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kreuz:1982:VGD

[KTKD82] K. F. Kreuz, A. V. Tyler, G. H. Kruse, and R. L. Demory. Variation in growth of Dover soles and English soles as related to upwelling. Transactions of the American Fisheries Society, 111 (2):180–192, March 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kushlan:1981:SCE

[Kus81] James A. Kushlan. Sampling characteristics of enclosure fish traps. Transactions of the American Fisheries Society, 110(4): 557–562, July 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kawamura:1989:OCB

[KW89] Gunzo Kawamura and Naoki Washiyama. Ontogenetic changes in behavior and sense organ morphogenesis in largemouth bass and Tilapia nilotica. Transactions of the American Fisheries Society, 118(2):203–213, March 1989. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kwain:1981:PDE

[Kwa81] Wen-Hwa Kwain. Population dynamics and exploitation of rainbow trout in Stokely Creek, Eastern Lake Superior. Transactions of the American Fisheries Society, 110(2):210–215, March 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

King:1981:FVR

[KWDS81] Terry A. King, J. C. Williams, William D. Davies, and William L. Shelton. Fixed versus random sampling of fishes in a large reservoir. Transactions of the American Fisheries Society, 110(4):563– 568, July 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Koppelman:1988:TPN

[KWP88] Jeffrey B. Koppelman, Gregory S. Whitt, and David P. Philipp. Thermal preferenda of northern, Florida, and reciprocal F_{1} hybrid largemouth bass. Transactions of the American Fisheries Society, 117(3):238–244, May 1988. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Karr:1987:STV

[KYFS87] James R. Karr, Philip R. Yant, Kurt D. Fausch, and Isaac J. Schlosser. Spatial and temporal variability of the index of biotic integrity in three Midwestern streams. Transactions of the American Fisheries Society, 116(1):1–11, January 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kaeding:1983:LHE

[KZ83] Lynn R. Kaeding and Marian A. Zimmerman. Life history and ecology of the humpback chub in the little Colorado and Colorado Rivers of the grand canyon. Transactions of the American Fisheries Society, 112(5):577–594, September 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Luey:1984:SSR

[LA84] James E. Luey and Ira R. Adelman. Stock structure of rainbow smelt in Western Lake Superior: Population characteristics. Transactions of the American Fisheries Society, 113(6):709–715, November 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lackey:1989:EOE

[Lac89] Robert T. Lackey. Ecology and our endangered life-support systems. Transactions of the American Fisheries Society, 118(6): 721–724, November 1989. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Leary:1985:DII

[LAK85] Robb F. Leary, Fred W. Allendorf, and Kathy L. Knudsen. Developmental instability as an indicator of reduced genetic variation in hatchery trout. Transactions of the American Fisheries Society, 114(2):230–235, March 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Leary:1987:GDI

[LAPK87] Robb F. Leary, Fred W. Allendorf, Stevan R. Phelps, and Kathy L. Knudsen. Genetic divergence and identification of seven cutthroat trout subspecies and rainbow trout. Transactions of the American Fisheries Society, 116(4):580–587, July 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Leary:1982:GSA

[LB82] Robb Leary and Henry E. Booke. Genetic stock analysis of yellow perch from Green Bay and Lake Michigan. Transactions of the American Fisheries Society, 111(1):52–57, January 1982. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Linder:1984:PDA

[LB84] Gregory Linder and Harold L. Bergman. Periodic depuration of anthracene metabolites by rainbow trout. Transactions of the American Fisheries Society, 113(4):513–520, July 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lindberg:1986:EDS

[LD86] Joan C. Lindberg and Serge I. Doroshov. Effect of diet switch between natural and prepared foods on growth and survival of white sturgeon juveniles. Transactions of the American Fisheries Society, 115(1):166–171, January 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Leider:1985:PTU

[Lei85] Steven A. Leider. Precise timing of upstream migrations by repeat steelhead spawners. Transactions of the American Fisheries Society, 114(6):906–908, November 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lemly:1985:SNF

[Lem85] A. Dennis Lemly. Suppression of native fish populations by green sunfish in first-order streams of Piedmont North Carolina. Transactions of the American Fisheries Society, 114(5):705–712, September 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Labar:1983:LMI

[LF83] George W. Labar and Douglas E. Facey. Local movements and inshore population sizes of American eels in Lake Champlain, Vermont. Transactions of the American Fisheries Society, 112 (1):111–116, January 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lorenz:1989:SHR

[LF89] J. Mitchel Lorenz and John H. Filer. Spawning habitat and redd characteristics of sockeye salmon in the glacial Taku River, British Columbia and Alaska. Transactions of the American Fisheries Society, 118(5):495–502, September 1989. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Limsuwan:1983:EAF

[LGP83] Chalor Limsuwan, John M. Grizzle, and John A. Plumb. Etomidate as an anesthetic for fish: Its toxicity and efficacy. Transactions of the American Fisheries Society, 112(4):544–550, July 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lee:1984:GRD

[LHA84] W. Y. Lee, G. J. Holt, and C. R. Arnold. Growth of red drum larvae in the laboratory. Transactions of the American Fisheries Society, 113(2):243–246, March 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lanka:1987:RGS

[LHW87] Robert P. Lanka, Wayne A. Hubert, and Thomas A. Wesche. Relations of geomorphology to stream habitat and trout standing stock in small rocky mountain streams. Transactions of the American Fisheries Society, 116(1):21–28, January 1987. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Larson:1986:BPN

[LJL86] Eric W. Larson, David L. Johnson, and William E. Lynch, Jr. A buoyant pop net for accurately sampling fish at artificial habitat structures. Transactions of the American Fisheries Society, 115 (2):351–355, March 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Loesch:1982:ELI

[LKF82] Joseph G. Loesch, William H. Kriete, Jr., and Eric J. Foell. Effects of light intensity on the catchability of juvenile anadromous Alosa species. Transactions of the American Fisheries Society, 111(1):41–44, January 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lovell:1982:ISD

[LL82] Richard T. Lovell and Tasanee Limsuwan. Intestinal synthesis and dietary nonessentiality of vitamin B_{12} for Tilapia nilotica. Transactions of the American Fisheries Society, 111(4):485–490, July 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lovell:1986:GMM

[LLBA86] Richard T. Lovell, Iwan Y. Lelana, Claude E. Boyd, and Martin S. Armstrong. Geosmin and musty-muddy flavors in pond-raised channel catfish. Transactions of the American Fisheries Society, 115(3):485–489, May 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Li:1981:EAS

[LM81] Hiram W. Li and Peter B. Moyle. Ecological analysis of species introductions into aquatic systems. Transactions of the American Fisheries Society, 110(6):772–782, November 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Li:1982:CHM

[LM82] Sifa Li and Jack A. Mathias. Causes of high mortality among cultured larval walleyes. Transactions of the American Fisheries Society, 111(6):710–721, November 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Larson:1985:EER

[LM85a] Gary L. Larson and Stephen E. Moore. Encroachment of exotic rainbow trout into stream populations of native brook trout in the Southern Appalachian mountains. Transactions of the American Fisheries Society, 114(2):195–203, March 1985. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Layher:1985:RBH

[LM85b] William G. Layher and O. Eugene Maughan. Relations between habitat variables and channel catfish populations in prairie streams. Transactions of the American Fisheries Society, 114(6): 771–781, November 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lewis:1986:TNF

[LM86] William M. Lewis, Jr. and Donald P. Morris. Toxicity of nitrite to fish: a review. Transactions of the American Fisheries Society, 115(2):183–195, March 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lyons:1987:EWP

[LM87] John Lyons and John J. Magnuson. Effects of walleye predation on the population dynamics of small littoral-zone fishes in a Northern Wisconsin lake. Transactions of the American Fisheries Society, 116(1):29–39, January 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lawler:1988:ETW

[LM88] Sharon P. Lawler and Peter J. Morin. The ecology of temporary waters. Transactions of the American Fisheries Society, 117(4): 417–420, July 1988. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Liang:1981:PPF

[LMW81] Yanling Liang, John M. Melack, and Ji Wang. Primary production and fish yields in Chinese ponds and lakes. Transactions of the American Fisheries Society, 110(3):346–350, May 1981. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Leaman:1987:AVR

[LN87] B. M. Leaman and D. A. Nagtegaal. Age validation and revised natural mortality rate for yellowtail rockfish. Transactions of the American Fisheries Society, 116(2):171–175, March 1987. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Leefers:1987:MUR

[LNK87] Larry A. Leefers, Frederick L. Norbury, and H. Fred Kaiser, Jr. Multiple-use resource economics: Theory and planning applications. Transactions of the American Fisheries Society, 116(3):

367–373, May 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Leslie:1983:EVC

[LNV83] Andrew J. Leslie, Jr., Larry E. Nall, and Jess M. Van Dyke. Effects of vegetation control by grass carp on selected water-quality variables in four Florida lakes. Transactions of the American Fisheries Society, 112(6):777–787, November 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Luczkovich:1983:FBP

[LO83] J. J. Luczkovich and B. L. Olla. Feeding behavior, prey consumption, and growth of juvenile red hake. Transactions of the American Fisheries Society, 112(5):629–637, September 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Leonard:1986:ATI

[LO86] Paul M. Leonard and Donald J. Orth. Application and testing of an index of biotic integrity in small, coolwater streams. Transactions of the American Fisheries Society, 115(3):401–414, May 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Loftus:1987:IST

[Lof87] K. H. Loftus. Inadequate science transfer: an issue basic to effective fisheries management. Transactions of the American Fisheries Society, 116(3):314–319, May 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Longhurst:1988:PAU

[Lon88] Alan R. Longhurst. The Peruvian anchoveta and its upwelling ecosystem: Three decades of change. Transactions of the American Fisheries Society, 117(3):317–320, May 1988. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Luecke:1986:OCF

[Lue86] Chris Luecke. Ontogenetic changes in feeding habits of juvenile cutthroat trout. Transactions of the American Fisheries Society, 115(5):703–710, September 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Leslie:1982:CVT

[LVNM82] Andrew J. Leslie, Jr., Jesse M. Van Dyke, Larry E. Nall, and Woodard W. Miley II. Current velocity for transport of grass carp eggs. Transactions of the American Fisheries Society, 111 (1):99–101, January 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mabee:1989:AUU

[Mab89] Paula M. Mabee. Assumptions underlying the use of ontogenetic sequences for determining character state order. Transactions of the American Fisheries Society, 118(2):151–158, March 1989. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Macdonald:1982:CCR

[Mac82] Craig D. Macdonald. Catch composition and reproduction of the spiny lobster Panulirus versicolor at Palau. Transactions of the American Fisheries Society, 111(6):694–699, November 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Maceina:1989:NDO

[Mac89] Michael J. Maceina. Notes: Differences in otolith morphology among the two subspecies of largemouth bass and their F_{1} hybrid. Transactions of the American Fisheries Society, 118(5): 573–575, September 1989. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mansfield:1984:RLM

[Man84] Pamela J. Mansfield. Reproduction by Lake Michigan fishes in a tributary stream. Transactions of the American Fisheries Society, 113(2):231–237, March 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Marley:1983:SDP

[Mar83] Robert D. Marley. Spatial distribution patterns of planktonic fish eggs in Lower Mobile Bay, Alabama. Transactions of the American Fisheries Society, 112(2):257–266, March 1983. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Martinez:1984:IBB

[Mar84] Anita M. Martinez. Identification of brook, brown, rainbow, and cutthroat trout larvae. Transactions of the American Fisheries Society, 113(2):252–259, March 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Marliave:1985:ISE

[Mar85] Jeffrey B. Marliave. International Symposium on the Early Life History of Fishes and Eighth Annual Larval Fish Conference. Transactions of the American Fisheries Society, 114(4):443–444, July 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Marliave:1986:LPD

[Mar86] Jeffrey B. Marliave. Lack of planktonic dispersal of rocky intertidal fish larvae. Transactions of the American Fisheries Society, 115(1):149–154, January 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Marsh:1987:DTC

[Mar87a] Paul C. Marsh. Digestive tract contents of adult razorback suckers in Lake Mohave, Arizona–Nevada. Transactions of the American Fisheries Society, 116(1):117–119, January 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Martin:1987:EIA

[Mar87b] Larry R. G. Martin. Economic impact analysis of a sport fishery on Lake Ontario: an appraisal of method. Transactions of the American Fisheries Society, 116(3):461–468, May 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Matlock:1981:NRT

[Mat81] Gary C. Matlock. Nonreporting of recaptured tagged fish by saltwater recreational boat anglers in Texas. Transactions of the American Fisheries Society, 110(1):90–92, January 1981. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Matthews:1984:ITI

[Mat84] William J. Matthews. Influence of turbid inflows on vertical distribution of larval shad and freshwater drum. Transactions of the American Fisheries Society, 113(2):192–198, March 1984. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Matthews:1985:SMS

[Mat85] William J. Matthews. Summer mortality of striped bass in reservoirs of the United States. Transactions of the American Fisheries Society, 114(1):62–66, January 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McCauley:1982:TTA

[MB82] Robert W. McCauley and Fred P. Binkowski. Thermal tolerance of the alewife. Transactions of the American Fisheries Society, 111(3):389–391, May 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McGee:1983:EIW

[MB83] Michael V. McGee and Claude E. Boyd. Evaluation of the influence of water exchange in channel catfish ponds. Transactions of the American Fisheries Society, 112(4):557–560, July 1983. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Moyle:1985:MUA

[MB85] Peter B. Moyle and Donald M. Baltz. Microhabitat use by an assemblage of California stream fishes: Developing criteria for instream flow determinations. Transactions of the American Fisheries Society, 114(5):695–704, September 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Millemann:1984:CAT

[MBB^{+}84] Raymond E. Millemann, Wesley J. Birge, Jeffrey A. Black, Robert M. Cushman, Karen L. Daniels, Paul J. Franco, Jeffrey M. Giddings, John F. McCarthy, and Arthur J. Stewart. Comparative acute toxicity to aquatic organisms of components of coalderived synthetic fuels. Transactions of the American Fisheries Society, 113(1):74–85, January 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mills:1985:SVF

[MBC85] Chris A. Mills, William R. C. Beaumont, and Ralph T. Clarke. Sources of variation in the feeding of larval dace Leuciscus leuciscus in an English River. Transactions of the American Fisheries Society, 114(4):519–524, July 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Macfarlane:1986:EFM

[MBM86] Richard D. Macfarlane, Graham L. Bullock, and John J. A. McLaughlin. Effects of five metals on susceptibility of striped bass to Flexibacter columnaris. Transactions of the American Fisheries Society, 115(2):227–231, March 1986. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Minckley:1989:OBS

[MBM89] W. L. Minckley, Donald G. Buth, and Richard L. Mayden. Origin of brood stock and allozyme variation in hatchery-reared bonytail, an endangered North American cyprinid fish. Transactions of the American Fisheries Society, 118(2):131–137, March 1989. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Moles:1981:RGC

[MBRK81] Adam Moles, Steve Bates, Stanley D. Rice, and Sid Korn. Reduced growth of Coho salmon fry exposed to two petroleum components, toluene and naphthalene, in fresh water. Transactions of the American Fisheries Society, 110(3):430–436, May 1981. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McGurk:1985:ENC

[McG85] Michael D. McGurk. Effects of net capture on the postpreservation morphometry, dry weight, and condition factor of Pacific herring larvae. Transactions of the American Fisheries Society, 114(3):348–355, May 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McDonough:1987:RCS

[MCH87] Maureen H. McDonough, Marci Cobb, and Donald F. Holecek. Role of communication science in social valuation of fisheries. Transactions of the American Fisheries Society, 116(3):519–524, May 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McLaren:1981:MHR

[MCHL81] James B. McLaren, Jon C. Cooper, Thomas B. Hoff, and Vincent Lander. Movements of Hudson River striped bass. Transactions of the American Fisheries Society, 110(1):158–167, January 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mills:1986:ZSY

[MCK86] Edward L. Mills, John L. Confer, and Donald W. Kretchmer. Zooplankton selection by young yellow perch: The influence of light, prey density, and predator size. Transactions of the American Fisheries Society, 115(5):716–725, September 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mills:1984:PSY

[MCR84] Edward L. Mills, John L. Confer, and Richard C. Ready. Prey selection by young yellow perch: The influence of capture success, visual acuity, and prey choice. Transactions of the American Fisheries Society, 113(5):579–587, September 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Muncy:1981:MWE

[MD81] Robert J. Muncy and Arthur P. D’Silva. Marking walleye eggs and fry. Transactions of the American Fisheries Society, 110(2): 300–305, March 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mummert:1986:EFS

[MD86] John R. Mummert and Ray W. Drenner. Effect of fish size on the filtering efficiency and selective particle ingestion of a filterfeeding clupeid. Transactions of the American Fisheries Society, 115(4):522–528, July 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Middaugh:1984:RRA

[MDS84] Douglas P. Middaugh, Richard G. Domey, and Geoffrey I. Scott. Reproductive rhythmicity of the Atlantic silverside. Transactions of the American Fisheries Society, 113(4):472–478, July 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Meier:1984:TSC

[Mei84] Albert H. Meier. Temporal synergism of circadian neuroendocrine oscillations regulates seasonal conditions in the Gulf killifish. Transactions of the American Fisheries Society, 113(4):422–431, July 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Moring:1986:EIS

[MENG86] John R. Moring, Paul D. Eiler, Mary T. Negus, and Elizabeth Gibbs. Ecological importance of submerged pulpwood logs in a Maine reservoir. Transactions of the American Fisheries Society, 115(2):335–342, March 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Merna:1986:CSF

[Mer86] James W. Merna. Contamination of stream fishes with chlorinated hydrocarbons from eggs of great lakes salmon. Transactions of the American Fisheries Society, 115(1):69–74, January 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mesick:1988:EFC

[Mes88] Carl F. Mesick. Effects of food and cover on numbers of Apache and brown trout establishing residency in artificial stream channels. Transactions of the American Fisheries Society, 117(5):421– 431, September 1988. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mills:1981:EFC

[MF81] Edward L. Mills and John L. Forney. Energetics, food consumption, and growth of young yellow perch in Oneida Lake, New York. Transactions of the American Fisheries Society, 110(4):479–488, July 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mills:1983:IDP

[MF83] Edward L. Mills and John L. Forney. Impact on Daphnia pulex of predation by young yellow perch in Oneida Lake, New York. Transactions of the American Fisheries Society, 112(2):154–161, March 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Moore:1988:RYY

[MG88] Kelly M. S. Moore and Stan V. Gregory. Response of young-of-the-year cutthroat trout to manipulation of habitat structure in a small stream. Transactions of the American Fisheries Society, 117(2):162–170, March 1988. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Matheson:1984:GMY

[MH84] Robert H. Matheson III and Gene R. Huntsman. Growth, mortality, and yield-per-recruit models for speckled hind and snowy grouper from the United States South Atlantic Bight. Transactions of the American Fisheries Society, 113(5):607–616, September 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Murphy:1981:ECM

[MHA81] Michael L. Murphy, Charles P. Hawkins, and N. H. Anderson. Effects of canopy modification and accumulated sediment on stream communities. Transactions of the American Fisheries Society, 110(4):469–478, July 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Matthews:1989:IWQ

[MHEG89] William J. Matthews, Loren G. Hill, David R. Edds, and Frances P. Gelwick. Influence of water quality and season on habitat use by striped bass in a large southwestern reservoir. Transactions of the American Fisheries Society, 118(3):243–250, May 1989. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mehrle:1982:RBB

[MHH^{+}82] Paul M. Mehrle, Terry A. Haines, Steve Hamilton, J. Larry Ludke, Foster L. Mayer, and Michael A. Ribick. Relationship between body contaminants and bone development in East-Coast striped bass. Transactions of the American Fisheries Society, 111(2):231–241, March 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Maceina:1987:AGA

[MHLL87] Michael J. Maceina, David N. Hata, Thomas L. Linton, and Andre M. Landry, Jr. Age and growth analysis of spotted seatrout from Galveston Bay, Texas. Transactions of the American Fisheries Society, 116(1):54–59, January 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Matthews:1985:DDS

[MHS85] William J. Matthews, Loren G. Hill, and Scott M. Schellhaass. Depth distribution of striped bass and other fish in Lake Texoma (Oklahoma–Texas) during summer stratification. Transactions of the American Fisheries Society, 114(1):84–91, January 1985. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McIntyre:1988:NSI

[MHS88] John D. McIntyre, Alan R. Hemmingsen, and Raymond C. Simon. Notes: Selection to increase survival of smolts in four successive broods of Coho salmon. Transactions of the American Fisheries Society, 117(1):90–92, January 1988. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Murawski:1989:YSU

[MI89] S. A. Murawski and J. S. Idoine. Yield sustainability under constant-catch policy and stochastic recruitment. Transactions of the American Fisheries Society, 118(4):349–367, July 1989. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Milstein:1981:ADJ

[Mil81] Charles B. Milstein. Abundance and distribution of juvenile Alosa species off Southern New Jersey. Transactions of the American Fisheries Society, 110(2):306–309, March 1981. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Millikin:1982:EDP

[Mil82] Mark R. Millikin. Effects of dietary protein concentration on growth, feed efficiency, and body composition of age-0 striped bass. Transactions of the American Fisheries Society, 111(3): 373–378, May 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Millikin:1983:IED

[Mil83] Mark R. Millikin. Interactive effects of dietary protein and lipid on growth and protein utilization of age-0 striped bass. Transactions of the American Fisheries Society, 112(2):185–193, March 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Minns:1989:FAF

[Min89] Charles K. Minns. Factors affecting fish species richness in Ontario lakes. Transactions of the American Fisheries Society, 118 (5):533–545, September 1989. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Meyer:1983:TOE

[MJ83] Fred P. Meyer and Ted A. Jorgenson. Teratological and other effects of malachite green on development of rainbow trout and rabbits. Transactions of the American Fisheries Society, 112(6): 818–824, November 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Manny:1989:FTB

[MJE89] Bruce A. Manny, David J. Jude, and Randy L. Eshenroder. Field test of a bioassay procedure for assessing habitat quality on fish spawning grounds. Transactions of the American Fisheries Society, 118(2):175–182, March 1989. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McCormick:1989:SBO

[MJL89] J. Howard McCormick, Kathleen M. Jensen, and Richard L. Leino. Survival, blood osmolality, and gill morphology of juvenile yellow perch, rock bass, black crappie, and largemouth bass exposed to acidified soft water. Transactions of the American Fisheries Society, 118(4):386–399, July 1989. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mansfield:1983:DAL

[MJM^{+}83] Pamela J. Mansfield, David J. Jude, David T. Michaud, Dan C. Brazo, and John Gulvas. Distribution and abundance of larval burbot and deepwater sculpin in Lake Michigan. Transactions of the American Fisheries Society, 112(2):162–172, March 1983. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Meador:1989:BML

[MK89] Michael R. Meador and William E. Kelso. Behavior and movements of largemouth bass in response to salinity. Transactions of the American Fisheries Society, 118(4):409–415, July 1989. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mathias:1982:FHW

[ML82] Jack A. Mathias and Sifa Li. Feeding habits of walleye larvae and juveniles: Comparative laboratory and field studies. Transactions of the American Fisheries Society, 111(6):722–735, November 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Moore:1984:PPR

[ML84] Christopher M. Moore and Ronald F. Labisky. Population parameters of a relatively unexploited stock of snowy grouper in the Lower Florida Keys. Transactions of the American Fisheries Society, 113(3):322–329, May 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Malhotra:1985:FFS

[MM85] Y. R. Malhotra and Sunita Munshi. First feeding and survival of Aspidoparia morar larvae (Cyprinidae). Transactions of the American Fisheries Society, 114(2):286–290, March 1985. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Maceina:1988:VWL

[MM88a] Michael J. Maceina and Brian R. Murphy. Variation in the weightto-length relationship among Florida and northern largemouth bass and their intraspecific F_{1} hybrid. Transactions of the American Fisheries Society, 117(3):232–237, May 1988. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mortensen:1988:OTH

[MM88b] Donald G. Mortensen and Parry D. Mothershead. Occurrence of two hemiuroid trematodes in the stomachs of juvenile Pacific salmon from the marine waters of Southeastern Alaska and British Columbia. Transactions of the American Fisheries Society, 117(5):452–455, September 1988. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mehrle:1981:KME

[MMB81] Paul M. Mehrle, Foster L. Mayer, and Denny R. Buckler. Kepone and mirex: Effects on bone development and swim bladder composition in fathead minnows. Transactions of the American Fisheries Society, 110(5):638–643, September 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Maceina:1988:FRF

[MMI88] Michael J. Maceina, Brian R. Murphy, and J. Jeffery Isely. Factors regulating Florida largemouth bass stocking success and hybridization with northern largemouth bass in Aquilla Lake, Texas. Transactions of the American Fisheries Society, 117(3):221–231, May 1988. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Martin:1981:SSW

[MMNW81] Dan B. Martin, Leslie J. Mengel, Jerry F. Novotny, and Charles H. Walburg. Spring and summer water levels in a Missouri River reservoir: Effects on age-0 fish and zooplankton. Transactions of the American Fisheries Society, 110(3):370–381, May 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Murray:1989:RDB

[MMR89] C. B. Murray, J. D. McPhail, and M. L. Rosenau. Reproductive and developmental biology of Kokanee from Upper Arrow Lake, British Columbia. Transactions of the American Fisheries Society, 118(5):503–509, September 1989. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mayer:1985:WTO

[MMW85] Kathleen S. Mayer, Foster L. Mayer, and Arthur Witt, Jr. Waste transformer oil and PCB toxicity to rainbow trout. Transactions of the American Fisheries Society, 114(6):869–886, November 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mathias:1982:HNQ

[MMY^{+}82] J. A. Mathias, J. Martin, M. Yurkowski, J. G. I. Lark, M. Papst, and J. L. Tabachek. Harvest and nutritional quality of Gammarus lacustris for trout culture. Transactions of the American Fisheries Society, 111(1):83–89, January 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Murphy:1983:UGT

[MNT83] Brian R. Murphy, Larry A. Nielsen, and Bruce J. Turner. Use of genetic tags to evaluate stocking success for reservoir walleyes. Transactions of the American Fisheries Society, 112(4):457–463, July 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McAllister:1986:IPN

[MO86] Philip E. McAllister and Winona J. Owens. Infectious pancreatic necrosis virus: Protocol for a standard challenge to brook trout. Transactions of the American Fisheries Society, 115(3):466–470, May 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McGovern:1988:PPF

[MO88] John C. McGovern and John E. Olney. Potential predation by fish and invertebrates on early life history stages of striped bass in the Pamunkey River, Virginia. Transactions of the American Fisheries Society, 117(2):152–161, March 1988. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Moles:1983:EPM

[Mol83] Adam Moles. Effect of parasitism by mussel glochidia on growth of Coho salmon. Transactions of the American Fisheries Society, 112(2):201–204, March 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Moss:1985:SST

[Mos85] Jerry L. Moss. Summer selection of thermal refuges by striped bass in Alabama reservoirs and tailwaters. Transactions of the American Fisheries Society, 114(1):77–83, January 1985. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Moles:1984:EPT

[MP84] Adam Moles and Jerome J. Pella. Effects of parasitism and temperature on salinity tolerance of the kelp shrimp eualus suckleyi. Transactions of the American Fisheries Society, 113(3):354–359, May 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Matthews:1986:SSC

[MPAR86] Gene M. Matthews, Donn L. Park, Stephen Achord, and Thomas E. Ruehle. Static seawater challenge test to measure relative stress levels in spring Chinook salmon smolts. Transactions of the American Fisheries Society, 115(2):236–244, March 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Magee:1982:BGA

[MPP82] Sheila M. Magee, David P. Philipp, and David P. Philipp. Biochemical genetic analyses of the grass carp [female sign] × bighead carp [male sign] F_{1} hybrid and the parental species. Transactions of the American Fisheries Society, 111(5):593–602, September 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mills:1989:IBP

[MPSC89] Edward L. Mills, Michael V. Pol, Ruth E. Sherman, and Teresa B. Culver. Interrelationships between prey body size and growth of age-0 yellow perch. Transactions of the American Fisheries Society, 118(1):1–10, January 1989. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Morgan:1982:ITS

[MR82] Raymond P. Morgan II and V. James Rasin, Jr. Influence of temperature and salinity on development of white perch eggs. Transactions of the American Fisheries Society, 111(3):396–398, May 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Moles:1983:ECO

[MR83] Adam Moles and Stanley D. Rice. Effects of crude oil and naphthalene on growth, caloric content, and fat content of pink salmon juveniles in seawater. Transactions of the American Fisheries Society, 112(2):205–211, March 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Murray:1989:RJC

[MR89] C. B. Murray and M. L. Rosenau. Rearing of juvenile Chinook salmon in nonnatal tributaries of the Lower Fraser River, British Columbia. Transactions of the American Fisheries Society, 118(3):284–289, May 1989. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Morgan:1981:TSE

[MRC81] Raymond P. Morgan II, V. James Rasin, Jr., and Ronald L. Copp. Temperature and salinity effects on development of striped bass eggs and larvae. Transactions of the American Fisheries Society, 110(1):95–99, January 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Muth:1987:SUF

[MRG87] Robert M. Muth, David E. Ruppert, and Ronald J. Glass. Subsistence use of fisheries resources in Alaska: Implications for great lakes fisheries management. Transactions of the American Fisheries Society, 116(3):510–518, May 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Morgan:1983:SEE

[MRN83] Raymond P. Morgan II, V. James Rasin, Jr., and Linda A. Noe. Sediment effects on eggs and larvae of striped bass and white perch. Transactions of the American Fisheries Society, 112(2): 220–224, March 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Michaletz:1982:FEG

[MRTR82] Paul H. Michaletz, Charles F. Rabeni, William W. Taylor, and Thomas R. Russell. Feeding ecology and growth of young-of-the-year paddlefish in hatchery ponds. Transactions of the American Fisheries Society, 111(6):700–709, November 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Miller:1982:DGR

[MS82] Steven J. Miller and Ted Storck. Daily growth rings in otoliths of young-of-the-year largemouth bass. Transactions of the American Fisheries Society, 111(4):527–530, July 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Miller:1984:TSD

[MS84a] Steven J. Miller and Ted Storck. Temporal spawning distribution of largemouth bass and young-of-year growth, determined from daily otolith rings. Transactions of the American Fisheries Society, 113(5):571–578, September 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Muth:1984:DTF

[MS84b] Robert T. Muth and James C. Schmulbach. Downstream transport of fish larvae in a shallow prairie river. Transactions of the American Fisheries Society, 113(2):224–230, March 1984. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Marty:1986:PME

[MS86] Gary D. Marty and Robert C. Summerfelt. Pathways and mechanisms for expulsion of surgically implanted dummy transmitters from channel catfish. Transactions of the American Fisheries Society, 115(4):577–589, July 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Marty:1988:IRC

[MS88a] Gary D. Marty and Robert C. Summerfelt. Inflammatory response of channel catfish to abdominal implants: a histological and ultrastructural study. Transactions of the American Fisheries Society, 117(4):401–416, July 1988. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Moffitt:1988:ADO

[MS88b] Christine M. Moffitt and Julie A. Schreck. Accumulation and depletion of orally administered erythromycin thiocyanate in tissues of Chinook salmon. Transactions of the American Fisheries Society, 117(4):394–400, July 1988. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mesa:1989:EMR

[MS89] Matthew G. Mesa and Carl B. Schreck. Electrofishing mark– recapture and depletion methodologies evoke behavioral and physiological changes in cutthroat trout. Transactions of the American Fisheries Society, 118(6):644–658, November 1989. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Margenau:1988:NLW

[MSB88] Terry L. Margenau, Stephen T. Schram, and William H. Blust. Notes: Lymphocystis in a walleye population. Transactions of the American Fisheries Society, 117(3):308–310, May 1988. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Maule:1988:PEC

[MSBB88] Alec G. Maule, Carl B. Schreck, C. Samuel Bradford, and Bruce A. Barton. Physiological effects of collecting and transporting emigrating juvenile Chinook salmon past dams on the Columbia River. Transactions of the American Fisheries Society, 117(3):245–261, May 1988. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mather:1986:EAM

[MSC86] Martha E. Mather, Roy A. Stein, and Robert F. Carline. Experimental assessment of mortality and hyperglycemia in tiger muskellunge due to stocking stressors. Transactions of the American Fisheries Society, 115(5):762–770, September 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mathur:1981:SAT

[MSPS81] Dilip Mathur, Robert M. Schutsky, Edmund J. Purdy, Jr., and Carl A. Silver. Similarities in acute temperature preferences of freshwater fishes. Transactions of the American Fisheries Society, 110(1):1–13, January 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mattice:1981:CTH

[MTB81] J. S. Mattice, S. C. Tsai, and M. B. Burch. Comparative toxicity of hypochlorous acid and hypochlorite ions to mosquitofish. Transactions of the American Fisheries Society, 110(4):519–525, July 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mattice:1981:TTC

[MTBB81] J. S. Mattice, S. C. Tsai, M. B. Burch, and J. J. Beauchamp. Toxicity of trihalomethanes to common carp embryos. Transactions of the American Fisheries Society, 110(2):261–269, March 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McCormick:1981:RLB

[MW81] J. Howard McCormick and Julie A. Wegner. Responses of largemouth bass from different latitudes to elevated water temperatures. Transactions of the American Fisheries Society, 110(3): 417–429, May 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mesing:1986:HRS

[MW86] Charles L. Mesing and Anton M. Wicker. Home range, spawning migrations, and homing of radio-tagged Florida largemouth bass in two central Florida lakes. Transactions of the American Fisheries Society, 115(2):286–295, March 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mundahl:1988:SDE

[MW88] Neal D. Mundahl and Thomas E. Wissing. Selection and digestive efficiencies of gizzard shad feeding on natural detritus and two laboratory diets. Transactions of the American Fisheries Society, 117(5):480–487, September 1988. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Matlock:1982:SNE

[MWG82] Gary C. Matlock, James E. Weaver, and Albert W. Green. Sampling nearshore estuarine fishes with rotenone. Transactions of the American Fisheries Society, 111(3):326–331, May 1982. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Martin:1985:IFS

[MWMSH85] F. Douglas Martin, David A. Wright, Jay C. Means, and Eileen M. Setzler-Hamilton. Importance of food supply to nutritional state of larval striped bass in the Potomac River Estuary. Transactions of the American Fisheries Society, 114(1):137–145, January 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Moser:1983:PHS

[MWSR83] Mike Moser, Jeannette Whipple, Judy Sakanari, and Carol Reilly. Protandrous hermaphroditism in striped bass from Coos Bay, Oregon. Transactions of the American Fisheries Society, 112(4): 567–569, July 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Moring:1981:HIS

[MYA81] John R. Moring, Ronald L. Youker, and Kathleen J. Anderson. High incidence of scale regeneration by potamodromous coastal cutthroat trout: Analytical implications. Transactions of the American Fisheries Society, 110(5):621–626, September 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Minello:1989:MYB

[MZM89] Thomas J. Minello, Roger J. Zimmerman, and Eduardo X. Martinez. Mortality of young brown shrimp Penaeus aztecus in estuarine nurseries. Transactions of the American Fisheries Society, 118(6):693–708, November 1989. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Nakashima:1987:RTV

[Nak87] Brian S. Nakashima. Regional and temporal variations in fecundity of capelin in Newfoundland waters. Transactions of the American Fisheries Society, 116(6):864–873, November 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Neitzel:1985:TEE

[NB85] Duane A. Neitzel and C. Dale Becker. Tolerance of eggs, embryos, and alevins of Chinook salmon to temperature changes and reduced humidity in dewatered redds. Transactions of the American Fisheries Society, 114(2):267–273, March 1985. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Newcombe:1981:PEC

[New81] C. Newcombe. A procedure to estimate changes in fish populations caused by changes in stream discharge. Transactions of the American Fisheries Society, 110(3):382–390, May 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ney:1981:FFM

[Ney81] John J. Ney. Forage-fish management in the United States: Evolution of forage-fish management in lakes and reservoirs. Transactions of the American Fisheries Society, 110(6):725–728, November 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Neilson:1981:ESS

[NG81] John D. Neilson and Glen H. Geen. Enumeration of spawning salmon from spawner residence time and aerial counts. Transactions of the American Fisheries Society, 110(4):554–556, July 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Neilson:1986:FYG

[NG86] John D. Neilson and Glen H. Geen. First-year growth rate of sixes river Chinook salmon as inferred from otoliths: Effects on mortality and age at maturity. Transactions of the American Fisheries Society, 115(1):28–33, January 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Nielsen:1983:VCY

[Nie83] Larry A. Nielsen. Variation in the catchability of yellow perch in an otter trawl. Transactions of the American Fisheries Society, 112(1):53–59, January 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Naesje:1986:DCW

[NJS86] Tor F. Næsje, Bror Jonsson, and Odd T. Sandlund. Drift of cisco and whitefish larvae in a Norwegian River. Transactions of the American Fisheries Society, 115(1):89–93, January 1986. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Nelson:1982:GMR

[NM82] Russell S. Nelson and Charles S. Manooch III. Growth and mortality of red snappers in the West-central Atlantic Ocean and Northern Gulf of Mexico. Transactions of the American Fisheries Society, 111(4):465–475, July 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Nemetz:1988:WHI

[NM88] T. G. Nemetz and J. R. Macmillan. Wound healing of incisions closed with a cyanoacrylate adhesive. Transactions of the American Fisheries Society, 117(2):190–195, March 1988. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Nammack:1985:LHS

[NMC85] Marta F. Nammack, J. A. Musick, and J. A. Colvocoresses. Life history of spiny dogfish off the Northeastern United States. Transactions of the American Fisheries Society, 114(3):367–376, May 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Nigro:1982:REL

[NN82] Anthony A. Nigro and John J. Ney. Reproduction and earlylife accommodations of landlocked alewives to a southern range extension. Transactions of the American Fisheries Society, 111 (5):559–569, September 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Noakes:1988:CSA

[Noa88] David L. G. Noakes. Common strategies of anadromous and catadromous fishes. American fisheries society symposium 1. Transactions of the American Fisheries Society, 117(5):515–518, September 1988. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Noble:1981:MFF

[Nob81] Richard L. Noble. Management of forage fishes in impoundments of the Southern United States. Transactions of the American Fisheries Society, 110(6):738–750, November 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Neves:1983:APF

[NP83] Richard J. Neves and Garland B. Pardue. Abundance and production of fishes in a small Appalachian stream. Transactions of the American Fisheries Society, 112(1):21–26, January 1983. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Noeske:1984:CFT

[NS84a] Teresa A. Noeske and Richard E. Spieler. Circadian feeding time affects growth of fish. Transactions of the American Fisheries Society, 113(4):540–544, July 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Norcross:1984:OET

[NS84b] Brenda L. Norcross and Richard F. Shaw. Oceanic and estuarine transport of fish eggs and larvae: a review. Transactions of the American Fisheries Society, 113(2):153–165, March 1984. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Oglesby:1982:MSO

[Ogl82] Ray T. Oglesby. The MEI symposium — overview and observations. Transactions of the American Fisheries Society, 111(2): 171–175, March 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Olney:1983:PIP

[OGS^{+}83] John E. Olney, George C. Grant, Fred E. Schultz, C. Lawrence Cooper, and John Hageman. Pterygiophore-interdigitation patterns in larvae of four Morone species. Transactions of the American Fisheries Society, 112(4):525–531, July 1983. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Osborne:1981:ECS

[OIWD81] L. L. Osborne, D. R. Iredale, F. J. Wrona, and R. W. Davies. Effects of chlorinated sewage effluents on fish in the Sheep River, Alberta. Transactions of the American Fisheries Society, 110(4): 536–540, July 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

OLeary:1986:BLS

[OK86] John A. O’Leary and Boyd Kynard. Behavior, length, and sex ratio of seaward-migrating juvenile American shad and blueback herring in the Connecticut River. Transactions of the American Fisheries Society, 115(4):529–536, July 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

OIestad:1985:MPA

[ØKF85] V. ØIestad, P. G. Kvenseth, and A. Folkvord. Mass production of Atlantic cod juveniles Gadus morhua in a Norwegian saltwater pond. Transactions of the American Fisheries Society, 114(4): 590–595, July 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Orth:1982:EIM

[OM82] Donald J. Orth and O. Eugene Maughan. Evaluation of the incremental methodology for recommending instream flows for fishes. Transactions of the American Fisheries Society, 111(4):413–445, July 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Orsi:1987:SVL

[Ors87] Joseph A. Orsi. Small versus large trolling lures for sampling juvenile Chinook salmon and Coho salmon. Transactions of the American Fisheries Society, 116(1):50–53, January 1987. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

OGorman:1986:DAL

[OS86] Robert O’Gorman and Clifford P. Schneider. Dynamics of alewives in Lake Ontario following a mass mortality. Transactions of the American Fisheries Society, 115(1):1–14, January 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Overholtz:1986:ESM

[OT86] William J. Overholtz and Albert V. Tyler. An exploratory simulation model of competition and predation in a demersal fish assemblage on Georges Bank. Transactions of the American Fisheries Society, 115(6):805–817, November 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Olson:1988:RPS

[OWNH88] Robert A. Olson, Jimmy D. Winter, David C. Nettles, and James M. Haynes. Resource partitioning in summer by salmonids in South-Central Lake Ontario. Transactions of the American Fisheries Society, 117(6):552–559, November 1988. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pine:1989:ESV

[PAH89] Robert T. Pine, Lars W. J. Anderson, and Silas S. O. Hung. Effects of static versus flowing water on aquatic plant preferences of triploid grass carp. Transactions of the American Fisheries Society, 118(3):336–344, May 1989. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Parker:1984:CAA

[Par84] Nick C. Parker. Chronobiologic approach to aquaculture. Transactions of the American Fisheries Society, 113(4):545–552, July 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pauly:1982:MES

[Pau82] Daniel Pauly. A method to estimate the stock-recruitment relationship of shrimps. Transactions of the American Fisheries Society, 111(1):13–20, January 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pimentel:1983:CTD

[PB83] R. Pimentel and R. V. Bulkley. Concentrations of total dissolved solids preferred or avoided by endangered Colorado River fishes. Transactions of the American Fisheries Society, 112(5):595–600, September 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Powell:1985:MIN

[PC85] Allyn B. Powell and Alexander J. Chester. Morphometric indices of nutritional condition and sensitivity to starvation of spot larvae. Transactions of the American Fisheries Society, 114(3): 338–347, May 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pierce:1985:IRS

[PCC85] Rodney B. Pierce, Daniel W. Coble, and Scott D. Corley. Influence of river stage on shoreline electrofishing catches in the Upper Mississippi River. Transactions of the American Fisheries Society, 114(6):857–860, November 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Philipp:1983:BGE

[PCW83] David P. Philipp, William F. Childers, and Gregory S. Whitt. A biochemical genetic evaluation of the northern and Florida subspecies of largemouth bass. Transactions of the American Fisheries Society, 112(1):1–20, January 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pikitch:1988:ASM

[PD88] Ellen K. Pikitch and Robert L. Demory. Assessment of scales as a means of aging Dover sole. Transactions of the American Fisheries Society, 117(4):345–349, July 1988. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Peck:1982:ERY

[Pec82] James W. Peck. Extended residence of young-of-the-year lake trout in shallow water. Transactions of the American Fisheries Society, 111(6):775–778, November 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pascho:1987:CFT

[PEMM87] R. J. Pascho, D. G. Elliott, R. W. Mallett, and D. Mulcahy. Comparison of five techniques for the detection of Renibacterium salmoninarum in adult Coho salmon. Transactions of the American Fisheries Society, 116(6):882–890, November 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Price:1985:NEC

[PFT^{+}85] Kent S. Price, David A. Flemer, Jay L. Taft, Gail B. Mackiernan, Willa Nehlsen, Robert B. Biggs, Ned H. Burger, and Dewey A. Blaylock. Nutrient enrichment of Chesapeake Bay and its impact on the habitat of striped bass: a speculative hypothesis. Transactions of the American Fisheries Society, 114(1):97–106, January 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Propst:1987:REI

[PG87] Dennis B. Propst and Dimitris G. Gavrilis. Role of economic impact assessment procedures in recreational fisheries management. Transactions of the American Fisheries Society, 116(3):450–460, May 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Prager:1989:SEA

[PH89] Michael H. Prager and John M. Hoenig. Superposed epoch analysis: a randomization test of environmental effects on recruitment with application to chub mackerel. Transactions of the American Fisheries Society, 118(6):608–618, November 1989. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Palawski:1985:SYS

[PHD85] Donald Palawski, Joseph B. Hunn, and F. James Dwyer. Sensitivity of young striped bass to organic and inorganic contaminants in fresh and saline waters. Transactions of the American Fisheries Society, 114(5):748–753, September 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pinto:1982:HBS

[Pin82] Lope Garcia Pinto. Hybridization between species of tilapia. Transactions of the American Fisheries Society, 111(4):481–484, July 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Parsons:1986:AGN

[PLC86] D. G. Parsons, G. R. Lilly, and G. J. Chaput. Age and growth of northern shrimp Pandalus borealis off Northeastern Newfoundland and Southern Labrador. Transactions of the American Fisheries Society, 115(6):872–881, November 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Prince:1981:ASS

[PM81] Eric D. Prince and Leslie J. Mengel. Aggregation of spottail shiners in the heated discharge of a nuclear power station. Transactions of the American Fisheries Society, 110(2):221–225, March 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Palm:1983:RBE

[PM83] Roger C. Palm, Jr. and Stephen P. Malvestuto. Relationships between economic benefit and sport-fishing effort on West point reservoir, Alabama-Georgia. Transactions of the American Fisheries Society, 112(1):71–78, January 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Paloheimo:1982:EFM

[PMB82] J. E. Paloheimo, Lyman L. McDonald, and Susan Ann Butler. Estimation of fish mortality. Transactions of the American Fisheries Society, 111(4):535–537, July 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Peterson:1982:WUA

[PMR82] R. H. Peterson and D. J. Martin-Robichaud. Water uptake by Atlantic salmon ova as affected by low pH. Transactions of the American Fisheries Society, 111(6):772–774, November 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Peterson:1986:GMI

[PMR86] R. H. Peterson and D. J. Martin-Robichaud. Growth and major inorganic cation budgets of Atlantic salmon alevins at three ambient acidities. Transactions of the American Fisheries Society, 115(2):220–226, March 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pajak:1987:HSF

[PN87] Paul Pajak and Richard J. Neves. Habitat suitability and fish production: a model evaluation for rock bass in two Virginia streams. Transactions of the American Fisheries Society, 116(6):839–850, November 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Powell:1982:AFO

[Pow82] Allyn B. Powell. Annulus formation on otoliths and growth of young summer flounder from Pamlico Sound, North Carolina. Transactions of the American Fisheries Society, 111(6):688–693, November 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Prager:1988:GMD

[Pra88] Michael H. Prager. Group method of data handling: a new method for stock identification. Transactions of the American Fisheries Society, 117(3):290–296, May 1988. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Probst:1984:RUS

[PRCM84] W. E. Probst, C. F. Rabeni, W. G. Covington, and R. E. Marteney. Resource use by stream-dwelling rock bass and smallmouth bass. Transactions of the American Fisheries Society, 113(3):283–294, May 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Prentice:1989:LTT

[Pre89] John A. Prentice. Low-temperature tolerance of southern flounder in Texas. Transactions of the American Fisheries Society, 118(1): 30–35, January 1989. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Peterman:1981:RBS

[PS81a] Randall M. Peterman and Gregory J. Steer. Relation between sport-fishing catchability coefficients and salmon abundance. Transactions of the American Fisheries Society, 110(5): 585–593, September 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Peters:1981:FRS

[PS81b] David S. Peters and William E. Schaaf. Food requirements and sources for juvenile Atlantic menhaden. Transactions of the American Fisheries Society, 110(3):317–324, May 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Patino:1986:ERC

[PSBZ86] Reynaldo Patiño, Carl B. Schreck, Joe L. Banks, and Wally S. Zaugg. Effects of rearing conditions on the developmental physiology of smolting Coho salmon. Transactions of the American Fisheries Society, 115(6):828–837, November 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pauley:1988:CTM

[PT88] Gilbert B. Pauley and David A. Troutt. Comparison of three methods of fluorescent dye application for marking juvenile steelhead. Transactions of the American Fisheries Society, 117(3): 311–313, May 1988. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Platts:1989:CSS

[PTMG89] William S. Platts, Richard J. Torquemada, Michael L. McHenry, and Charles K. Graham. Changes in salmon spawning and rearing habitat from increased delivery of fine sediment to the South Fork Salmon River, Idaho. Transactions of the American Fisheries Society, 118(3):274–283, May 1989. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Petering:1988:REP

[PV88] Raymond W. Petering and Michael J. Van Den Avyle. Relative efficiency of a pump for sampling larval gizzard and threadfin shad. Transactions of the American Fisheries Society, 117(1): 78–83, January 1988. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pierce:1981:AFE

[PWM81a] Robert J. Pierce, Thomas E. Wissing, and Bernard A. Megrey. Aspects of the feeding ecology of gizzard shad in acton lake, Ohio. Transactions of the American Fisheries Society, 110(3):391–395, May 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pierce:1981:RMG

[PWM81b] Robert J. Pierce, Thomas E. Wissing, and Bernard A. Megrey. Respiratory metabolism of gizzard shad. Transactions of the American Fisheries Society, 110(1):51–55, January 1981. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Phillips:1989:PDC

[PZI89] Ruth B. Phillips, Kerry D. Zajicek, and Peter E. Ihssen. Population differences in chromosome-banding polymorphisms in lake trout. Transactions of the American Fisheries Society, 118(1): 64–73, January 1989. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Quiros:1985:FAR

[QB85] Rolando Quirós and Claudio Baigún. Fish abundance related to organic matter in the Plata River Basin, South America. Transactions of the American Fisheries Society, 114(3):377–387, May 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Quantz:1985:UEE

[Qua85] Gerrit Quantz. Use of endogenous energy sources by larval turbot Scophthalmus maximus. Transactions of the American Fisheries Society, 114(4):558–563, July 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ross:1986:DDG

[RA86] Michael R. Ross and Frank P. Almeida. Density-dependent growth of silver hakes. Transactions of the American Fisheries Society, 115(4):548–554, July 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Radtke:1984:FSC

[Rad84] Richard L. Radtke. Formation and structural composition of larval striped mullet otoliths. Transactions of the American Fisheries Society, 113(2):186–191, March 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rahel:1981:SZT

[Rah81] Frank J. Rahel. Selection for zinc tolerance in fish: Results from laboratory and wild populations. Transactions of the American Fisheries Society, 110(1):19–28, January 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rankin:1986:HSS

[Ran86] Edward T. Rankin. Habitat selection by smallmouth bass in response to physical characteristics in a natural stream. Transactions of the American Fisheries Society, 115(2):322–334, March 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Randall:1987:TEV

[Ran87] Alan Randall. Total economic value as a basis for policy. Transactions of the American Fisheries Society, 116(3):325–335, May 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ratliff:1981:CSE

[Rat81] D. E. Ratliff. Ceratomyxa shasta: Epizootiology in Chinook salmon of central Oregon. Transactions of the American Fisheries Society, 110(4):507–513, July 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Reash:1986:FTM

[RB86] Rob J. Reash and Tim M. Berra. Fecundity and trace-metal content of creek chubs from a metal-contaminated stream. Transactions of the American Fisheries Society, 115(2):346–351, March 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Riley:1982:ESS

[RC82] L. M. Riley and R. F. Carline. Evaluation of scale shape for the identification of walleye stocks from Western Lake Erie. Transactions of the American Fisheries Society, 111(6):736–741, November 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Richards:1987:UAR

[RC87] R. Anne Richards and J. Stanley Cobb. Use of avoidance responses to keep spider crabs out of traps for American lobsters. Transactions of the American Fisheries Society, 116(2):282–285, March 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rice:1985:EOA

[RCB85] James A. Rice, Larry B. Crowder, and Fred P. Binkowski. Evaluating otolith analysis for bloater Coregonus hoyi: Do otoliths ring true? Transactions of the American Fisheries Society, 114(4): 532–539, July 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rice:1987:EMR

[RCH87] James A. Rice, Larry B. Crowder, and Mark E. Holey. Exploration of mechanisms regulating larval survival in Lake Michigan bloater: a recruitment analysis based on characteristics of individual larvae. Transactions of the American Fisheries Society, 116 (5):703–718, September 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Reckahn:1986:LTC

[Rec86] J. A. Reckahn. Long-term cyclical trends in growth of lake whitefish in South Bay, Lake Huron. Transactions of the American Fisheries Society, 115(6):787–804, November 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Reinitz:1983:IDF

[Rei83] Gary Reinitz. Influence of diet and feeding rate on the performance and production cost of rainbow trout. Transactions of the American Fisheries Society, 112(6):830–833, November 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Restrepo:1989:GEM

[Res89] Victor R. Restrepo. Growth estimates for male stone crabs along the Southwest Coast of Florida: a synthesis of available data and methods. Transactions of the American Fisheries Society, 118(1): 20–29, January 1989. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rettig:1987:BMD

[Ret87] R. Bruce Rettig. Bioeconomic models: Do they really help fishery managers? Transactions of the American Fisheries Society, 116(3):405–411, May 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rieman:1981:EEM

[RF81] B. E. Rieman and C. M. Falter. Effects of the establishment of Mysis relicta on the macrozooplankton of a large lake. Transactions of the American Fisheries Society, 110(5):613–620, September 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Restrepo:1988:PUS

[RF88] Victor R. Restrepo and William W. Fox, Jr. Parameter uncertainty and simple yield-per-recruit analysis. Transactions of the American Fisheries Society, 117(3):282–289, May 1988. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rach:1986:DAR

[RG86] J. J. Rach and W. H. Gingerich. Distribution and accumulation of rotenone in tissues of warmwater fishes. Transactions of the American Fisheries Society, 115(2):214–219, March 1986. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ross:1982:AGM

[RH82] Jeffrey L. Ross and Gene R. Huntsman. Age, growth, and mortality of blueline tilefish from North Carolina and South Carolina. Transactions of the American Fisheries Society, 111(5): 585–592, September 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rosen:1982:BEP

[RHU82] Rudolph A. Rosen, Donald C. Hales, and Dennis G. Unkenholz. Biology and exploitation of paddlefish in the Missouri River below gavins point dam. Transactions of the American Fisheries Society, 111(2):216–222, March 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rigler:1982:RBF

[Rig82] F. H. Rigler. The relation between fisheries management and limnology. Transactions of the American Fisheries Society, 111 (2):121–132, March 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rimmer:1985:ERD

[Rim85] David M. Rimmer. Effects of reduced discharge on production and distribution of age-0 rainbow trout in seminatural channels. Transactions of the American Fisheries Society, 114(3):388–396, May 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rombough:1985:RWU

[RJ85] P. J. Rombough and J. O. T. Jensen. Reduced water uptake and resistance to deformation in acid-exposed eggs of steelhead Salmo gairdneri. Transactions of the American Fisheries Society, 114(4):571–576, July 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rehnberg:1985:OSD

[RJS85] Bradley G. Rehnberg, Brian Jonasson, and Carl B. Schreck. Olfactory sensitivity during parr and smolt developmental stages of Coho salmon. Transactions of the American Fisheries Society, 114(5):732–736, September 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rosenau:1987:IDA

[RM87a] Marvin L. Rosenau and J. D. McPhail. Inherited differences in agonistic behavior between two populations of Coho salmon. Transactions of the American Fisheries Society, 116(4):646–654, July 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rulifson:1987:FSB

[RM87b] Roger A. Rulifson and Sean A. McKenna. Food of striped bass in the Upper Bay of Fundy, Canada. Transactions of the American Fisheries Society, 116(1):119–122, January 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ready:1985:NEF

[RMC85] Richard C. Ready, Edward L. Mills, and John L. Confer. A new estimator of, and factors influencing, the sampling variance of the linear index of food selection. Transactions of the American Fisheries Society, 114(2):258–266, March 1985. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rogier:1985:EAG

[RNT85] Catherine G. Rogier, John J. Ney, and Bruce J. Turner. Electrophoretic analysis of genetic variability in a landlocked striped bass population. Transactions of the American Fisheries Society, 114(2):244–249, March 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ross:1988:AGM

[Ros88] Steve W. Ross. Age, growth, and mortality of Atlantic croaker in North Carolina, with comments on population dynamics. Transactions of the American Fisheries Society, 117(5):461–473, September 1988. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ross:1989:DBD

[Ros89] Steve W. Ross. Diet of the banded drum in North Carolina. Transactions of the American Fisheries Society, 118(6):680–686, November 1989. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Riley:1989:MMV

[RPI89] S. C. Riley, G. Power, and P. E. Ihssen. Meristic and morphometric variation in parr of ouananiche and anadromous Atlantic salmon from rivers along the North Shore of the Gulf of St. Lawrence. Transactions of the American Fisheries Society, 118 (5):515–522, September 1989. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Redding:1983:IAS

[RS83] J. Michael Redding and Carl B. Schreck. Influence of ambient salinity on osmoregulation and cortisol concentration in yearling Coho salmon during stress. Transactions of the American Fisheries Society, 112(6):800–807, November 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Redding:1987:PEC

[RSE87] J. Michael Redding, Carl B. Schreck, and Fred H. Everest. Physiological effects on Coho salmon and steelhead of exposure to suspended solids. Transactions of the American Fisheries Society, 116(5):737–744, September 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Roberts:1985:CVR

[RTP85] Kenneth J. Roberts, Mark E. Thompson, and Perry W. Pawlyk. Contingent valuation of recreational diving at petroleum rigs, Gulf of Mexico. Transactions of the American Fisheries Society, 114(2):214–219, March 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rogers:1984:FNU

[RTV84] S. Gordon Rogers, Timothy E. Targett, and Scott B. Van Sant. Fish-nursery use in Georgia salt-Marsh estuaries: The influence of springtime freshwater conditions. Transactions of the American Fisheries Society, 113(5):595–606, September 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ruggerone:1986:CMJ

[Rug86] Gregory T. Ruggerone. Consumption of migrating juvenile salmonids by gulls foraging below a Columbia River dam. Transactions of the American Fisheries Society, 115(5):736–742, September 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rijnsdorp:1985:STT

[RVV85] A. D. Rijnsdorp, M. Van Stralen, and H. W. Van Der Veer. Selective tidal transport of North Sea plaice larvae Pleuronectes platessa in coastal nursery areas. Transactions of the American Fisheries Society, 114(4):461–470, July 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rogers:1981:LSE

[RW81a] Bruce A. Rogers and Deborah T. Westin. Laboratory studies on effects of temperature and delayed initial feeding on development of striped bass larvae. Transactions of the American Fisheries Society, 110(1):100–110, January 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ross:1981:WMF

[RW81b] Marvin Jon Ross and Jimmy D. Winter. Winter movements of four fish species near a thermal plume in Northern Minnesota. Transactions of the American Fisheries Society, 110(1):14–18, January 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Reiser:1983:ECR

[RW83] Dudley W. Reiser and Robert G. White. Effects of complete redd dewatering on salmonid egg-hatching success and development of juveniles. Transactions of the American Fisheries Society, 112(4): 532–540, July 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rago:1986:DPA

[RW86] Paul J. Rago and James G. Wiener. Does pH affect fish species richness when lake area is considered? Transactions of the American Fisheries Society, 115(3):438–447, May 1986. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ryder:1982:MIU

[Ryd82] R. A. Ryder. The morphoedaphic index — use, abuse, and fundamental concepts. Transactions of the American Fisheries Society, 111(2):154–164, March 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smagula:1982:DDV

[SA82] Carl M. Smagula and Ira R. Adelman. Day-to-day variation in food consumption by largemouth bass. Transactions of the American Fisheries Society, 111(5):543–548, September 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Shelton:1983:CFC

[SB83] James L. Shelton, Jr. and Claude E. Boyd. Correction factors for calculating oxygen-transfer rates of pond aerators. Transactions of the American Fisheries Society, 112(1):120–122, January 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Scarnecchia:1986:PHT

[SB86a] Dennis L. Scarnecchia and Eric P. Bergersen. Production and habitat of threatened greenback and Colorado River cutthroat trouts in rocky mountain headwater streams. Transactions of the American Fisheries Society, 115(3):382–391, May 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sorensen:1986:ECF

[SB86b] Peter W. Sorensen and Marco L. Bianchini. Environmental correlates of the freshwater migration of elvers of the American eel in a rhode Island brook. Transactions of the American Fisheries Society, 115(2):258–268, March 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stewart:1986:DCF

[SB86c] Donald J. Stewart and Fred P. Binkowski. Dynamics of consumption and food conversion by Lake Michigan alewives: an energetics-modeling synthesis. Transactions of the American Fisheries Society, 115(5):643–661, September 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sigler:1984:ECT

[SBE84] John W. Sigler, T. C. Bjornn, and Fred H. Everest. Effects of chronic turbidity on density and growth of steelheads and Coho salmon. Transactions of the American Fisheries Society, 113(2): 142–150, March 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sullivan:1983:PTH

[SBJD83] Craig V. Sullivan, Scott D. Brewer, Gregory P. Johnston, and Walton W. Dickhoff. Plasma thyroid-hormone concentrations and gill (Na+K)-ATPase activities in postemergent pink salmon. Transactions of the American Fisheries Society, 112(6):825–829, November 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Shireman:1981:ERS

[SCD81] Jerome V. Shireman, Douglas E. Colle, and Douglas F. Durant. Efficiency of rotenone sampling with large and small block nets in vegetated and open-water habitats. Transactions of the American Fisheries Society, 110(1):77–80, January 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stein:1981:LBP

[SCH81] Roy A. Stein, Robert F. Carline, and Robert S. Hayward. Largemouth bass predation on stocked tiger muskellunge. Transactions of the American Fisheries Society, 110(5):604–612, September 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Schramm:1989:FAO

[Sch89] Harold L. Schramm, Jr. Formation of annuli in otoliths of bluegills. Transactions of the American Fisheries Society, 118(5): 546–555, September 1989. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Scott:1985:DGF

[Sco85] Alasdair Scott. Distribution, growth, and feeding of postemergent grayling Thymallus thymallus in an English River. Transactions of the American Fisheries Society, 114(4):525–531, July 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Scoppettone:1988:GLC

[Sco88] G. Gary Scoppettone. Growth and longevity of the cui-ui and longevity of other catostomids and cyprinids in Western North America. Transactions of the American Fisheries Society, 117(3): 301–307, May 1988. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sechnick:1986:HSS

[SCSR86] Carolyn W. Sechnick, Robert F. Carline, Roy A. Stein, and Edward T. Rankin. Habitat selection by smallmouth bass in response to physical characteristics of a simulated stream. Transactions of the American Fisheries Society, 115(2):314–321, March 1986. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Shirvell:1983:MCB

[SD83] C. S. Shirvell and R. G. Dungey. Microhabitats chosen by brown trout for feeding and spawning in rivers. Transactions of the American Fisheries Society, 112(3):355–367, May 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stynes:1987:STC

[SD87] Daniel J. Stynes and Dennis M. Donnelly. Simplifying the travel cost method. Transactions of the American Fisheries Society, 116(3):432–440, May 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Saila:1987:EAM

[SE87] Saul B. Saila and Karim Erzini. Empirical approach to multispecies stock assessment. Transactions of the American Fisheries Society, 116(4):601–611, July 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Serns:1982:IVF

[Ser82a] Steven L. Serns. Influence of various factors on density and growth of age-0 walleyes in Escanaba Lake, Wisconsin, 1958–1980. Transactions of the American Fisheries Society, 111(3):299–306, May 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Serns:1982:RTP

[Ser82b] Steven L. Serns. Relation of temperature and population density to first-year recruitment and growth of smallmouth bass in a Wisconsin lake. Transactions of the American Fisheries Society, 111 (5):570–574, September 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Serns:1984:FSS

[Ser84] Steven L. Serns. First-summer survival, eggs to juveniles, of smallmouth bass in Nebish Lake, Wisconsin. Transactions of the American Fisheries Society, 113(3):304–307, May 1984. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Serns:1986:CAI

[Ser86] Steven L. Serns. Cohort analysis as an indication of walleye yearclass strength in Escanaba Lake, Wisconsin, 1956–1974. Transactions of the American Fisheries Society, 115(6):849–852, November 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Soderberg:1983:EAG

[SFS83] Richard W. Soderberg, John B. Flynn, and H. R. Schmittou. Effects of ammonia on growth and survival of rainbow trout in intensive static-water culture. Transactions of the American Fisheries Society, 112(3):448–451, May 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Savitz:1983:EFH

[SFW83] J. Savitz, Pamela A. Fish, and R. Weszely. Effects of forage on home-range size of largemouth bass. Transactions of the American Fisheries Society, 112(6):772–776, November 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Simco:1989:ISG

[SGK^{+}89] Bill A. Simco, Cheryl A. Goudie, Gerald T. Klar, Nick C. Parker, and Kenneth B. Davis. Influence of sex on growth of channel catfish. Transactions of the American Fisheries Society, 118(4): 427–434, July 1989. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Southall:1984:HUA

[SH84] Peter D. Southall and Wayne A. Hubert. Habitat use by adult paddlefish in the Upper Mississippi River. Transactions of the American Fisheries Society, 113(2):125–131, March 1984. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Simonson:1986:EAE

[SH86] Jenny L. Simonson and Randall J. Hochberg. Effects of air exposure and claw breaks on survival of stone crabs Menippe mercenaria. Transactions of the American Fisheries Society, 115(3): 471–477, May 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Setzler-Hamilton:1981:STD

[SHBM^{+}81] Eileen M. Setzler-Hamilton, Walter R. Boynton, Joseph A. Mihursky, Tibor T. Polgar, and Kathyrn V. Wood. Spatial and temporal distribution of striped bass eggs, larvae, and juveniles in the Potomac estuary. Transactions of the American Fisheries Society, 110(1):121–136, January 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Storr:1983:DRB

[SHCMS83] John F. Storr, Patricia J. Hadden-Carter, Julian M. Myers, and A. Garry Smythe. Dispersion of rock bass along the south shore of Lake Ontario. Transactions of the American Fisheries Society, 112(5):618–628, September 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Saxton:1984:SNP

[SHI84] A. M. Saxton, W. K. Hershberger, and R. N. Iwamoto. Smoltification in the net-pen culture of Coho salmon: Quantitative genetic analysis. Transactions of the American Fisheries Society, 113(3): 339–347, May 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Starnes:1983:LFT

[SHM83] Lynn B. Starnes, Peter A. Hackney, and Thomas A. McDonough. Larval fish transport: a case study of white bass. Transactions of the American Fisheries Society, 112(3):390–397, May 1983. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Shardlow:1985:DDC

[SHP^{+}85] Tom Shardlow, Ray Hilborn, Randall M. Peterman, Gregory J. Steer, and Michael J. Bradford. Density-dependent catchability coefficients. Transactions of the American Fisheries Society, 114(3):436–440, May 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Shingleton:1981:TPN

[SHS81] Michael V. Shingleton, Charles H. Hocutt, and Jay R. Stauffer, Jr. Temperature preference of the new river shiner. Transactions of the American Fisheries Society, 110(5):660–661, September 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Scarnecchia:1989:ORI

[SÍW89] Dennis L. Scarnecchia, Árni Ísaksson, and S. E. White. Oceanic and riverine influences on variations in yield among Icelandic stocks of Atlantic salmon. Transactions of the American Fisheries Society, 118(5):482–494, September 1989. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Schneeberger:1981:UFL

[SJ81] Philip J. Schneeberger and David J. Jude. Use of fish larva morphometry to predict exclusion capabilities of small-mesh screens at cooling-water intakes. Transactions of the American Fisheries Society, 110(2):246–252, March 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Savitz:1982:UGS

[SJ82] J. Savitz and John Janssen. Utilization of green sunfish and bluegills by largemouth bass: Influence of ingestion time. Transactions of the American Fisheries Society, 111(4):462–464, July 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Schramm:1989:EMF

[SJ89a] Harold L. Schramm, Jr. and Kurt J. Jirka. Epiphytic macroinvertebrates as a food resource for bluegills in Florida lakes. Transactions of the American Fisheries Society, 118(4):416–426, July 1989. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:1989:GMF

[SJ89b] Barry D. Smith and Glen S. Jamieson. Growth of male and female Dungeness crabs near Tofino, British Columbia. Transactions of the American Fisheries Society, 118(5):556–563, September 1989. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stoffle:1987:CBS

[SJR87] Richard W. Stoffle, Florence V. Jensen, and Danny L. Rasch. Cultural basis of sport Anglers’ response to reduced lake trout catch limits. Transactions of the American Fisheries Society, 116(3): 503–509, May 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Serns:1981:RAE

[SK81] Steven L. Serns and James J. Kempinger. Relationship of angler exploitation to the size, age, and sex of walleyes in Escanaba Lake, Wisconsin. Transactions of the American Fisheries Society, 110 (2):216–220, March 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Spieler:1984:RF

[SK84] Richard E. Spieler and Robert L. Kendall. Rhythmicity in fishes. Transactions of the American Fisheries Society, 113(4):411–413, July 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Senta:1985:LJF

[SK85] Tetsushi Senta and Izumi Kinoshita. Larval and juvenile fishes occurring in surf zones of Western Japan. Transactions of the American Fisheries Society, 114(4):609–618, July 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stier:1986:URT

[SK86] David J. Stier and Boyd Kynard. Use of radio telemetry to determine the mortality of Atlantic salmon smolts passed through a 17-MW kaplan turbine at a low-head hydroelectric dam. Transactions of the American Fisheries Society, 115(5):771–775, September 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stewart:1981:FFT

[SKC81] Donald J. Stewart, James F. Kitchell, and Larry B. Crowder. Forage fishes and their salmonid predators in Lake Michigan. Transactions of the American Fisheries Society, 110(6):751–763, November 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stevens:1985:DSB

[SKMK85] Donald E. Stevens, David W. Kohlhorst, Lee W. Miller, and D. W. Kelley. The decline of striped bass in the sacramentosan joaquin estuary, California. Transactions of the American Fisheries Society, 114(1):12–30, January 1985. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sheehan:1986:IPA

[SL86] Robert J. Sheehan and William M. Lewis. Influence of pH and ammonia salts on ammonia toxicity and water balance in young channel catfish. Transactions of the American Fisheries Society, 115(6):891–899, November 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Schreiner:1984:SSR

[SLJ^{+}84] Donald R. Schreiner, James E. Luey, Lawrence D. Jacobson, Charles C. Krueger, and Ira R. Adelman. Stock structure of rainbow smelt in Western Lake Superior: Biochemical genetics. Transactions of the American Fisheries Society, 113(6):701–708, November 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Summerfelt:1984:TES

[SM84] Robert C. Summerfelt and Doyle Mosier. Transintestinal expulsion of surgically implanted dummy transmitters by channel catfish. Transactions of the American Fisheries Society, 113(6):760– 766, November 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:1984:TVB

[Smi84] Susan E. Smith. Timing of vertebral-band deposition in tetracycline-injected leopard sharks. Transactions of the American Fisheries Society, 113(3):308–313, May 1984. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:1985:SCW

[Smi85] Eric P. Smith. Statistical comparison of weighted overlap measures. Transactions of the American Fisheries Society, 114(2): 250–257, March 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:1989:RMS

[Smi89] David G. Smith. Role of museums in studies of early life history stages of fishes. Transactions of the American Fisheries Society, 118(2):214–217, March 1989. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stickney:1983:RYC

[SMLM83] Robert R. Stickney, Robert B. McGeachin, Donald H. Lewis, and Jon Marks. Response of young channel catfish to diets containing purified fatty acids. Transactions of the American Fisheries Society, 112(5):665–669, September 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Spieler:1984:EPF

[SN84] Richard E. Spieler and Teresa A. Noeske. Effects of photoperiod and feeding schedule on diel variations of locomotor activity, cortisol, and thyroxine in goldfish. Transactions of the American Fisheries Society, 113(4):528–539, July 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Snyder:1983:ICL

[Sny83] Darrel E. Snyder. Identification of catostomid larvae in Pyramid Lake and the Truckee River, Nevada. Transactions of the American Fisheries Society, 112(2):333–348, March 1983. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Snyder:1984:SLF

[Sny84] Darrel E. Snyder. The seventh larval fish conference. Transactions of the American Fisheries Society, 113(2):151–152, March 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sosiak:1982:BCB

[Sos82] A. J. Sosiak. Buoyancy comparisons between juvenile Atlantic salmon and brown trout of wild and hatchery origin. Transactions of the American Fisheries Society, 111(3):307–311, May 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sowden:1985:PRT

[SP85] Terry K. Sowden and G. Power. Prediction of rainbow trout embryo survival in relation to groundwater seepage and particle size of spawning substrates. Transactions of the American Fisheries Society, 114(6):804–812, November 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Strange:1987:NCC

[SP87] Richard J. Strange and Janice C. Pelton. Nutrient content of clupeid forage fishes. Transactions of the American Fisheries Society, 116(1):60–66, January 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Swanson:1988:PEF

[SP88] R. Lawrence Swanson and Charles A. Parker. Physical environmental factors contributing to recurring hypoxia in the New York Bight. Transactions of the American Fisheries Society, 117(1): 37–47, January 1988. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Schlesinger:1982:CMI

[SR82] David A. Schlesinger and Henry A. Regier. Climatic and morphoedaphic indices of fish yields from natural lakes. Transactions of the American Fisheries Society, 111(2):141–150, March 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sakanari:1983:TLP

[SRM83] Judy A. Sakanari, Carol A. Reilly, and Mike Moser. Tubercular lesions in Pacific Coast populations of striped bass. Transactions of the American Fisheries Society, 112(4):565–566, July 1983. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Savino:1982:PPI

[SS82a] Jacqueline F. Savino and Roy A. Stein. Predator-prey interaction between largemouth bass and bluegills as influenced by simulated, submersed vegetation. Transactions of the American Fisheries Society, 111(3):255–266, May 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sower:1982:VIF

[SS82b] Stacia A. Sower and Carl B. Schreck. In vitro induction of final maturation of oocytes from Coho salmon. Transactions of the American Fisheries Society, 111(3):399–403, May 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Specker:1984:IAS

[SS84] Jennifer L. Specker and Carl B. Schreck. Influence of ambient salts on plasma thyroxine concentration in juvenile Coho salmon. Transactions of the American Fisheries Society, 113(1):90–94, January 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Scarano:1984:HRS

[SSGT84] G. Scarano, M. G. Saroglia, R. H. Gray, and E. Tibaldi. Hematological responses of sea bass Dicentrarchus labrax to sublethal nitrite exposures. Transactions of the American Fisheries Society, 113(3):360–364, May 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Scott:1986:EUD

[SSS86] J. B. Scott, C. R. Steward, and Q. J. Stober. Effects of urban development on fish population dynamics in Kelsey Creek, Washington. Transactions of the American Fisheries Society, 115(4): 555–567, July 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Seeb:1986:UGM

[SSU86] J. E. Seeb, L. W. Seeb, and F. M. Utter. Use of genetic marks to assess stock dynamics and management programs for chum salmon. Transactions of the American Fisheries Society, 115(3): 448–454, May 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Schwedler:1983:ERB

[ST83] Thomas E. Schwedler and Craig S. Tucker. Empirical relationship between percent methemoglobin in channel catfish and dissolved nitrite and chloride in ponds. Transactions of the American Fisheries Society, 112(1):117–119, January 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stanley:1988:NSD

[Sta88] David R. Stanley. Notes: Sexual dimorphism of pelvic fin shape in four species of Catostomidae. Transactions of the American Fisheries Society, 117(6):600–602, November 1988. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stickney:1989:SFH

[Sti89] Robert R. Stickney. Salmon farming handbook. Transactions of the American Fisheries Society, 118(3):345–348, May 1989. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Storck:1986:IGS

[Sto86] Ted W. Storck. Importance of gizzard shad in the diet of largemouth bass in lake shelbyville, Illinois. Transactions of the American Fisheries Society, 115(1):21–27, January 1986. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Spigarelli:1982:FGF

[STP82] S. A. Spigarelli, M. M. Thommes, and W. Prepejchal. Feeding, growth, and fat deposition by brown trout in constant and fluctuating temperatures. Transactions of the American Fisheries Society, 111(2):199–209, March 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Strauss:1982:IRS

[Str82] Richard E. Strauss. Influence of replicated subsamples and subsample heterogeneity on the linear index of food selection. Transactions of the American Fisheries Society, 111(4):517–522, July 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Scheerer:1987:PDS

[STS87] Paul D. Scheerer, Gary H. Thorgaard, and James E. Seeb. Performance and developmental stability of triploid tiger trout (brown trout [female sign] × brook trout [male sign]). Transactions of the American Fisheries Society, 116(1):92–97, January 1987. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:1985:BSK

[SW85] Joseph W. Smith and Charles A. Wenner. Biology of the southern kingfish in the South Atlantic bight. Transactions of the American Fisheries Society, 114(3):356–366, May 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Seelbach:1988:ICW

[SW88a] Paul W. Seelbach and Gary E. Whelan. Identification and contribution of wild and hatchery steelhead stocks in Lake Michigan tributaries. Transactions of the American Fisheries Society, 117 (5):444–451, September 1988. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sigismondi:1988:CAR

[SW88b] Linda A. Sigismondi and Lavern J. Weber. Changes in avoidance response time of juvenile Chinook salmon exposed to multiple acute handling stresses. Transactions of the American Fisheries Society, 117(2):196–201, March 1988. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Scoppettone:1983:REC

[SWCB83] G. Gary Scoppettone, Gary A. Wedemeyer, Mark Coleman, and Howard Burge. Reproduction by the endangered cui-ui in the Lower Truckee River. Transactions of the American Fisheries Society, 112(6):788–793, November 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Salvatore:1987:EWT

[SWM87] Scott R. Salvatore, Thomas E. Wissing, and Neal D. Mundahl. Effect of water temperature on food evacuation rate and feeding activity of age-0 gizzard shad. Transactions of the American Fisheries Society, 116(1):67–70, January 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Shuter:1985:AEM

[SWRM85] B. J. Shuter, D. A. Wismer, H. A. Regier, and J. E. Matuszek. An application of ecological modelling: Impact of thermal effluent on a smallmouth bass population. Transactions of the American Fisheries Society, 114(5):631–651, September 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Shaw:1985:TLG

[SWT^{+}85] Richard F. Shaw, William J. Wiseman, Jr., R. Eugene Turner, Lawrence J. Rouse, Jr., Richard E. Condrey, and Francis J. Kelly, Jr. Transport of larval Gulf menhaden Brevoortia patronus in Continental Shelf waters of Western Louisiana: a hypothesis. Transactions of the American Fisheries Society, 114(4):452–460, July 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Schramm:1985:ECP

[SZ85] Harold L. Schramm, Jr. and Alexander V. Zale. Effects of cover and prey size on preferences of juvenile largemouth bass for blue tilapias and bluegills in tanks. Transactions of the American Fisheries Society, 114(5):725–731, September 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Shrode:1982:EST

[SZS82] Joy B. Shrode, Kim E. Zerba, and John S. Stephens, Jr. Ecological significance of temperature tolerance and preference of some inshore California fishes. Transactions of the American Fisheries Society, 111(1):45–51, January 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Talhelm:1987:RSS

[Tal87] Daniel R. Talhelm. Recommendations from the SAFR symposium. Transactions of the American Fisheries Society, 116(3): 537–540, May 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tanaka:1985:FAI

[Tan85] Masaru Tanaka. Factors affecting the inshore migration of pelagic larval and demersal juvenile red sea bream Pagrus major to a nursery ground. Transactions of the American Fisheries Society, 114(4):471–477, July 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Talbot:1986:FSC

[TAS86] C. W. Talbot, K. W. Able, and J. K. Shisler. Fish species composition in New Jersey salt marshes: Effects of Marsh alterations for mosquito control. Transactions of the American Fisheries Society, 115(2):269–278, March 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Taylor:1984:LSF

[Tay84] Malcolm H. Taylor. Lunar synchronization of fish reproduction. Transactions of the American Fisheries Society, 113(4):484–493, July 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Taylor:1988:WTV

[Tay88] Eric B. Taylor. Water temperature and velocity as determinants of microhabitats of juvenile Chinook and Coho salmon in a laboratory stream channel. Transactions of the American Fisheries Society, 117(1):22–28, January 1988. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Teska:1981:ZPL

[TB81] Jeffrey D. Teska and David J. Behmer. Zooplankton preference of larval lake whitefish. Transactions of the American Fisheries Society, 110(3):459–461, May 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Trippel:1987:CPI

[TB87] E. A. Trippel and F. W. H. Beamish. Characterizing piscivory from ingested remains. Transactions of the American Fisheries Society, 116(5):773–776, September 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tschirhart:1987:EVE

[TC87] John Tschirhart and Thomas D. Crocker. Economic valuation of ecosystems. Transactions of the American Fisheries Society, 116(3):469–478, May 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tiersch:1989:CEI

[TC89] Terrence R. Tiersch and Robert W. Chandler. Chicken erythrocytes as an internal reference for analysis of DNA content by flow cytometry in grass carp. Transactions of the American Fisheries Society, 118(6):713–717, November 1989. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Thayer:1983:ELF

[TCKW83] Gordon W. Thayer, David R. Colby, Martin A. Kjelson, and Michael P. Weinstein. Estimates of larval-fish abundance: Diurnal variation and influences of sampling gear and towing speed. Transactions of the American Fisheries Society, 112(2):272–279, March 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Teichert-Coddington:1988:LRT

[TCS88] David R. Teichert-Coddington and R. Oneal Smitherman. Lack of response by Tilapia nilotica to mass selection for rapid early growth. Transactions of the American Fisheries Society, 117(3): 297–300, May 1988. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Talhelm:1987:ISA

[TDG^{+}87] Daniel R. Talhelm, Jack A. Donnan, Delano R. Graff, A. P. Grima, Lawrence W. Libby, Scott Milliman, and Margaret Ross Dochoda. Introduction to the social assessment of fisheries resources proceedings: SAFR steering committee. Transactions of the American Fisheries Society, 116(3):289–292, May 1987. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Thomas:1986:EFT

[TGC^{+}86] Robert E. Thomas, Jessica A. Gharrett, Mark G. Carls, Stanley D. Rice, Adam Moles, and Sid Korn. Effects of fluctuating temperature on mortality, stress, and energy reserves of juvenile Coho salmon. Transactions of the American Fisheries Society, 115(1):52–59, January 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Thune:1982:EPE

[TGRA82a] R. L. Thune, T. E. Graham, L. M. Riddle, and R. L. Amborski. Extracellular products and endotoxin from Aeromonas hydrophila: Effects on age-0 channel catfish. Transactions of the American Fisheries Society, 111(3):404–408, May 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Thune:1982:EPA

[TGRA82b] R. L. Thune, T. E. Graham, L. M. Riddle, and R. L. Amborski. Extracellular proteases from Aeromonas hydrophila: Partial purification and effects on age-0 channel catfish. Transactions of the American Fisheries Society, 111(6):749–754, November 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Thorogood:1986:NTS

[Tho86] John Thorogood. New technique for sampling otoliths of sashimigrade scombrid fishes. Transactions of the American Fisheries Society, 115(6):913–914, November 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Talhelm:1987:PTC

[THV87] Daniel R. Talhelm, J. Edward Hanna, and Peter Victor. Product travel cost approach: Estimating acid rain damage to sportfishing in Ontario. Transactions of the American Fisheries Society, 116(3):420–431, May 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tisa:1989:PFA

[Tis89] Mark S. Tisa. Present and future Atlantic salmon management: Measuring progress toward international cooperation. Transactions of the American Fisheries Society, 118(5):576–577, September 1989. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tin:1983:DGL

[TJ83] Heang T. Tin and David J. Jude. Distribution and growth of larval rainbow smelt in Eastern Lake Michigan, 1978–1981. Transactions of the American Fisheries Society, 112(4):517–524, July 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Thorgaard:1981:PIH

[TJS81] Gary H. Thorgaard, Mary Ellen Jazwin, and Alton R. Stier. Polyploidy induced by heat shock in rainbow trout. Transactions of the American Fisheries Society, 110(4):546–550, July 1981. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Taylor:1985:MJA

[TK85] Ralph E. Taylor and Boyd Kynard. Mortality of juvenile American shad and blueback herring passed through a low-head kaplan hydroelectric turbine. Transactions of the American Fisheries Society, 114(3):430–435, May 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Talhelm:1987:STV

[TL87] Daniel R. Talhelm and Lawrence W. Libby. In search of a total value assessment framework: SAFR symposium overview and synthesis. Transactions of the American Fisheries Society, 116(3): 293–301, May 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Taylor:1985:BSS

[TM85a] Eric B. Taylor and J. D. McPhail. Burst swimming and sizerelated predation of newly emerged Coho salmon Oncorhynchus kisutch. Transactions of the American Fisheries Society, 114(4): 546–551, July 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Taylor:1985:OSR

[TM85b] Eric B. Taylor and J. D. McPhail. Ontogeny of the startle response in young Coho salmon Oncorhynchus kisutch. Transactions of the American Fisheries Society, 114(4):552–557, July 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tonn:1983:CAF

[TMF83] William M. Tonn, John J. Magnuson, and Anne M. Forbes. Community analysis in fishery management: an application with Northern Wisconsin lakes. Transactions of the American Fisheries Society, 112(3):368–377, May 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Thurston:1983:ATAa

[TR83] Robert V. Thurston and Rosemarie C. Russo. Acute toxicity of ammonia to rainbow trout. Transactions of the American Fisheries Society, 112(5):696–704, September 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Todd:1989:MHU

[TR89] Brian L. Todd and Charles F. Rabeni. Movement and habitat use by stream-dwelling smallmouth bass. Transactions of the American Fisheries Society, 118(3):229–242, May 1989. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Thurston:1984:CTA

[TRL^{+}84] Robert V. Thurston, Rosemarie C. Russo, Robert J. Luedtke, Charlie E. Smith, Elizabeth L. Meyn, Charles Chakoumakos, Kenneth C. Wang, and C. J. D. Brown. Chronic toxicity of ammonia to rainbow trout. Transactions of the American Fisheries Society, 113(1):56–73, January 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Thurston:1986:CTA

[TRM^{+}86] Robert V. Thurston, Rosemarie C. Russo, Elizabeth L. Meyn, Richard K. Zajdel, and Charlie E. Smith. Chronic toxicity of ammonia to fathead minnows. Transactions of the American Fisheries Society, 115(2):196–207, March 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Trotter:1989:CCT

[Tro89] Patrick C. Trotter. Coastal cutthroat trout: a life history compendium. Transactions of the American Fisheries Society, 118 (5):463–473, September 1989. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Thurston:1983:ATAb

[TRP83] Robert V. Thurston, Rosemarie C. Russo, and Glenn R. Phillips. Acute toxicity of ammonia to fathead minnows. Transactions of the American Fisheries Society, 112(5):705–711, September 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Thrower:1984:FAR

[TS84] F. P. Thrower and W. W. Smoker. First adult return of pink salmon tagged as emergents with binary-coded wires. Transactions of the American Fisheries Society, 113(6):803–804, November 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tomcko:1984:PTM

[TSC84] Cynthia M. Tomcko, Roy A. Stein, and Robert F. Carline. Predation by tiger muskellunge on bluegill: Effects of predator experience, vegetation, and prey density. Transactions of the American Fisheries Society, 113(5):588–594, September 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Timmons:1981:EGM

[TSD81] Tom J. Timmons, W. L. Shelton, and W. D. Davies. Early growth and mortality of largemouth bass in West point reservoir, Alabama-Georgia. Transactions of the American Fisheries Society, 110(4):489–494, July 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Taubert:1982:VFA

[TT82] Bruce D. Taubert and John A. Tranquilli. Verification of the formation of annuli in otoliths of largemouth bass. Transactions of the American Fisheries Society, 111(4):531–534, July 1982. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tzeng:1989:VDG

[TY89] W.-N. Tzeng and S.-Y. Yu. Validation of daily growth increments in otoliths of milkfish larvae by oxytetracycline labeling. Transactions of the American Fisheries Society, 118(2):168–174, March 1989. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tyler:1988:DMF

[Tyl88] A. V. Tyler. Dynamics of marine fish populations. Transactions of the American Fisheries Society, 117(1):95–101, January 1988. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tyus:1987:DRH

[Tyu87] Harold M. Tyus. Distribution, reproduction, and habitat use of the razorback sucker in the Green River, Utah, 1979–1986. Transactions of the American Fisheries Society, 116(1):111–116, January 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Uphoff:1989:EES

[Uph89] James H. Uphoff, Jr. Environmental effects on survival of eggs, larvae, and juveniles of striped bass in the Choptank River, Maryland. Transactions of the American Fisheries Society, 118(3): 251–263, May 1989. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Usher:1987:BR

[Ush87a] Anthony J. Usher. Book reviews. Transactions of the American Fisheries Society, 116(5):779–784, September 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Usher:1987:OLW

[Ush87b] Anthony J. Usher. Ontario lake of the Woods fishery: Economic and social analysis. Transactions of the American Fisheries Society, 116(3):352–366, May 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Vouglitois:1987:LHP

[VAKT87] James J. Vouglitois, Kenneth W. Able, Robert J. Kurtz, and Kenneth A. Tighe. Life history and population dynamics of the bay anchovy in New Jersey. Transactions of the American Fisheries Society, 116(2):141–153, March 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Vanderpool:1987:SAT

[Van87a] Christopher K. Vanderpool. Social action, total economic value, and environmental policy: The problem of rationality. Transactions of the American Fisheries Society, 116(3):336–338, May 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Vanderpool:1987:SIA

[Van87b] Christopher K. Vanderpool. Social impact assessment and fisheries. Transactions of the American Fisheries Society, 116(3): 479–485, May 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Vetter:1982:SMC

[Vet82] Russell D. Vetter. Seasonal metabolic compensation in sympatric seatrout: Adaptation to the estuary. Transactions of the American Fisheries Society, 111(2):193–198, March 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Vetter:1986:CGA

[VHH86] Russell D. Vetter, Huey-Min Hwang, and Robert E. Hodson. Comparison of glycogen and adenine nucleotides as indicators of metabolic stress in mummichogs. Transactions of the American Fisheries Society, 115(1):47–51, January 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Vollestad:1986:LHC

[VJ86] Leif Asbjørn Vøllestad and Bror Jonsson. Life-history characteristics of the European eel Anguilla anguilla in the Imsa River, Norway. Transactions of the American Fisheries Society, 115(6): 864–871, November 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Vogelbein:1987:HIA

[VO87] Wolfgang K. Vogelbein and Robin M. Overstreet. Histopathology of the internal anchor tag in spot and spotted seatrout. Transactions of the American Fisheries Society, 116(5):745–756, September 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

VanDenAvyle:1988:ITN

[VP88] Michael J. Van Den Avyle and Raymond W. Petering. Inundated timber as nursery habitat for larval gizzard and threadfin shad in a new pumped storage reservoir. Transactions of the American Fisheries Society, 117(1):84–89, January 1988. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wallace:1981:ADO

[Wal81] Richard K. Wallace, Jr. An assessment of diet-overlap indexes. Transactions of the American Fisheries Society, 110(1):72–76, January 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Waldman:1986:DVM

[Wal86a] John R. Waldman. Diagnostic value of Morone dentition. Transactions of the American Fisheries Society, 115(6):900–907, November 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wallus:1986:PRC

[Wal86b] Robert Wallus. Paddlefish reproduction in the Cumberland and Tennessee River Systems. Transactions of the American Fisheries Society, 115(3):424–428, May 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Waring:1984:AGM

[War84] Gordon T. Waring. Age, growth, and mortality of the little skate off the Northeast Coast of the United States. Transactions of the American Fisheries Society, 113(3):314–321, May 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Waters:1983:RBT

[Wat83] Thomas F. Waters. Replacement of brook trout by brown trout over 15 years in a Minnesota stream: Production and abundance. Transactions of the American Fisheries Society, 112(2):137–146, March 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wydoski:1981:FSL

[WB81] Richard S. Wydoski and David H. Bennett. Forage species in lakes and reservoirs of the Western United States. Transactions of the American Fisheries Society, 110(6):764–771, November 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wilmot:1985:GDA

[WB85] Richard L. Wilmot and Carl V. Burger. Genetic differences among populations of Alaskan sockeye salmon. Transactions of the American Fisheries Society, 114(2):236–243, March 1985. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wilson:1987:CFM

[WBD87] Charles A. Wilson, Daniel W. Beckman, and John Mark Dean. Calcein as a fluorescent marker of otoliths of larval and juvenile fish. Transactions of the American Fisheries Society, 116(4):668– 670, July 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Watenpaugh:1985:TTN

[WBH85] D. E. Watenpaugh, T. L. Beitinger, and D. W. Huey. Temperature tolerance of nitrite-exposed channel catfish. Transactions of the American Fisheries Society, 114(2):274–278, March 1985. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wurtsbaugh:1983:GAJ

[WC83] Wayne A. Wurtsbaugh and Joseph J. Cech, Jr. Growth and activity of juvenile mosquitofish: Temperature and ration effects. Transactions of the American Fisheries Society, 112(5):653–660, September 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Williamson:1986:DRH

[WC86] J. Holt Williamson and Gary J. Carmichael. Differential response to handling stress by Florida, northern, and hybrid largemouth bass. Transactions of the American Fisheries Society, 115(5):756– 761, September 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wicker:1988:MBS

[WCM88] Anton M. Wicker, Benjamin M. Currin, and John M. Miller. Migration of brown shrimp in nursery areas of Pamlico Sound, North Carolina, as indicated by changes in length-frequency distributions. Transactions of the American Fisheries Society, 117(1): 92–94, January 1988. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Williamson:1986:NBG

[WCSM86] J. Holt Williamson, Gary J. Carmichael, Maureen E. Schmidt, and Donald C. Morizot. New biochemical genetic markers for largemouth bass. Transactions of the American Fisheries Society, 115(3):460–465, May 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wangila:1988:IGT

[WD88] B. C. C. Wangila and T. A. Dick. Influence of genotype and temperature on the relationship between specific growth rate and size of rainbow trout. Transactions of the American Fisheries Society, 117(6):560–564, November 1988. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wismer:1985:EMS

[WDS85] Donald A. Wismer, Donald L. Deangelis, and B. J. Shuter. An empirical model of size distributions of smallmouth bass. Transactions of the American Fisheries Society, 114(5):737–742, September 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wetterer:1988:ELP

[Wet88] James K. Wetterer. Effect of localized prey depletion on prey selectivity by fish. Transactions of the American Fisheries Society, 117(3):313–316, May 1988. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Willis:1981:ICL

[WF81] David W. Willis and Stephen A. Flickinger. Intensive culture of largemouth bass fry. Transactions of the American Fisheries Society, 110(5):650–655, September 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Woodward:1989:SES

[WFM^{+}89] Daniel F. Woodward, Aida M. Farag, Mary E. Mueller, Edward E. Little, and Frank A. Vertucci. Sensitivity of endemic Snake River cutthroat trout to acidity and elevated aluminum. Transactions of the American Fisheries Society, 118(6):630–643, November 1989. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wetherbee:1987:XRO

[WGR87] Bradley M. Wetherbee, Samuel H. Gruber, and Anna L. Ramsey. X-radiographic observations of food passage through digestive tracts of lemon sharks. Transactions of the American Fisheries Society, 116(5):763–767, September 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Weithman:1982:SVT

[WH82a] A. Stephen Weithman and Mark A. Haas. Socioeconomic value of the trout fishery in Lake Taneycomo, Missouri. Transactions of the American Fisheries Society, 111(2):223–230, March 1982. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wiener:1982:GCB

[WH82b] James G. Wiener and Wendy R. Hanneman. Growth and condition of bluegills in Wisconsin lakes: Effects of population density and Lake pH. Transactions of the American Fisheries Society, 111 (6):761–767, November 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Weithman:1984:EDO

[WH84] A. Stephen Weithman and Mark A. Haas. Effects of dissolvedoxygen depletion on the rainbow trout fishery in Lake Taneycomo, Missouri. Transactions of the American Fisheries Society, 113(2): 109–124, March 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wicker:1987:RAF

[WJ87] Anton M. Wicker and William E. Johnson. Relationships among fat content, condition factor, and first-year survival of Florida largemouth bass. Transactions of the American Fisheries Society, 116(2):264–271, March 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wolters:1981:ITC

[WLC81] William R. Wolters, George S. Libey, and Charles L. Chrisman. Induction of triploidy in channel catfish. Transactions of the American Fisheries Society, 110(2):310–312, March 1981. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wolters:1982:ETG

[WLC82] William R. Wolters, George S. Libey, and C. Larry Chrisman. Effect of triploidy on growth and gonad development of channel catfish. Transactions of the American Fisheries Society, 111 (1):102–105, January 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Witzel:1983:RSS

[WM83] Larry D. Witzel and Hugh R. Maccrimmon. Redd-site selection by brook trout and brown trout in Southwestern Ontario streams. Transactions of the American Fisheries Society, 112(6):760–771, November 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Weld:1984:CRG

[WM84] Madeline M. Weld and Albert H. Meier. Circadian responses of gonads and fat stores to handling of Gulf killifish. Transactions of the American Fisheries Society, 113(4):521–527, July 1984. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wertheimer:1986:GSA

[WM86] Alex C. Wertheimer and Roy M. Martin. Growth and survival to adulthood of chum salmon vaccinated as juveniles against vibriosis. Transactions of the American Fisheries Society, 115(2): 343–346, March 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Woodward:1981:ASE

[WMM81] Daniel F. Woodward, Paul M. Mehrle, Jr., and Wilbur L. Mauck. Accumulation and sublethal effects of a Wyoming crude oil in cutthroat trout. Transactions of the American Fisheries Society, 110(3):437–445, May 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Weisberg:1986:ENI

[WMRH86] Stephen B. Weisberg, Richard P. Morin, Eric A. Ross, and Michael F. Hirshfield. Eustrongylides (Nematoda) infection in mummichogs and other fishes of the Chesapeake Bay region. Transactions of the American Fisheries Society, 115(5):776–783, September 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Weinstein:1986:EMJ

[WO86] M. P. Weinstein and S. P. O’Neil. Exchange of marked juvenile spots between adjacent tidal creeks in the York River Estuary, Virginia. Transactions of the American Fisheries Society, 115(1): 93–97, January 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wright:1983:NEZ

[WOL83] David I. Wright, W. John O’Brien, and Chris Luecke. A new estimate of zooplankton retention by gill rakers and its ecological significance. Transactions of the American Fisheries Society, 112 (5):638–646, September 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Westin:1985:EPD

[WOR85] Deborah T. Westin, Charles E. Olney, and Bruce A. Rogers. Effects of parental and dietary organochlorines on survival and body burdens of striped bass larvae. Transactions of the American Fisheries Society, 114(1):125–136, January 1985. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Whitman:1982:ISV

[WQB82] Randall P. Whitman, Thomas P. Quinn, and Ernest L. Brannon. Influence of suspended volcanic ash on homing behavior of adult Chinook salmon. Transactions of the American Fisheries Society, 111(1):63–69, January 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wrenn:1984:SBR

[Wre84] William B. Wrenn. Smallmouth bass reproduction in elevated temperature regimes at the Species’ native southern limit. Transactions of the American Fisheries Society, 113(3):295–303, May 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Woodward:1985:LRO

[WRH^{+}85] D. F. Woodward, R. G. Riley, M. G. Henry, J. S. Meyer, and T. R. Garland. Leaching of retorted oil shale: Assessing the toxicity to Colorado squawfish, fathead minnows, and two food-chain organisms. Transactions of the American Fisheries Society, 114 (6):887–894, November 1985. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wagner:1982:DAP

[WS82] Wilbert C. Wagner and Thomas M. Stauffer. Distribution and abundance of pink salmon in Michigan tributaries of the great lakes, 1967–1980. Transactions of the American Fisheries Society, 111(4):523–526, July 1982. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Whitworth:1983:GPS

[WS83] Wilbur E. Whitworth and Richard J. Strange. Growth and production of sympatric brook and rainbow trout in an Appalachian stream. Transactions of the American Fisheries Society, 112(4): 469–475, July 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Woodward:1987:PSR

[WS87] Colleen Caldwell Woodward and Richard J. Strange. Physiological stress responses in wild and hatchery-reared rainbow trout. Transactions of the American Fisheries Society, 116(4):574–579, July 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wahl:1988:SPT

[WS88] David H. Wahl and Roy A. Stein. Selective predation by three esocids: The role of prey behavior and morphology. Transactions of the American Fisheries Society, 117(2):142–151, March 1988. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Weiner:1986:ELP

[WSL86] Gary S. Weiner, Carl B. Schreck, and Hiram W. Li. Effects of low pH on reproduction of rainbow trout. Transactions of the American Fisheries Society, 115(1):75–82, January 1986. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wurstbaugh:1988:MMF

[WT88] Wayne A. Wurstbaugh and Rene Alfaro Tapia. Mass mortality of fishes in Lake Titicaca (Peru–Bolivia) associated with the protozoan parasite Ichthyophthirius multifiliis. Transactions of the American Fisheries Society, 117(2):213–217, March 1988. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Webb:1986:FLM

[WW86a] Paul W. Webb and Daniel Weihs. Functional locomotor morphology of early life history stages of fishes. Transactions of the American Fisheries Society, 115(1):115–127, January 1986. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wiley:1986:EBD

[WW86b] Michael J. Wiley and Lynn D. Wike. Energy balances of diploid, triploid, and hybrid grass carp. Transactions of the American Fisheries Society, 115(6):853–863, November 1986. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Yarbrough:1987:UPT

[Yar87] C. Jarrell Yarbrough. Using political theory in fishery management. Transactions of the American Fisheries Society, 116(3): 532–536, May 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Youngs:1982:ACM

[YH82] William D. Youngs and Douglas G. Heimbuch. Another consideration of the morphoedaphic index. Transactions of the American Fisheries Society, 111(2):151–153, March 1982. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Young:1989:SAS

[YHW89] Michael K. Young, Wayne A. Hubert, and Thomas A. Wesche. Substrate alteration by spawning brook trout in a Southeastern Wyoming stream. Transactions of the American Fisheries Society, 118(4):379–385, July 1989. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Young:1983:FPF

[YMH83] Leroy M. Young, James P. Monaghan, Jr., and Roy C. Heidinger. Food preferences, food intake, and growth of the F_{1} hybrid of grass carp [female sign] × bighead carp [male sign]. Transactions of the American Fisheries Society, 112(5):661–664, September 1983. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Yasutake:1986:PBE

[YMH86] W. T. Yasutake, J. D. McIntyre, and A. R. Hemmingsen. Parasite burdens in experimental families of Coho salmon. Transactions of the American Fisheries Society, 115(4):636–640, July 1986. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Zohar:1984:ADC

[ZB84] Y. Zohar and R. Billard. Annual and daily changes in plasma gonadotropin and sex steroids in relation to teleost gonad cycles. Transactions of the American Fisheries Society, 113(4):444–451, July 1984. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Zale:1989:ESC

[ZG89] Alexander V. Zale and Richard W. Gregory. Effect of salinity on cold tolerance of juvenile blue tilapias. Transactions of the American Fisheries Society, 118(6):718–720, November 1989. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Zimmerman:1981:IHM

[ZR81] Earl G. Zimmerman and M. Carol Richmond. Increased heterozygosity at the mdh-b locus in fish inhabiting a rapidly fluctuating thermal environment. Transactions of the American Fisheries Society, 110(3):410–416, May 1981. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Zhang:1988:BBC

[ZS88] Chang Ik Zhang and Patrick J. Sullivan. Biomass-based cohort analysis that incorporates growth. Transactions of the American Fisheries Society, 117(2):180–189, March 1988. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Zinkl:1987:TBC

[ZSNC87] Joseph G. Zinkl, Patrick J. Shea, Rodney J. Nakamoto, and Josh Callman. Technical and biological considerations for the analysis of brain cholinesterase of rainbow trout. Transactions of the American Fisheries Society, 116(4):570–573, July 1987. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).